### v2

# Pedro Montealegre Universidad Adolfo Ibañez, Santiago, Chile Ivan Rapaport Universidad de Chile, Santiago, Chile Ioan Todinca LIFO, Université d’Orléans and INSA Centre-Val de Loire, Orléans, France

Abstract
The sparsity theory of Nešetřil and Ossona de Mendez provides generic combinatorial and algorithmic methods suitable for many classes of sparse graphs. As such, it has already driven major advances in model checking, parameterized complexity, and approximation algorithms. We show that this theory is equally well suited to the design of distributed algorithms, providing the first generic algorithmic results for distributed first-order (FO) model checking on graphs of bounded expansion. The latter form a rich graph family which subsumes planar, bounded-genus, bounded-treedepth, bounded-treewidth, and bounded-degree graphs, as well as graphs excluding a fixed minor or topological minor, sparse Erdős–Rényi graphs (a.a.s.), and several network models such as stochastic block models for suitable parameter ranges.
Our main results are the following algorithmic meta-theorems for distributed algorithms in the standard CONGEST model on every graph class $\mathcal{G}$ of bounded expansion.
  • First, we resolve an open problem originally posed by Nešetřil and Ossona de Mendez (Distributed Computing 2016), and reiterated by Pilipczuk, Siebertz, and Toruńczyk (LICS 2018). Recall that a formula $\varphi(x)$ is local if the satisfaction of $\varphi(x)$ depends only on the $r$-neighbourhood of its free variable $x$, for some fixed $r$. For instance, the formula “$x$ belongs to a triangle” is local. We show that, for every local FO formula $\varphi(x)$, there exists a deterministic algorithm that, for every $n$-vertex graph $G\in\mathcal{G}$, identifies all vertices $v\in V(G)$ such that $G\models\varphi(v)$, in $\mathcal{O}(\log n)$ rounds.
  • Second, we show that, for every FO formula $\varphi$, there is a deterministic algorithm that, for every $n$-vertex graph $G\in\mathcal{G}$, decides whether $G\models\varphi$, in $\mathcal{O}(D+\log n)$ rounds, where $D$ is the diameter of $G$. Note that the dependence on $D$ is unavoidable, as even the simple FO formula “there exist two vertices of degree 3” needs $\Omega(D)$ rounds in CONGEST, even on trees.
The techniques behind these meta-theorems extend to distributed counting, optimization, and certification problems on graphs of bounded expansion.
1 Introduction
1.1 Context and Objective
What is a sparse graph? The answer to this question depends on the perspective. For instance, the graph class in which every graph is obtained from a complete graph by subdividing every edge once is “combinatorially sparse” (it has a number of edges linear in the number of vertices), but is not necessarily “effectively sparse” from an algorithmic perspective. The sparsity theory by Nešetřil and Ossona de Mendez [53, 54] provided us with an approach of sparsity that can be effectively used algorithmically.
1.1.1 Graphs of Bounded Expansion
Within the sparsity theory, a crucial role is assigned to the graph classes of bounded expansion. Intuitively, a graph class $\mathcal{G}$ has bounded expansion if there exists a function $f:\mathbb{N}\to\mathbb{N}$ such that, for every $G\in\mathcal{G}$, and any $r\geq 0$, every minor $H$ of $G$ obtained by contracting disjoint subgraphs of radius at most $r$ is $f(r)$-degenerate (i.e., every subgraph of $H$ has a vertex of degree at most $f(r)$). Bounded expansion can be thought of as strengthened degeneracy, which must persist even after performing “local minor operations”. For instance, the aforementioned graph class containing every graph obtained from a complete graph by subdividing every edge once is not of bounded expansion.
Graph classes with bounded expansion form a rich graph family (see Figure 1). For instance, the class of planar graphs, the class of graphs with bounded genus, the class of graphs of bounded degree, and the class of minor-free graphs and topologically minor-free graphs are all of bounded expansion, as well as sparse Erdős–Rényi graphs (a.a.s.). Moreover, [15] demonstrated that many network models (e.g., stochastic block models) exhibit bounded expansion for some ranges of parameters, and these findings are even supported with empirical measurements on a corpus of real-world networks. All these points make graph classes of bounded expansion quite attractive from both conceptual and practical perspectives. In particular, it has already driven major advances in model checking [18], parameterized complexity [53], kernelization [16, 20] and approximation algorithms [43].
Bounded degeneracyBounded expansionExcluded topological subgraphExcluded minorBounded degreeBounded genusBounded treewidthPlanarBounded treedepth
Figure 1: Different classes of sparse graphs, with their mutual inclusion relations. A arrow from a class $\mathcal{G}$ to a class $\mathcal{G}^{\prime}$ indicates that $\mathcal{G}^{\prime}$ is (strictly) included in $\mathcal{G}$.
1.1.2 Distributed Computing in Graphs of Bounded Expansion
Nešetřil and Ossona de Mendez also showed that bounded expansion can also be approached from a distributed computing perspective (see [55]). Their finding is based on the following characterization [54]. A class $\mathcal{G}$ of graphs has bounded expansion if and only if there exists a function $g:\mathbb{N}\to\mathbb{N}$ such that, for every $G\in\mathcal{G}$, and for every $k\in\mathbb{N}$, there exists a coloring of the vertices of $G$ with colors in $[g(k)]=\{1,\dots,g(k)\}$ such that, for every set $U\subseteq[g(k)]$ of cardinality at most $k$, the subgraph of $G$ induced by the nodes colored with the colors in $U$ has treedepth at most $|U|$. (Recall that a graph $G=(V,E)$ has treedepth at most $d$ if there exists a forest $F=(V,E^{\prime})$ of rooted trees, on the same set of vertices as $G$, such that (1) every tree in $F$ has depth at most $d$, and (2) for every edge in $E$, one of its end points is an ancestor of the other in some tree of $F$.)
Nešetřil and Ossona de Mendez [55] established that for every graph class $\mathcal{G}$ of bounded expansion $f$, there exists a distributed algorithm which, for every positive integer $k$, and for every $G\in\mathcal{G}$, computes a $g(k)$-coloring of the nodes of $G$ satisfying the desired treedepth property, in $\mathcal{O}(\log n)$ rounds in the CONGEST model [^2] The CONGEST model [57] is a standard distributed computing model in which the nodes of the graph perform in a sequence of synchronous rounds, during each of which every pair of adjacent nodes can exchange a message on $\mathcal{O}(\log n)$ bits, e.g., a constant number of node identifiers.. The function $g$ depends on the given bound $f$ on the expansion of $\mathcal{G}$.
1.1.3 Deciding Local First-Order Formulas in CONGEST
The fact that a coloring satisfying the requirement of bounded expansion can be computed in a distributed manner yields the question of whether this ability can be used as a building block for solving “hard problems” efficiently in CONGEST, in graph classes of bounded expansion. In the context of CONGEST, “hard problems” are usually thought as problems requiring at least $\mbox{poly}(n)$ rounds to be solved, whereas “easy problems” are problems solvable in $\mbox{poly}\log(n)$ rounds. As a concrete illustration, it is known [17] that deciding $C_{4}$-freeness is hard in CONGEST, as it requires $\tilde{\Omega}(\sqrt{n})$ rounds — in the distributed setting, deciding a property asks all nodes to accept if the property is satisfied, and at least one node to reject if the property is violated. Is it possible to overcome the $\tilde{\Omega}(\sqrt{n})$ lower bound for deciding $C_{4}$-freeness in graphs of bounded expansion?
Nešetřil and Ossona de Mendez [55] actually answered positively to this question, by highlighting a general approach for deciding $H$-freeness, in $\mathcal{O}(\log n)$ rounds, for any fixed (connected) graph $H$. In fact, this approach can even be used to mark all nodes $v$ such that $v$ belongs to some subgraph of $G$ isomorphic to $H$. However, they asked a broader and more challenging question, reiterated by Pilipczuk, Siebertz, and Toruńczyk [59]. Recall that a formula $\varphi(x)$ is local if the satisfaction of $\varphi(x)$ depends only on the $r$-neighborhood of its free variable $x$, for some fixed $r$. For instance, in graphs, the formula “node $x$ belongs to a 4-cycle” is indeed local. The question in [55, 59] goes way beyond deciding $H$-freeness. It can be formulated as follows:
Open Problem (Nešetřil and Ossona de Mendez [55]):
For every local first-order (FO) formula $\varphi(x)$, and for every graph class $\mathcal{G}$ of bounded expansion, is there a distributed algorithm that, for every $n$-node graph $G\in\mathcal{G}$, marks all vertices $v\in V(G)$ such that $G\models\varphi(v)$, in $\mathcal{O}(\log n)$ rounds in the CONGEST model?
An example of such local formula $\varphi(x)$ is, for fixed positive integers $d$ and $k$, marking all nodes $v$ such that the distance-$d$ neighborhood of $v$ in $G$ has a dominating set of size at most $k$.
1.2 Our Results
Our first main result is a positive answer to the open problem in [55, 59]. That is, we show that, for every local FO formula $\varphi(x)$, and for every graph class $\mathcal{G}$ of bounded expansion, there exists a deterministic algorithm that, for every $n$-vertex graph $G\in\mathcal{G}$, marks all vertices $v\in V(G)$ such that $G\models\varphi(v)$, in $\mathcal{O}(\log n)$ rounds in the CONGEST model. In fact, this result also holds for FO formulas on labeled graphs, i.e., graphs in which every node is provided with labels drawn from a finite set $\Lambda$. That is, the formulas may also include unary predicates on node labels. Let us call a $\Lambda$-labeled graph any graph whose nodes are label by labels in $\Lambda$. A $\Lambda$-labeled graph is thus a pair $(G,\ell)$ where $G=(V,E)$ is a graph, and $\ell:V\to 2^{\Lambda}$ is a label-assignment to the nodes of $G$. Our first main theorem is the following:
Theorem 1.1.
Let $\Lambda$ be a finite set. For every local FO formula $\varphi(x)$ on $\Lambda$-labeled graphs, and for every class of graphs $\mathcal{G}$ of bounded expansion, there exists a deterministic algorithm that, for every $n$-node network $G=(V,E)\in\mathcal{G}$, and every $\ell:V\to 2^{\Lambda}$, marks all vertices $v\in V$ such that $(G,\ell)\models\varphi(v)$, in $\mathcal{O}(\log n)$ rounds under the CONGEST model.
Before summarizing our techniques used to solve the open problem in [55, 59], it is worth mentioning that we extend this result in several directions.
1.2.1 Extension 1: Deciding FO formulas
An immediate corollary of our result on local FO formulas is that, for every formula $\varphi$ such as $\varphi=\forall x\;\psi(x)$ where $\psi(x)$ is a local FO formula (e.g., $H$-freeness), and for every graph class $\mathcal{G}$ of bounded expansion, deciding whether $G\models\varphi$ can be done in $\mathcal{O}(\log n)$ rounds in the CONGEST model (by merely applying our algorithm to $\neg\psi(x)$, and letting every marked node reject). However, Gaifman’s theorem [29] tells us that not all FO formulas are as simple as that. For instance, even the simple FO formula “there exist two vertices of degree 3” needs $\Omega(D)$ rounds to be checked in CONGEST, where $D$ denotes the diameter of the graph — this lower bound holds even for trees. In other words, a dependence on $D$ is unavoidable, even for trees. Nevertheless, our second main result states that, for every FO formula $\varphi$, and for every graph class $\mathcal{G}$ of bounded expansion, there exists a deterministic algorithm that, for every $n$-vertex graph $G\in\mathcal{G}$ of diameter $D$, decides whether $G\models\varphi$, in $\mathcal{O}(D+\log n)$ rounds. Again, our result holds for labeled graphs too, and is formally stated as follows:
Theorem 1.2.
Let $\Lambda$ be a finite set. For every FO formula $\varphi$ on $\Lambda$-labeled graphs, and for every class of graphs $\mathcal{G}$ of bounded expansion, there exists a distributed algorithm that, for every $n$-node network $G=(V,E)\in\mathcal{G}$ of diameter $D$, and every $\ell:V\to 2^{\Lambda}$, decides whether $(G,\ell)\models\varphi$ in $\mathcal{O}(D+\log n)$ rounds under the CONGEST model.
Note that bounded degeneracy would not suffice to establish the same result as above. Indeed, we show that the mere FO formula $C_{6}$-freeness requires $\tilde{\Omega}(\sqrt{n})$ rounds to be checked in CONGEST, even in graphs of degeneracy 2 and of diameter 6 (see Theorem 7.4). Moreover, our result is the best that one can hope regarding the class of formulas, in the sense that it cannot be extended to, e.g., MSO. Indeed, we show that non-3-colorability requires $\tilde{\Omega}(n)$ rounds to be checked in bounded degree graphs with logarithmic diameter (see Theorem 7.2).
1.2.2 Extension 2: Counting, Optimization, and Certification
Our results extend beyond model checking, in several directions. First, we consider the problem of counting solutions of FO formulas with free variables. We actually consider two forms of counting:
  • Local counting: Given a local formula $\varphi(x_{1},\dots,x_{k})$ with $k\geq 1$ free variables, such as $\textsf{adj}(x_{1},x_{2})\land\textsf{adj}(x_{2},x_{3})\land\textsf{adj}(x_{3},x_{1})$ (i.e., $x_{1},x_{2},x_{3}$ form a triangle), each node $v$ of the considered graph $G=(V,E)$ outputs a (possibly negative) integer $\nu(v)$ satisfying that the number of tuples $v_{1},\dots,v_{k}$ for which $(G,\ell)\models\varphi(v_{1},\dots,v_{k})$ equals $\sum_{v\in V}\nu(v)$.
  • Global counting: Given a general formula $\varphi(x_{1},\dots,x_{k})$ with $k\geq 1$ free variables, all nodes output a same value equal to the number of tuples $v_{1},\dots,v_{k}$ for which $(G,\ell)\models\varphi(v_{1},\dots,v_{k})$. E.g., counting the $k$-tuples of nodes of degree 3.
Theorem 1.3.
Let $\Lambda$ be a finite set. For every FO formula $\varphi(x_{1},\dots,x_{k})$ on $\Lambda$-labeled graphs with $k\geq 1$ free variables $x_{1},\dots,x_{k}$, and for every class of graphs $\mathcal{G}$ of bounded expansion, there exists a distributed algorithm that, for every $n$-node network $G=(V,E)\in\mathcal{G}$ of diameter $D$ and every $\ell:V\to 2^{\Lambda}$, performs global counting in ${O(D+\log n)}$ rounds in CONGEST. If the formula $\varphi(x_{1},\dots,x_{k})$ is local then there exists a distributed algorithm that performs local counting in $\mathcal{O}(\log n)$ rounds in CONGEST.
Note that, in contrast to our Theorem 1.3, the best known algorithm for (locally) counting triangles in general graphs takes $\tilde{\mathcal{O}}(n^{1/3})$ rounds [9].
Some optimization problems can also be expressed as finding optimal solutions for FO formulas with free variables, over weighted graphs, e.g., finding a maximum-weight triangle. (If no triangles exist in the graph, then all nodes reject.) Note that, as usual in the literature on CONGEST [57], the weights assigned to the nodes or edges are assumed to be polynomial in the network size so that any weight can be transmitted between neighbors in a single round of communication. We prove the following theorem.
Theorem 1.4.
Let $\Lambda$ be a finite set. For every FO formula $\varphi(x_{1},\dots,x_{k})$ on $\Lambda$-labeled graphs with $k\geq 1$ free variables $x_{1},\dots,x_{k}$, and for every class graphs $\mathcal{G}$ of bounded expansion, there exists a distributed algorithm that, for every $n$-node network $G=(V,E)\in\mathcal{G}$ of diameter $D$, every $\ell:V\to 2^{\Lambda}$, and every weight function $\omega:V\to\mathbb{N}$, computes a $k$-tuple of vertices $(v_{1},\dots,v_{k})$ such that $(G,\ell)\models\varphi(v_{1},\dots,v_{k})$ and $\sum_{i=1}^{k}\omega(v_{i})$ is maximum, in $\mathcal{O}(D+\log n)$ rounds under the CONGEST model. (The same holds if replacing maximum by minimum.) If no tuples $(v_{1},\dots,v_{k})$ exist such that $(G,\ell)\models\varphi(v_{1},\dots,v_{k})$, then all nodes reject.
Theorem 1.4 allows to find, for example, a monochromatic triangle of minimum weight, or a multi-colored $k$-independent set of maximum weight, in a class $\mathcal{G}$ of $c$-colored graphs of bounded expansion.
Last but not least, our results extend to distributed certification. Roughly, a distributed certification scheme [36, 48] for a graph property $\varphi$ is a pair prover-verifier. The prover is a centralized, computationally unlimited oracle that assigns certificates to the nodes. The verifier is a 1-round distributed algorithm whose role is to check whether the certificates assigned to the nodes form a proof that the graph satisfies the property. A distributed certification scheme is correct if it satisfies completeness (i.e., if $(G,\ell)\models\varphi$, then the prover can assign certificates such that the verifier accepts at all nodes), and soundness (i.e., if $(G,\ell)\not\models\varphi$, then, no matter the certificates assigned to the nodes, the verifier rejects in at least one node). The main complexity measure is the size of the certificates assigned by the prover on legal instances.
Theorem 1.5.
Let $\Lambda$ be a finite set. For every FO formula $\varphi$ on $\Lambda$-labeled graphs, and for every class of graphs $\mathcal{G}$ of bounded expansion, there exists a distributed certification scheme that, for every $n$-node network $G=(V,E)\in\mathcal{G}$ and every $\ell:V\rightarrow 2^{\Lambda}$, certifies $(G,\ell)\models\varphi$ using certificates on $\mathcal{O}(\log n)$ bits.
Note that, with $O(\log n)$-bit certificates, the verifier can be implemented in a single round of communication in the CONGEST model, as desired. Note also that our certification scheme is actually a proof-labeling scheme [48], which is a more demanding form of distributed certification than, e.g., locally checkable proofs [36]. As a direct consequence of Theorem 1.5, certifying $C_{4}$-freeness can be done with $\mathcal{O}(\log n)$-bit certificates in graphs of bounded expansion. In contrast, certifying $C_{4}$-freeness requires certificates on $\tilde{\Omega}(\sqrt{n})$ bits in general graphs, using the reduction from [17]. More generally, [22] shows that every MSO formula can be certified with $O(\log n)$-bit certificates in graphs of bounded treedepth. Theorem 1.5 extends this result. Indeed, while Theorem 1.5 is for FO, and the result of [22] is for MSO, both logics have identical expressive power on graphs of bounded treedepth. Last but not least, as for our model-checking algorithm, Theorem 1.5 is the best that can be achieved. Indeed, we show that it does not extend to MSO, as certifying non-3-colorability requires certificates of $\tilde{\Omega}(\sqrt{n})$ bits in planar graphs (see Theorem 7.3), and even certificates of $\tilde{\Theta}(n)$ bits in bounded degree graphs (see Theorem 7.2).
Remark.
The open problem by Nešetřil and Ossona de Mendez was actually stated for the BROADCAST CONGEST model, i.e., a weaker version of the CONGEST model in which, at every round, each node $v$ must send the same $\mathcal{O}(\log n)$-bit message to all its neighbors (see, e.g.,  [17, 47]). However, (1) the problem is not necessarily easier to solve in the general version of the CONGEST model, and (2) we solve it anyway by the affirmative using an algorithm designed for the BROADCAST CONGEST model, and the complexity of our algorithm is expressed under the latter. For the sake of simplifying the notations, we simply refer to CONGEST in the statements of our theorems, but all our upper bound results (but Theorem 1.4) hold under the BROADCAST CONGEST model too [^3] The main reason why Theorem 1.4 does not hold in BROADCAST CONGEST is that our algorithm involves an information dissemination phase going downward a tree, in which different messages are sent to different children. .
1.3 Our Techniques
In a nutshell, our technical contribution is twofold:
  • The first contribution is an adaptation of the quantifier elimination method [18, 38, 59] from centralized computing to the distributed setting. On a general level, we primarily follow the steps of the centralized approach. Yet, this adaptation is highly non-trivial, as we had to perform all operations locally without incurring a high volume of communication between adjacent nodes.
  • The second, and perhaps main, contribution of the paper is the novel locality preserving quantifier elimination method. Specifically, given a local FO formula $\varphi(x)$, we show how to eliminate its quantifiers, i.e., producing a quantifier-free formula $\widehat{\varphi}(x)$ that is equivalent to $\varphi(x)$, without losing locality. This is achieved by generalizing the notion of locality to formulas with more than a single free variable in a specific way, thereby enabling the application of induction on the number of quantifiers while preserving locality.
More specifically, our approach to establish Theorem 1.1 is mainly based on the quantifier elimination methods from [59]. Recall that, given an FO formula $\varphi$, and a class $\mathcal{G}$ of graphs with expansion $f$, quantifier elimination transforms every FO formula $\varphi$ into a quantifier-free FO formula $\widehat{\varphi}$ such that, for every $G=(V,E)\in\mathcal{G}$, and every labeling $\ell:V\to 2^{\Lambda}$, one can construct a node-labeling $\widehat{\ell}$ of $G$ such that
$(G,\ell)\models\varphi\iff(G,\widehat{\ell})\models\widehat{\varphi}.$
In full generality, quantifier elimination allows the graph $G$ to be transformed into a graph $\widehat{G}$, with $(G,\ell)\models\varphi\iff(\widehat{G},\widehat{\ell})\models\widehat{\varphi}$. However, as we shall discuss later, the approach in [59] preserves the graph $G$. Yet, we have to overcome two main difficulties for computing $\widehat{\ell}$.
First, we have to implement quantifier elimination in the distributed setting. That is, we show how to compute the new labeling $\widehat{\ell}$ of $G$, and the new formula $\widehat{\varphi}$ efficiently in the CONGEST model. Note that for an algorithm to be efficient in CONGEST, not only should the exchange of information be restricted to nearby nodes in the graph, but also the amount of information exchanged between neighbors should be small. For instance, in $\mathcal{O}(\log n)$ rounds, a node with degree $\gg\log n$ cannot directly send the list of all its neighbors’ identifiers to any of its neighbors.
Second, we have to adapt quantifier elimination for preserving locality. Given $G,\ell$, and a local FO formula $\varphi$, we show how to compute $\widehat{\ell}$ and $\widehat{\varphi}$ so that not only $\widehat{\varphi}$ is quantifier-free, and satisfies, for every $v\in V$,
$(G,\ell)\models\varphi(v)\iff(G,\widehat{\ell})\models\widehat{\varphi}(v),$
but also every stage of the quantifier elimination process preserves locality. This is achieved, in particular, by extending the natural “semantic” notion of local formula $\varphi(x)$ with a unique free variable, to a “syntactic” notion of local formula $\psi(x,\overline{y})$ involving more than one free variables, i.e., $\bar{y}=(y_{1},\dots,y_{k})$ is a vector of free variables. This syntactical notion is based on relativization, which is, in our framework, defining a new structure by restricting its domain to elements satisfying a form of mutual proximity.
Let us now enter a bit more into the details.
1.3.1 Eliminating Existential Quantifiers
In order to build the pair $(\widehat{\varphi},\widehat{\ell})$ from $(\varphi,G,\ell)$, the main issue is the elimination of the existential quantifiers (universal quantifiers can be handled as negated existential quantifiers). The technique of [59] proceeds gradually, by eliminating each existential quantifier one after the other. The basic step is therefore the elimination of one existential quantifier. Let us consider a formula $\psi(x,\bar{y})=\exists z\,\psi^{\prime}(x,\bar{y},z)$ where $\bar{y}=(y_{1},\dots,y_{k})$ is a vector of free variables — these variables are actually quantified in the original formula $\varphi(x)$, but the corresponding quantifiers will be eliminated at further steps. Let us assume that a pair $(\widehat{\psi^{\prime}},\ell^{\prime})$ has already been computed from $(\psi^{\prime},G,\ell)$, i.e., $\widehat{\psi^{\prime}}$ is quantifier-free, and $\ell^{\prime}$ is the appropriate new node-labeling of $G$. The goal is then to compute $(\widehat{\psi},\widehat{\ell})$ from $(\widehat{\psi^{\prime}},G,\ell^{\prime})$ where $\widehat{\psi}$ is quantifier-free, and $\widehat{\ell}$ is the appropriate new node-labeling. The existential quantifier in $\psi(x,\bar{y})=\exists z\,\widehat{\psi^{\prime}}(x,\bar{y},z)$ is eliminated in two stages.
First, recall that the aforementioned characterization theorem from [53] states that, for any graph class $\mathcal{G}$ of expansion $f$, there exists a function $g:\mathbb{N}\to\mathbb{N}$ satisfying that, for every $k\geq 1$, the vertices of any graph $G\in\mathcal{G}$ can be colored using a palette of $g(k)$ colors such that, for any set $U$ of at most $k$ colors, the subgraph of $G$ induced by the vertices with colors in $U$ has treedepth at most $k$. By setting $k=|\psi|$, i.e., by letting $k$ be the number of variables occurring in $\psi$, it is guaranteed that at least one set $U$ of $k$ colors captures $k$ nodes for which $\psi$ holds whenever such nodes exist. For instance, for $C_{4}$ freeness, for any coloring of the nodes with at least 4 colors, there exists a set $U$ of at most 4 colors such that, whenever there is a 4-cycle in the graph, this 4-cycle appears in the subgraph induced by the vertices colored with some color in $U$. Assuming given an appropriate $g(k)$-coloring, the quantifier elimination can proceed by considering all sets $U$ of $k$ colors, and eliminating the existential quantifier in graphs of bounded treedepth.
The second stage involves eventually eliminating the existential quantifier by transforming a formula on labeled graphs of bounded treedepth with an existential quantifier into a formula on labeled forests of bounded depth without existential quantifiers. This is achieved by encoding the existential quantifier into a disjunction of quantifier-free formulas, each referring to a specific finite subgraph pattern. To encode the subgraph pattern, new unary and binary predicates are introduced. These unary predicates are called label predicates. Their aim is to specify the presence of certain symbols in the labels assigned to the nodes. The new binary predicates are called least common ancestor predicates, or simply lca predicate for short. Their aim is to specify the depth of the least common ancestor of two nodes in certain trees of the forest. The subgraph patterns are encoded as Boolean combinations of label and lca predicates.
1.3.2 Locality-Preserving Quantifier Elimination
A first novelty of this paper is to adapt the approach of [59] such that each step preserves the locality property of the considered formula. To that end, we extend the notion of local formula to more that one free variables: Roughly, a formula $\psi(x,\bar{y})$ is local if (1) $(G,\ell)\models\varphi(x,\bar{y})$ implies that $\bar{y}\subseteq B_{G}(x,r)$, i.e., each $y_{i}$ is at distance at most $r$ from $x$, and (2) for every two graphs $G$ and $G^{\prime}$, and every two vertices $x$ and $x^{\prime}$ of $G$ and $G^{\prime}$, respectively, if there is a (label preserving) isomorphism $h:B_{G}(x,r)\to B_{G^{\prime}}(x^{\prime},r)$ satisfying $h(x)=x^{\prime}$, then $(G,\ell)\models\varphi(x,\bar{y})\iff(G^{\prime},\ell)\models\varphi(h(x),h(\bar{y}))$. This “semantic” definition is difficult to use in practice, which is why we introduce a “syntactic” notion of local form for local formulas. For this purpose, we introduce new predicates, called distance predicates. These predicates are of the form $\mathrm{ball}(y,x,r)$, interpreted as “$y$ is in the ball of radius $r$ around $x$ in the graph”. The distance predicates allow us to syntactically enforce the locality property of a formula. For instance, in a formula $\mathrm{ball}(y,x,r)\wedge\varphi(x,y)$, the distance predicate enforces the fact that $x$ and $y$ are at bounded distance. At each step of the quantifier elimination process, we make sure that both the original formula $\varphi$, and the transformed formula $\widehat{\varphi}$ explicitly contain distance predicates that guarantee their locality.
The presence of negations in the formula, potentially induced by universal quantifiers treated as negated existential quantifiers, must be treated with care. Indeed, the locality property is not stable by negation. For instance, the negation of $\mathrm{ball}(y,x,r)\wedge\varphi(x,y)$ is $\neg\mathrm{ball}(y,x,r)\vee\neg\varphi(x,y)$, which is not local because $\neg\mathrm{ball}(y,x,r)$ is satisfied for nodes $y$ that are far away from $x$. To overcome this issue, we introduce redundancy in the syntax of local formulas. That is, instead of adding distance predicates once in the formula (for instance, immediately after the innermost quantifier), we duplicate the distance predicates wherever they are needed.
For instance, let us consider a formula $\varphi(x,\bar{y})=\forall z\,\psi(x,\bar{y},z)$ that is assumed to be local. As said before, quantifier elimination proceeds by eliminating existential quantifiers, and thus this formula is rewritten as $\varphi(x,\bar{y})=\neg\exists z\,\neg\psi(x,\bar{y},z)$. Now, we make the locality of $z$ explicit by rewriting $\varphi$ as
$\varphi(x,\bar{y})=\neg\exists z\,\big{(}\mathrm{ball}(z,x,r)\land\neg\psi(x,\bar{y},z)\big{)}.$
Moreover, we use redundancy, and reinforce the locality of $\bar{y}$ by introducing another distance predicate:
$\varphi(x,\bar{y})=\mathrm{ball}(\bar{y},x,r)\land\neg\exists z\,\big{(}\mathrm{ball}(z,x,r)\land\neg\psi(x,\bar{y},z)\big{)}.$
By distributing the distance predicate on $\bar{y}$, we get
$\varphi(x,\bar{y})=\mathrm{ball}(\bar{y},x,r)\land\neg\exists z\,\big{(}\mathrm{ball}(\bar{y},x,r)\land\mathrm{ball}(z,x,r)\land\neg\psi(x,\bar{y},z)\big{)},$
which can be rewritten as
$\varphi(x,\bar{y})=\mathrm{ball}(\bar{y},x,r)\land\neg\exists z\,\big{(}\mathrm{ball}((\bar{y},z),x,r)\land\neg\psi(x,\bar{y},z)\big{)}.$
As a result, the potential loss of locality resulting from the negation of $\psi(x,\bar{y},z)$ is recovered thanks to the (redundant) distance predicate, as $\mathrm{ball}((\bar{y},z),x,r)\land\neg\psi(x,\bar{y},z)$ is local. Eliminating the existential quantifier in this formula yields a (local) quantifier-free formula. Negation is, however, applied a second time, which may result in losing locality. Yet, locality is again recovered using the other (redundant) distance predicate $\mathrm{ball}(\bar{y},x,r)$.
1.3.3 Quantifier Elimination in CONGEST
The second novelty of this paper is the implementation of quantifier elimination, including the new constraint of preserving locality, in CONGEST, specifically in BROADCAST CONGEST. We had to tackle two main difficulties.
First, the distributed implementation of the techniques in [59] requires technical attention in several places. There are indeed operations that are very simple to perform in the centralized setting (they are even often implicit when described in the centralized algorithm), but it is not necessarily clear whether all these operations are local (i.e., involving nodes that are close in the graph), and frugal in term of bandwidth (i.e., requiring neighboring nodes to exchange a limited amount of information). As a consequence, we had to carefully revisit the entire construction and ’open all boxes’ to ensure that each step of the construction can be implemented efficiently in (BROADCAST) CONGEST.
Second, model checking over graphs of bounded expansion actually refers to model checking of relational structures whose Gaifman graphs are of bounded expansion. The techniques from [59] (as well as other related approaches) involve enriching the structure relative to an arbitrary first-order formula with new symbols and predicates to obtain a quantifier-free formula. Enriching the structure naturally modifies Gaifman graphs. In contrast, in distributed computing models such as CONGEST, nodes establish communications over a network that does not change. When enriching the structure, one must therefore be very careful to avoid creating connections between nodes that cannot directly communicate in the network.
1.3.4 Model Checking, Counting, Optimization, and Certification
Model Checking.
For checking whether $(G,\ell)\models\varphi$ in CONGEST, for a given FO formula $\varphi$, and a given labeled graph $G\in\mathcal{G}$, we use the same techniques as the ones used to mark all nodes satisfying a given local FO formula. That is, the proof of Theorem 1.2 follows the same guidelines as the ones used to establish Theorem 1.1. On the one hand, the approach is simplified as one does not need to care about relativization. On the other hand, the ”non-locality” of the formula requires checking properties that potentially occur at nodes that are far away from each other in the graph (e.g., the presence of at least two nodes of degree 3). In the context of quantifier elimination, this translates into counting the number of nodes in which unary predicates are satisfied. This counting is merely achieved along the edges of a breadth-first search (BFS) tree, in $\mathcal{O}(D)$ rounds. We show that all other operations involved in quantifier elimination can be performed in $\mathcal{O}(\log n)$ rounds, for a total of $\mathcal{O}(D+\log n)$ rounds.
Counting and Optimization.
Establishing Theorem 1.3, i.e., showing how to efficiently count the number of $k$-tuples of vertices $(v_{1},\dots,v_{k})$ such that $(G,\ell)\models\varphi(v_{1},\dots,v_{k})$, requires more work. For the global counting problem, we proceed in two stages.
  • Quantifier-free formulas: As for model checking, we use the fact that there exists an integer function $g$ depending only on the class $\mathcal{G}$ such that, for any graph $G\in\mathcal{G}$, and any constant $p$, $G$ admits a coloring with $g(p)$ colors such that any subset of at most $p$ colors induces a subgraph of treedepth at most $p$. For $p=k$, each $k$-uple of vertices $(v_{1},\dots,v_{k})$ of the labeled graph in $(G,\ell)$ satisfying the formula $\varphi(x_{1},\dots,x_{k})$ covers some set $U$ of colors, with $|U|\leq p$. Assuming given an $\mathcal{O}(D+\log n)$-round algorithm counting the solutions of $\varphi(x_{1},\dots,x_{k})$ on labeled subgraphs $(G[U],\ell)$ of treedepth at most $k$, we show how to count the solutions on the whole labeled graph $(G,\ell)$. For any two sets of colors $U_{1},U_{2}$, the set of solutions of $\varphi(x_{1},\dots,x_{k})$ on $(G[U_{1}\cap U_{2}],\ell)$ is exactly the intersection of the set of solutions for $(G[U_{1}],\ell)$ and for $(G[U_{2}],L)$. This holds thanks to the fact that $\varphi(x_{1},\dots,x_{k})$ has no variables other than $(x_{1},\dots,x_{k})$. As a consequence, by counting the solutions on $(G[U],\ell)$ for every set $U$ of at most $k$ colors, and then using the inclusion-exclusion principle, we are able to count solutions of $\varphi$ on $(G,\ell)$.
  • General formulas: In order to use the previous stage on quantifier free formulas, we first proceed with quantifier elimination. This introduces new variables and new predicates, especially lca predicates, which prevents us from directly applying the previous stage. To overcome this difficulty, we eliminate the lca predicates by introducing new variables, which are existentially quantified. To avoid over-counting, we enforce unicity, i.e., at most one instance of each newly introduced variable may satisfy the formula. Informally, it is as if every existential quantifier “there exists $y$ such that” was replaced by “there exists a unique $y$ such that”.
To establish Theorem 1.4, we first utilize the entire quantifier elimination machinery developed for distributed model-checking in CONGEST. We emphasize that, during the process, the variables of the quantifier-free formula $\widehat{\varphi}$ are precisely the free variables of $\varphi(\bar{x})$, but $\widehat{\varphi}$ may have a new set $\overline{y}$ of existentially-quantified variables. Let $p=|\overline{x}|+|\overline{y}|$. We construct now a coloring of $G$ using $g(p)$ colors in $\mathcal{O}(\log n)$ rounds, and we observe that, for any $k$-uple of vertices $(v_{1},\dots,v_{k})$ of $G$, we have $(G,\ell)\models\varphi(v_{1},\dots,v_{k})$ if and only if there exists a set $U$ of at most $p$ colors such that $(G[U],\widehat{\ell})\models\widehat{\varphi}(v_{1},\dots,v_{k})$. Indeed, $U$ is obtained by guessing not only the colors of $v_{1},\dots,v_{k}$, but also the colors of the tuple of vertices $u_{1},\dots,u_{p-k}$ assigned to $\overline{y}$. Optimizing the solution of $\varphi$ on $(G,\ell)$ boils down to optimizing over all optimal solutions of $\widehat{\varphi}$ on graphs $(G[U],\widehat{\ell})$ of treedepth at most $p$.
Based on the above, it remains to deal with optimization and global counting for FO formulas on labeled subgraphs $G[U]$ of $G$, each of which has a constant treedepth. For this purpose, we extend the results from [25] stating that counting and optimization of FO formulae can be achieved on connected labeled graphs of bounded treedepth, in a constant number of rounds. Indeed, we show that one can perform optimization and counting for FO formulae over an induced (not necessarily connected) subgraph of constant treedepth of the labeled graph $G$, in $\mathcal{O}(D)$ rounds. In [25], the algorithm starts by constructing a tree decomposition of small width of the graph, and then performs dynamic programming over this tree decomposition. If the input graph of small treedepth is connected, this directly enables the construction of a rooted tree decomposition of constant depth and constant width. Moreover, the decomposition tree is a spanning tree of the graph, which allows communications to proceed with dynamic programming bottom-up the tree in a constant number of rounds. The root of the tree can then compute the output by simulating the centralized algorithm for optimization and counting on graphs of bounded treewidth of [6]. However, in our case, the graphs $G[U]$ may not be connected, and thus we must deal with a forest decomposition $F$ instead of a mere tree decomposition. A BFS tree rooted at an arbitrary vertex can span all the roots of the trees in the forest $F$, resulting in a rooted tree of depth $D+\mathcal{O}(1)$. This tree can serve as a base for constructing a tree-decomposition of $G[U]$. Moreover, since the nodes of the tree-decomposition correspond to vertices of $G$, the centralized dynamic programming algorithm of [6] can be simulated for counting and optimization on $G[U]$, in $D+\mathcal{O}(1)$ rounds. The information stored in the dynamic programming tables can indeed be computed in a sequence of rounds, starting from the leaves, upward to the root of the tree decomposition by communicating along the edges of $G$. Altogether, the whole process for counting and optimization for formula $\varphi$ on the pair $(G,\ell)$ costs $\mathcal{O}(D+\log n)$ rounds in CONGEST.
Our method for local counting follows the same lines as that for global counting. That is, we first reduce the study to quantifier-free formulas, using the (locality preserving) quantifier elimination method. Similarly to the case of global counting, this reduction introduces lca-predicates, which need to be eliminated in a locality-preserving fashion. It turns out that the method applied to the case of global counting happens to preserve locality, although the proof is somewhat technical. For quantifier-free formulas, we essentially apply the same idea as for global counting, based on the inclusion-exclusion principle. However, we also use the fact that the formula is local, by considering some set of colors $U$, and a tuple of vertices $\overline{v}$ such that $(G[U],\ell)$ satisfies $\varphi(\overline{v})$. Then, since $\varphi$ is local, all vertices in $\overline{v}$ belong to the same connected component of $G[U]$, which implies that we only need to count the solutions to $\varphi$ in the connected components of $G[U]$. The latter are connected graphs of bounded treedepth, and thus counting is achieved in $\mathcal{O}(1)$ rounds using the algorithm in [25].
Distributed Certification.
Let us start with a preliminary remark. At first glance, it may seem overcomplicated to use quantifier elimination for certifying a formula of the type $\varphi(\bar{x})=\exists y\,\psi(\bar{x},y)$. Indeed, to certify that $\varphi(\bar{x})=\top$, it may seem sufficient to certify the existence of a vertex $y$ that yield $\psi(\bar{x},y)=\top$, which can be done easily thanks to a spanning tree rooted at $y$, consuming $\mathcal{O}(\log n)$-bit certificates [48]. The situation is, however, more complex than it seems, for at least two reasons. First, in the intermediate steps $\psi$ may refers to free variables $(x_{1},\dots,x_{\ell})=\bar{x}$, which makes the identification of a vertex $y$ such that $\psi(\bar{x},y)=\top$ non-trivial. Second, there may be negations in the global formula, that is, one may have to consider also the case where $\varphi(\bar{x})=\neg\exists y\,\psi(\bar{x},y)$, and certifying the non existence of a vertex $y$ is not straightforward. A typical example is the case of certifying the existence of a 4-cycle, which is trivial using $\mathcal{O}(\log n)$-bit certificate, in contrast to certifying the non existence of a 4-cycle, which requires certificates on $\tilde{\Omega}(\sqrt{n})$ bits in general graphs.
At a high level, our certification scheme for an FO formula $\varphi$ and a class of graphs $\mathcal{G}$ with bounded expansion essentially consists of the prover providing each node with the transcript of the execution of our distributed model-checking algorithm at this node. The role of the verifier is then to check (in a distributed manner) whether this transcript encodes a correct execution of our algorithm. Note that the certification scheme does not have to certify the bounded expansion property of $\mathcal{G}$, which is a promise. However, given a coloring of the input graph $G=(V,E)\in\mathcal{G}$, the certification scheme must certify that any set $U$ of $k=|\varphi|$ colors does induce a subgraph with treedepth $k$. This is done in a specific manner, using an algorithm by Bousquet et al. [7] for certifying MSO formulas on graphs of bounded treedepth with certificates of size $\mathcal{O}(\log n)$. The main difficulty is to encode the transcript of the quantifier elimination procedure, i.e., the transformation from $(\varphi,G,\ell)$ to $(\widehat{\varphi},\widehat{\ell})$ with $\widehat{\varphi}$ quantifier-free. Each lca predicate refers to a local condition, and the value of such a predicate can be certified locally at each node, in forests of bounded depth. On the other hand, the unary predicates may refer to non-local properties. Nevertheless, we demonstrate that verifying the correctness of the unary predicates can be accomplished using a spanning tree whose root is capable of making decisions. (Certify a spanning tree is standard using $\mathcal{O}(\log n)$-bit certificates [48].)
1.4 Related Work
The results of our paper lie at the intersection of two research directions. The first is the study of distributed computing and distributed certification schemes on sparse networks, such as planar graphs, or graphs excluding a fixed minor, for fundamental graph optimization problems. The second is the development of meta-theorems for sparse graph classes in centralized computation.
1.4.1 Distributed Computing on Sparse Networks
One of the popular areas of research in distributed computing concerns the study of classical combinatorial problems on sparse networks, including planar graphs, bounded-degree graphs, graphs of bounded treewidth, and graphs excluding a fixed minor. In such settings, a central goal is to develop distributed algorithms that exploit structural sparsity to achieve improved efficiency—compared to general networks—particularly in terms of communication rounds. Typical problems studied in this context include coloring, matching, dominating sets, cuts, spanning trees, and network decompositions. Research directions focus on understanding how sparsity affects the complexity of these problems in standard distributed models such as LOCAL and CONGEST, and on designing algorithms with optimal or near-optimal round complexity. This line of work often builds on structural properties of sparse networks, revealing deep and fruitful connections between structural graph theory and distributed computation.
Examples of the work in this direction include distributed algorithms for maximal matchings, maximal independent sets (MIS), and colourings in graphs of bounded vertex degrees [56], depth-first search (DFS) for planar graphs [34], minimum spanning tree (MST) and min-cut for planar, bounded genus, and $H$-minor-free graphs [31, 32, 40, 41, 42, 33], minimum dominating set on planar and $H$-minor-free graphs [1, 4, 12, 13, 52], and MIS for graphs excluding a fixed minor [8].
The round complexity $\tilde{\mathcal{O}}(D)$ achieved in Theorem 1.2, is considered a “gold standard” for algorithms on sparse graphs in the CONGEST model. It has been obtained, for instance, for the Minimum Spanning Tree (MST) problem on planar graphs and, more generally, on $H$-minor-free graphs [31, 32, 40, 41, 42, 33]. Algorithms with $\tilde{\mathcal{O}}(D)$ complexity in the CONGEST model have also been developed for bounded-treewidth graphs for a variety of problems, including matching, the directed single-source shortest paths problem, bipartite unweighted maximum matching, and girth [46].
We are not aware of any meta-theorem that combines logical and combinatorial constraints within the CONGEST model, with one exception. The paper [25] establishes that every MSO formula on graphs of bounded treedepth is decidable in a constant number of rounds in the CONGEST model. The results of [25] are incomparable with Theorems 1.1 and 1.2. Indeed, Theorems 1.1 and 1.2 apply to graph classes that are significantly broader than graphs of bounded treedepth, whereas the number of rounds in the algorithm of [25] is independent of the number of vertices in the input graph. We incorporate some of the methods developed in [25] in our work.
Regarding distributed algorithms on graphs of bounded expansion, the foundational work in this area is due to Nešetřil and Ossona de Mendez [55]. In the CONGEST model, they presented a distributed algorithm for computing a low treedepth decomposition of graphs of bounded expansion in $O(\log n)$ rounds. We use this construction as a subroutine in our algorithms. In the same paper, Nešetřil and Ossona de Mendez also used this decomposition to design an algorithm for checking whether a graph of bounded expansion contains a given connected subgraph $H$ of constant size, also in $O(\log n)$ rounds. Theorem 1.1 significantly extends this result.
Approximation algorithms for minimum dominating set on graphs of bounded expansion in LOCAL model were developed in [1, 44, 50]. In [62], Siebertz and Vigny discuss the complexity of the model-checking problem for first-order logic on graphs of bounded expansion, in the CONGESTED CLIQUE model.
Distributed certification is a very active area of research, see [23] for a survey. Several meta-theorems are known for distributed certification. A significant result in the field was achieved in [22], who demonstrated that any MSO formula can be locally certified on graphs with bounded treedepth using a logarithmic number of bits per node, representing a gold standard in certification. This theorem has numerous implications for certification; for more details, we refer to [22]. Notably, the FO property of $C_{4}$-freeness and the MSO property of non-3-colorability, which typically require polynomial-size certificates in general graphs, can be certified with just $\mathcal{O}(\log n)$-bit certificates in graphs of bounded treedepth. The result by Bousquet et al. has been extended to more comprehensive classes of graphs, including graphs excluding a small minor [7], graphs of bounded treewidth [2, 10, 27], and graphs of bounded cliquewidth [26].
1.4.2 Centralized Meta-Theorems
Meta-theorems, according to Grohe and Kreutzer [38], are the results that certain problems of model checking for formulas of some logic are tractable (in some sense) on some class of graphs. One of the most celebrated examples of a meta-theorem is Courcelle’s theorem, which asserts that graph properties definable in MSO are decidable in linear time on graphs of bounded treewidth [11]. There is a long line of meta-theorems for FO logic on sparse graph classes in the centralized model of computation. The starting point is Seese’s theorem [61] on linear time decidability of FO on graphs of bounded degree. [28] gave linear time algorithms for planar graphs and all apex-minor-free graph classes, as well as an $\mathcal{O}(n^{1+\varepsilon})$ algorithms for graphs of bounded local tree-width. [24] proved that deciding FO properties is fixed-parameter tractable on graph classes with excluded minors. The next step to classes of graphs locally excluding a minor was done by [14].
The Sparsity Theory of [54] opened new horizons for model checking on sparse graphs. [18] proved that FO properties can be decided in linear time on graph classes of bounded expansion. Further extensions of these results are known for nowhere-dense graphs  [39] and graphs of bounded twin-width [5]. See [49] and [37] for surveys.
From a technical perspective, the most important aspect for our work is the quantifier elimination procedure introduced in [18] and [38], as well as in [59]. Similar to these works, we apply the idea of modifying the structure by adding new unary relations and functions without changing the Gaifman graph.
2 Model and Definitions
In this section, we recall basic concepts of distributed computing, logic, and graphs. Readers familiar with these frameworks may skip the first two sub-sections, but Section 2.3, which introduces and extends the notion of local formulas, and Section 2.4, which recalls properties of graphs with bounded expansion, may require special attention.
2.1 The CONGEST Model
The CONGEST model is a standard model of distributed computation in networks [57]. The network is modeled as a connected simple (i.e., no self-loops nor multiple edges) graph $G=(V,E)$. The $n\geq 1$ vertices of $G$ model the processing nodes in the networks, and its edges models the communication links. In the following, we may indistinguishably use “node” of “vertex” for referring to the processing elements in the network. Each node is assigned an identifier, which is assumed to be unique in the network. The identifiers may not be between 1 and $n$, but the CONGEST model assumes that each identifier can be stored on $\mathcal{O}(\log n)$ bits in $n$-node networks. In fact, it is usually assumed that all the nodes of a graph $G$ are given a polynomial upper bound $N=\mbox{poly}(n)$ on the number $n$ of vertices in $G$, with all identifiers in $\{1,\dots,N\}$.
Computation proceeds in lockstep, as a sequence of synchronous rounds. All nodes start simultaneously, at the same round. At each round, every node $v$ sends an $\mathcal{O}(\log n)$-bit message to each of its neighbors, receives the messages sent by all its neighbors, and performs some individual computation. The round-complexity of an algorithm $A$ is the function $R_{A}(n)=\max_{|V(G)|\leq n}R_{A}(G)$ where $R_{A}(G)$ denotes the number of rounds performed by the nodes executing $A$ in $G$ until all of them terminate.
As said before, it is worth mentioning the BROADCAST CONGEST model, that is, the weaker variant of the CONGEST model in which, at every round, each node must send the same $\mathcal{O}(\log n)$-bit message to all its neighbors. For the sake of simplifying the notations, we solely refer to the CONGEST model in the statements of our results. Yet, all our upper bound results also holds in the weaker BROADCAST CONGEST model, apart from Theorem 1.4.
Remark.
Throughout the paper, we will use the following convention regarding the inputs and outputs of our algorithms. When an algorithm receives a “global value” as input, whether it be an integer $p$, a logical formula $\varphi$, or a function (e.g., a bound $f:\mathbb{N}\to\mathbb{N}$ on the expansion), we assume that all nodes are given this global value as input. Instead, when an algorithm receives as input a node labeling function $\ell$, we assume that each node $v$ is only given its label $\ell(v)$. Similarly, when an algorithm returns a node-labeling function $\widehat{\ell}$, we mean that each node $v$ returns its label ^$\widehat{\ell}(v)$.
2.2 First-Order Formulas on Labeled Graphs
2.2.1 Definition
Recall that a first-order (FO) formula over graphs is a combination (conjunction, disjunction, and negations) of Boolean predicates over variables, which may be quantified or not. Every variable represents a vertex of a graph, and there are two binary predicates: equality $x=y$, and adjacency $\textsf{adj}(x,y)$, where the latter states whether or not there is an edge between the two vertices $x$ and $y$. The variables of an FO formula can be quantified with existential or universal quantifiers. Non-quantified variables are called free variables. Given an FO formula $\varphi$ with free variables $x_{1},\dots,x_{k}$, we make explicit that the latter are free by denoting $\varphi=\varphi(x_{1},\dots,x_{k})$. In the following, $\mathrm{FO}$ denotes the set of all first order formulas over undirected graphs.
Throughout the paper, a tuple $(x_{1},\dots,x_{k})$ is abbreviated in $\overline{x}$. The length $k$ of a tuple $(x_{1},\dots,x_{k})$ is then denoted by $|\overline{x}|$. Expressions such as $\exists\overline{x}$ or $\forall\overline{x}$ are shorthand for $\exists x_{1}\dots\exists x_{k}$ and $\forall x_{1}\dots\forall x_{k}$, respectively. We may also abuse notation in context where no confusions may occur by letting $\overline{x}$ denote the set $\{x_{1},\dots,x_{k}\}$.
Given an FO formula $\varphi$ and a graph $G$, $G\models\varphi$ denotes that $G$ satisfies $\varphi$ (i.e., $G$ is a model for $\varphi$). The negation of $G\models\varphi$ is denoted by $G\not\models\varphi$. Given an FO formula $\varphi(\overline{x})$ with free variables, a graph $G=(V,E)$, and a set of vertices $\overline{v}\in V^{k}$ where $k=|\overline{x}|$, we say that $G$ satisfies $\varphi(\overline{x})$ for vertices $\overline{v}$, denoted by $G\models\varphi(\overline{v})$, if the formula is satisfied by $G$ whenever assigning $\overline{v}$ to the free variables of $\varphi(\overline{x})$.
A graph property is called a first-order (FO) property if it can be expressed by a first-order formula with no free variables. For example, the property “the graph contains a triangle” is clearly an FO property: it holds whenever there is a set of three pairwise adjacent vertices in the graph. Indeed, it can be written as the formula
$\varphi\;=\;\exists x_{1}\,\exists x_{2}\,\exists x_{3}\,\bigl{(}\mathrm{adj}(x_{1},x_{2})\;\wedge\;\mathrm{adj}(x_{2},x_{3})\;\wedge\;\mathrm{adj}(x_{1},x_{3})\bigr{)}.$
More generally, the property of containing a fixed graph $H$ as a subgraph (sometimes called “subgraph isomorphism”) is an FO property. Its negation—namely, the property that a graph does not contain $H$ as a subgraph (often called “$H$-freeness”)—is also an FO property, since first-order logic is closed under negation.
A formula with free variables expresses properties over graphs and tuples of vertices. As an example, the following formula $\psi(x)$ holds whenever $x$ is a node in a triangle, formulated as
$\psi(x)=\exists y_{1}\exists y_{2},\textsf{adj}(x,y_{1})\wedge\textsf{adj}(x,y_{2})\wedge\textsf{adj}(y_{1},y_{2}).$
The graph $G$ with vertex set $V=\{a,b,c,d\}$, and edge set $E=\{\{a,b\},\{a,c\},\{b,c\},\{c,d\}\}$ satisfies $G\models\varphi$ as the vertices $\{a,b,c\}$ form a triangle. Similarly, $G\models\psi(a)$, but $G\not\models\psi(d)$ as $d$ is not in any triangle.
2.2.2 FO Formulas for Labeled Graphs
The formalism of FO properties can be extended to labeled graphs. Let $\Lambda$ be a finite set, whose elements are called labels. A $\Lambda$-labeled graph is a pair $(G,\ell)$ where $G=(V,E)$ is a graph, and $\ell:V\rightarrow 2^{\Lambda}$ is an assignment of a set of labels to each vertex of $G$. An FO formula over $\Lambda$-labeled graphs is a first order formula $\varphi$ over graphs, extended with predicates of the form $\textsf{lab}_{\lambda}(x)$ for every $\lambda\in\Lambda$. The predicate $\textsf{lab}_{\lambda}(x)$ is true if vertex $x$ contains $\lambda$ as one of its labels. Such predicates are called label predicates. Whenever $\textsf{lab}_{\lambda}(x)$ holds, we say that $x$ is labeled with $\lambda$. Note that since $\ell$ assigns a set of labels to each vertex, it may be the case that $\textsf{lab}_{\lambda}(x)\land\textsf{lab}_{\lambda^{\prime}}(x)$ holds for two different labels $\lambda$ and $\lambda^{\prime}$. It may thus be convenient to define the predicates $\textsf{lab}_{S}$ for sets $S\subseteq\Lambda$, as follows. Given $S\in\Lambda$, $\textsf{lab}_{S}(x)$ holds if and only if $\ell(x)=S$. In particular, $\textsf{lab}_{S}(x)\land\textsf{lab}_{S^{\prime}}(x)$ is necessarily false whenever $S\neq S^{\prime}$. The set of all first-order logic formulas over $\Lambda$-labeled graphs is denoted $\mathrm{FO}[\Lambda]$.
Given a $\Lambda$-labeled graph $(G,\ell)$, and a subgraph $G^{\prime}$ of $G$, we abuse notation and denote $(G^{\prime},\ell)$ the $\Lambda$-labeled graph defined by the restriction of $\ell$ to the nodes in $G^{\prime}$. For instance, let $\Lambda=\{0,1\}$, and let us consider the following graph property on labeled graphs: “existence of a triangle with a node labeled $1$”. This property can be expressed by the formula:
$\varphi=\exists x_{1}\exists x_{2}\exists x_{3},\textsf{adj}(x_{1},x_{2})\wedge\textsf{adj}(x_{2},x_{3})\wedge\textsf{adj}(x_{1},x_{3})\wedge\textsf{lab}_{1}(x_{1}).$
We may also consider formula with free variables, such as
$\psi(x)=\exists y_{1}\exists y_{2},\textsf{adj}(x,y_{1})\wedge\textsf{adj}(x,y_{2})\wedge\textsf{adj}(y_{1},y_{2})\wedge\textsf{lab}_{1}(x).$
The graph $G$ with vertex set $V=\{a,b,c,d\}$ and edge set $E=\{\{a,b\},\{a,c\},\{b,c\},\{c,d\}\}$ with labels $\ell(a)=\{0\},\ell(b)=\{0\},\ell(c)=\{1\}$, and $\ell(d)=\{1\}$ satisfies $(G,\ell)\models\varphi$ as the nodes $\{a,b,c\}$ form a triangle where $c$ is labeled $1$. For the same reason, we have that $(G,\ell)\models\psi(c)$. However, $(G,\ell)\not\models\psi(a)$ as $a$ is not labeled $1$ (even if it is in a triangle), and $(G,\ell)\not\models\psi(d)$ as node $d$ is not in a triangle (even if it is labeled $1$).
For $\varphi(x_{1},\dots,x_{k})\in\mathrm{FO}[\Lambda]$, and a $\Lambda$-labeled graph $(G,\ell)$, we denote by $\textsf{true}(\varphi,G,\ell)$ the set of all tuples of vertices $(v_{1},\dots,v_{k})$ satisfying that $(G,\ell)\models\varphi(v_{1},\dots,v_{k})$. If $\varphi$ has no free variables then $\textsf{true}(\varphi,G,\ell)$ is equal to the truth value of $(G,\ell)\models\varphi$. Two $\mathrm{FO}[\Lambda]$ formulas $\varphi_{1}$ and $\varphi_{2}$ are equivalent if $\textsf{true}(\varphi_{1},G,\ell)=\textsf{true}(\varphi_{2},G,\ell)$ for every $\Lambda$-labeled graph $(G,\ell)$. The following basic definitions play a central role in this paper.
  • A formula $\varphi(\overline{x})\in\mathrm{FO}[\Lambda]$ is called quantifier-free if it has no quantifiers, i.e., it can be expressed as a boolean combination of predicates involving only the free variables in $\overline{x}$.
  • A formula $\varphi(\overline{x})\in\mathrm{FO}[\Lambda]$ is existential if it can be written as $\varphi(\overline{x})=\exists\overline{y}\,\psi(\overline{x},\overline{y})$, where $\psi(\overline{x},\overline{y})$ is quantifier-free formula in $\mathrm{FO}[\Lambda]$.
Furthermore, a formula $\varphi(\overline{x})$ is in prenex normal form if it can be written in the form
$\varphi(\overline{x})=Q_{1}y_{1}\,\dots\,Q_{t}y_{t}\;\psi(\overline{x},\overline{y}),$
where $\psi(\overline{x},\overline{y})\in\mathrm{FO}[\Lambda]$, and $Q_{1},\dots,Q_{t}$ are existential or universal quantifiers. The value of $t$ is the quantifier depth of $\varphi$. Every $\mathrm{FO}$ formula admits an equivalent formula in prenex normal form (see, e.g., [21]). This fact can be easily established by induction on the length of the formula, and it can be extended to $\mathrm{FO}[\Lambda]$, i.e., first order formulas for labeled graphs, in a straightforward manner.
Notation.
For every family $\mathcal{G}$ of graphs, and every finite set $\Lambda$, we denote by $\mathcal{G}[\Lambda]$ the set of all $\Lambda$-labeled graphs $(G,\ell)$ with $G=(V,E)\in\mathcal{G}$, and $\ell:V\to 2^{\Lambda}$.
2.3 Local Formulas
Local formulas can be defined as follows. For a graph $G=(V,E)$ and a vertex $v\in V$, let $B_{G}(v,r)$ denote the $r$-neighborhood of $v$ in $G$, that is, the subgraph of $G$ induced by vertices at distance at most $r$ from $v$. By extension, when considering a labeled graph $(G,\ell)$, $B_{G}(v,r)$ is also a labeled graph: each node $u\in B_{G}(v,r)$ is labeled by the same set of labels $\ell(u)$.
Definition 2.1.
Let $\Lambda$ be a finite set of labels, and let $r\geq 0$ be an integer. A formula $\varphi(x)\in FO[\Lambda]$ with one free variable is $r$-local if, for any two $\Lambda$-labeled graphs $(G,\ell)$ and $(G^{\prime},\ell^{\prime})$, and for any two vertices $v\in V(G)$ and $v^{\prime}\in V(G^{\prime})$, we have
$B_{G}(v,r)=B_{G^{\prime}}(v^{\prime},r)\implies\Bigl{(}(G,\ell)\models\varphi(v)\iff(G^{\prime},\ell^{\prime})\models\varphi(v^{\prime})\Bigr{)}.$
where $B_{G}(v,r)=B_{G^{\prime}}(v^{\prime},r)$ means that there exists a label-preserving isomorphism $h:B_{G}(v,r)\to B_{G^{\prime}}(v^{\prime},r)$ such that $h(v)=v^{\prime}$.
We say $\varphi(x)$ is local if there exists $r\geq 0$ such that $\varphi(x)$ is $r$-local.
For technical reasons that will appear clear later, we need to extend the notion of local formulas to formulas $\varphi(x,\overline{y})$ with more than a single free variable. Our notion of locality is defined with respect to the first free variable $x$ of the formula: intuitively, we require that each variable is in a ball centered on $x$. This is different from the notion of Gaifman-locality [29], in which quantified variables are required to be in a ball centered on any of the free variables.
Definition 2.2.
Let $\Lambda$ be a finite set of labels, and let $r\geq 0$ be an integer. A formula $\varphi(x,\overline{y})\in FO[\Lambda]$ is $r$-local if, for any $\Lambda$-labeled graph $(G,\ell)$ with $G=(V,E)$, the following two properties are satisfied:
  • for every pair $(v,\overline{u})\in\textsf{true}(\varphi,G,\ell)$, it holds that $\overline{u}\subseteq B_{G}(v,r)$;
  • for every $\Lambda$-labeled graph $(G^{\prime},\ell^{\prime})$ with $G^{\prime}=(V^{\prime},E^{\prime})$, for every vertices $v\in V,v^{\prime}\in V^{\prime}$, $\overline{u}\subseteq B_{G}(v,r)$ and $\overline{u^{\prime}}\subseteq B_{G^{\prime}}(v^{\prime},r)$, if there exists a label-preserving isomorphism $h:B_{G}(v,r)\to B_{G^{\prime}}(v^{\prime},r)$ such that $h(v)=h(v^{\prime})$ and $h(\overline{u})$ = $h(\overline{u^{\prime}})$, then:
    $(G,\ell)\models\varphi(v,\overline{u})\iff(G^{\prime},\ell^{\prime})\models\varphi(v^{\prime},\overline{u^{\prime}})$
We shall often write $\varphi(\overline{x})$ instead of $\varphi(x,\overline{y})$, in which case it is $x_{1}$ that takes the particular role of $x$ in the definition above.
The above definition is, unfortunately, very difficult in practice, because it provides little information on the form of the considered formula. We therefore define the notion of local form in Definition 2.3. To that end, let us introduce the new predicate $\mathrm{ball}(y,x,r)$ defined as
$G\models\mathrm{ball}(u,v,r)\iff u\in B_{G}(v,r).$
More generally, for $\overline{y}=(y_{1},\dots,y_{k})$, we define
$\mathrm{ball}(\overline{y},x,r)=\bigwedge_{i=1}^{|\overline{y}|}\mathrm{ball}(y_{i},x,r).$
These predicates will be referred to as distance predicates. Note that they can easily be written as existential FO formulas, see Lemma 2.5. We use distance predicates to relativize quantifiers with respect to the $r$-neighborhood of a given vertex (relativization of quantifiers is standard technique, see for example [45, 19]): let us then consider an existential $r$-local formula $\varphi(x)=\exists\overline{y}\,\psi(x,\overline{y})$. Since $\varphi(x)$ is $r$-local, it can be rewritten as
$\varphi(x)=\exists\overline{y}\;\big{(}\psi(x,\overline{y})\wedge\mathrm{ball}(\overline{y},x,r)\big{)}.$
Such a formula, where the locality is enforced explicitly by a distance predicate, is referred to as a formula in $r$-local form. The following definition formalizes this notion.
Definition 2.3.
Let $r\geq 0$ be an integer. A formula $\varphi(x,\overline{y})\in FO[\Lambda]$ with at least one free variable is in $r$-local form if
  • $\varphi(x,\overline{y})=\psi(x,\overline{y})\wedge\mathrm{ball}(\overline{y},x,r)$, where $\psi\in FO[\Lambda]$ is quantifier-free, or
  • $\varphi(x,\overline{y})=\exists z\;\psi(x,\overline{y},z)$, where $\psi\in FO[\Lambda]$ is in $r$-local form, or
  • $\varphi(x,\overline{y})=\mathrm{ball}(\overline{y},x,r)\wedge\forall z\;\big{(}\neg\mathrm{ball}((\overline{y},z),x,r)\vee\psi(x,\overline{y},z)\big{)}$ where $\psi\in FO[\Lambda]$ is in $r$-local form.
We say $\varphi(x,\overline{y})$ is in local form if there exists $r\geq 0$ such that $\varphi(x,\overline{y})$ is in $r$-local form.
When considering a formula with several free variables, we say it is ($r$-)local if it can be rewritten in ($r$-)local form.
Remark.
The last case of Definition 2.3 introduces a lot of redundancy, with two distance predicates, in addition to requiring that $\psi$ must be in $r$-local form. The main motivation for this redundancy is to guarantee that every “subformula” of $\varphi$ remains local. Thanks to this redundancy, all the formulas below are local, despite the fact that negation does not preserve locality for formulas with several free variables. In particular, the following two formulas are equivalent, where the universal quantifier is merely replaced by a negated existential quantifier:
$\mathrm{ball}(\overline{y},x,r)$ $\wedge\forall z\;\big{(}\neg\mathrm{ball}((\overline{y},z),x,r)\vee\psi(x,\overline{y},z)\big{)}$
$\equiv\mathrm{ball}(\overline{y},x,r)\wedge\neg\exists z\;\big{(}\mathrm{ball}((\overline{y},z),x,r)\wedge\neg\psi(x,\overline{y},z)\big{)}.$
Similarly, the formulas
$\exists z\;\big{(}\mathrm{ball}((\overline{y},z),x,r)\wedge\neg\psi(x,\overline{y},z)\big{)},\;\text{and}\;\mathrm{ball}((\overline{y},z),x,r)\wedge\neg\psi(x,\overline{y},z)$
are both local. In other words, the implicit or explicit presence of negations forces us to introduce redundancies in the distance predicates for “balancing” the potential loss of the locality property.
The following result establishes the consistency of Definitions 2.2 and 2.3.
Lemma 2.4.
For every integer $r\geq 0$, every $r$-local formula can be rewritten in $r$-local form.
Proof.
Let $\varphi(x,\overline{y})$ be $r$-local formula, and assume, w.l.o.g, that it is written in prenex normal form $\varphi(x,\overline{y})=Q_{1}z_{1}\,\dots\,Q_{t}z_{t}\;\psi(\overline{x},\overline{y},\overline{z})$. First, we rewrite $\varphi$ by relativizing the quantifier with respect to the $r$-neighborhood of $x$. Intuitively, we replace existential quantification $\exists z$ by $\exists z\in B_{G}(x,r)$, i.e., $z$ must be in the ball of radius $r$ around $x$ in $G$. Similarly universal quantification $\forall z$ is replaced by $\forall z\in B_{G}(x,r)$, i.e., the formula is checked only for nodes $z$ close to $x$. Formally, for the $r$-local formula $\varphi(x,\overline{y})$, we define $\varphi^{r}(x,\overline{y})$ by induction on quantifier depth. If $\varphi$ is quantifier free, then $\varphi^{r}=\varphi$. If $\varphi(x,\overline{y})=\exists z,\zeta(x,\overline{y},z)$, then $\varphi^{r}(x,\overline{y})=\exists z,\mathrm{ball}(z,x,r)\wedge\zeta^{r}(x,\overline{y},z)$. Finally, if $\varphi(x,\overline{y})=\forall z,\zeta(x,\overline{y},z)$, then $\varphi^{r}(x,\overline{y})=\neg\exists z,\mathrm{ball}(z,x,r)\wedge\neg\zeta^{r}(x,\overline{y},z)$.
By definition, for $v\in V$ and $\overline{u}\subseteq B_{G}(v,r)$,
$G\models\varphi^{r}(v,\overline{u})\iff B_{G}(v,r)\models\varphi^{r}(v,\overline{u})\iff B_{G}(v,r)\models\varphi(v,\overline{u}).$
By locality of $\varphi$, we deduce that
$G\models\mathrm{ball}(\overline{u},v,r)\wedge\varphi^{r}(v,\overline{u})\iff B_{G}(v,r)\models\varphi(v,\overline{u})\iff G\models\varphi(v,\overline{u}).$
It only remains to show that the formula $\mathrm{ball}(\overline{y},x,r)\wedge\varphi^{r}(x,\overline{y})$ can be rewritten in $r$-local form. Once again, we proceed by induction on quantifier depth. By definition of $\varphi^{r}$, there are only three cases to consider.
Case 1:
$\varphi$ is quantifier free. Then $\varphi^{r}=\varphi$ and $\mathrm{ball}(\overline{y},x,r)\wedge\varphi(x,\overline{y})$ is already in local form.
Case 2:
$\varphi(x,\overline{y})=\exists z\,\psi(x,\overline{y},z)$. Then,
$\mathrm{ball}(\overline{y},x,r)\wedge\varphi^{r}(x,\overline{y})$ $=\mathrm{ball}(\overline{y},x,r)\wedge\exists z\,\big{(}\mathrm{ball}(z,x,r)\wedge\psi^{r}(x,\overline{y},z)\big{)}$
$=\exists z\,\big{(}\mathrm{ball}((\overline{y},z),x,r)\wedge\psi^{r}(x,\overline{y},z)\big{)}$
By induction, we can rewrite $\mathrm{ball}((\overline{y},z),x,r)\wedge\psi^{r}(x,\overline{y},z)$ in $r$-local form.
Case 3:
$\varphi(x,\overline{y})=\forall z,\psi(x,\overline{y},z)$. Then,
$\mathrm{ball}(\overline{y},x,r)\wedge\varphi^{r}(x,\overline{y})$ $=\mathrm{ball}(\overline{y},x,r)\wedge\neg\exists z\,\big{(}\mathrm{ball}(z,x,r)\wedge\neg\psi^{r}(x,\overline{y},z)\big{)}$
$=\mathrm{ball}(\overline{y},x,r)\wedge\neg\Big{(}\mathrm{ball}(\overline{y},x,r)\wedge\exists z\,\big{(}\mathrm{ball}(z,x,r)\wedge\neg\psi^{r}(x,\overline{y},z)\big{)}\Big{)}$
$=\mathrm{ball}(\overline{y},x,r)\wedge\neg\exists z\,\big{(}\mathrm{ball}((\overline{y},z),x,r)\wedge\neg\psi^{r}(x,\overline{y},z)\big{)}$
$=\mathrm{ball}(\overline{y},x,r)\wedge\forall z\,\big{(}\neg\mathrm{ball}((\overline{y},z),x,r)\vee\psi^{r}(x,\overline{y},z)\big{)}$
$=\mathrm{ball}(\overline{y},x,r)\wedge\forall z\,\Big{(}\neg\mathrm{ball}((\overline{y},z),x,r)\vee\big{(}\mathrm{ball}((\overline{y},z),x,r)\wedge\psi^{r}(x,\overline{y},z)\big{)}\Big{)}$
By induction, $\mathrm{ball}((\overline{y},z),x,r)\wedge\psi^{r}(x,\overline{y},z)$ can be rewritten in $r$-local form.
The study of these three cases completes the proof. ∎
We conclude this section by some simple observation that will be useful at several places in the paper.
Lemma 2.5.
Let $r\in\mathbb{N}$, and let $\overline{x}=(x_{1},\dots,x_{k})$. There exists an existential FO formula
$\beta_{r}(\overline{x})=\exists\overline{y}\,\psi(\overline{x},\overline{y}),$
where $\psi$ is quantifier free such that, for every graph $G$, and every $k$-tuple $\bar{v}$ of vertices of $G$,
$G\models\beta_{r}(\overline{v})\iff G\models\mathrm{ball}(\overline{v},v_{1},r).$
Moreover, for every $r^{\prime}\geq r$, every graph $G$, and every $k$-tuple $\bar{v}$ of vertices of $G$,
$G\models\beta_{r}(\overline{v})\iff G\models\beta_{r,r^{\prime}}(\overline{v}),$
where $\beta_{r,r^{\prime}}(\overline{x})=\exists\overline{y}\,\big{(}\psi(\overline{x},\overline{y})\wedge\mathrm{ball}((\overline{x},\overline{y}),x_{1},r^{\prime})\big{)}$.
Proof.
It is sufficient to describe $\beta_{r}(x,y)$ with two variables only, as $\beta_{r}(x_{1},\dots,x_{k})$ can be written as $\bigwedge_{j=2}^{k}\beta_{r}(x_{1},x_{j})$. We define
$\beta_{r}(x,y)=\exists z_{0}\,\dots\,\exists z_{r}\,\left[(x=z_{0})\wedge\bigvee_{j=0}^{r}\Big{(}(y=z_{j})\wedge\bigwedge_{i=0}^{j-1}\textsf{adj}(z_{i},z_{i+1})\Big{)}\right].$
In this formula, $j$ is the distance between $x$ and $y$, and $z_{0},\dots,z_{j}$ represents the path from $x$ to $y$. This formula satisfies the first item of Lemma 2.5. Let us now prove that it satisfies the second item. Let $r^{\prime}\geq r$. By definition, if $G\models\beta_{r,r^{\prime}}(u,v)$ then $G\models\beta_{r}(u,v)$. Conversely, let us assume that $G\models\beta_{r}(u,v)$. Then, there exists some $\overline{w}=w_{0},\dots,w_{r}$ such that $G\models\psi(u,v,\overline{w})$. Moreover, there exists $j\in[0,r]$ such that
$G\models(u=w_{0})\wedge(v=w_{j})\wedge\bigwedge_{i=0}^{j-1}\textsf{adj}(w_{i},w_{i+1}).$
For every $i\in[0,j]$, let $w^{\prime}_{i}=w_{i}$, and, for every $i\in[j+1,r]$, let $w^{\prime}_{i}=u$. By definition, every vertex $w^{\prime}_{i}$ satisfies $\mathrm{ball}(w^{\prime}_{i},u,r)$, and thus $\mathrm{ball}(w^{\prime}_{i},u,r^{\prime})$ since $r^{\prime}\geq r$. This implies that
$G\models(u=w^{\prime}_{0})\wedge(v=w^{\prime}_{j})\wedge\bigwedge_{i=0}^{j-1}\textsf{adj}(w^{\prime}_{i},w^{\prime}_{i+1})\wedge\mathrm{ball}((u,v,\overline{w^{\prime}}),u,r^{\prime}).$
As a consequence, $G\models\psi(u,v,\overline{w^{\prime}})\wedge\mathrm{ball}((u,v,\overline{w^{\prime}}),u,r^{\prime})$, and thus $G\models\beta_{r,r^{\prime}}(u,v)$. ∎
Lemma 2.6.
Let $\Lambda$ be a set of labels, and let $\varphi(\overline{x})\in\mathrm{FO}[\Lambda]$ be a $r$-local, quantifier-free formula. If $\varphi(\overline{x})=\varphi_{1}(\overline{x})\vee\varphi_{2}(\overline{x})$ for some formulas $\varphi_{1}$ and $\varphi_{2}$, then both $\varphi_{1}$ and $\varphi_{2}$ must be $r$-local.
Proof.
Since $\varphi_{1}$ and $\varphi_{2}$ play symmetric roles, we only prove that $\varphi_{1}$ is local. Since $\varphi$ is quantifier free, and therefore $\varphi_{1}$ as well, checking the $r$-locality of $\varphi_{1}$ is equivalent to checking that, for all $\Lambda$-labeled graphs $(G,\ell)$,
$(G,\ell)\models\varphi_{1}(\overline{v})\implies\overline{v}\subseteq B_{G}(v_{1},r).$
This holds trivially as, by locality of $\varphi$,
$(G,\ell)\models\varphi_{1}(\overline{v})\implies(G,\ell)\models\varphi(\overline{v})\implies\overline{v}\subseteq B_{G}(v_{1},r),$
which completes the proof. ∎
2.4 Graphs of Bounded Expansion
Recall that a graph $G=(V,E)$ has treedepth at most $d$ if there exists a forest $F$ of rooted trees with depth at most $d$ satisfying that there exists a one-to-one mapping $f:V\to V(F)$ such that, for every edge $\{u,v\}\in E$, $f(u)$ is an ancestor of $f(v)$ in a tree of $F$, or $f(v)$ is an ancestor of $f(u)$ in a tree of $F$. The treedepth of a graph is the smallest $d$ such that such a mapping $f$ exists.
It has been recently shown in [25] that decision, counting and optimization of FO formulas can be performed in $\mathcal{O}(1)$ rounds in CONGEST, for connected labeled graphs of bounded treedepth.
Proposition 2.7 ([25]).
Let $\mathcal{G}[\Lambda]$ be a class of connected labeled graphs of bounded treedepth, and let $\varphi(x_{1},\dots,x_{k})\in FO[\Lambda]$. Then, for any and any $(G,\ell)\in\mathcal{G}[\Lambda]$, the following problems can be solved by a distributed algorithm running in $\mathcal{O}(1)$ rounds in CONGEST:
- Deciding $\varphi$:
deciding whether it exists a $k$-uple of vertices $(v_{1},\dots,v_{k})$ such that$(G,\ell)\models\varphi(v_{1},\dots,v_{k})$.
- Counting the solutions of $\varphi$:
computing the number of $k$-uples $(v_{1},\dots,v_{k})$ such that$(G,\ell)\models\varphi(v_{1},\dots,v_{k})$.
- Optimizing $\varphi$:
[…]
Graphs of bounded expansion [54] can be defined in term of local minors, and degeneracy, as pointed out in the Introduction. It is however more convenient for us to use the following characterization. A class $\mathcal{G}$ of graphs has bounded expansion if there exists a function $f:\mathbb{N}\to\mathbb{N}$ such that, for every $G\in\mathcal{G}$, and for every $k\in\mathbb{N}$, there exists a coloring of the vertices of $G$ with colors in $[f(k)]=\{1,\dots,f(k)\}$ such that, for every set $U\subseteq[f(k)]$ of cardinality at most $k$ the subgraph of $G$ induced by the nodes colored by a color in $U$ has treedepth at most $|U|$. In particular, each subset of at most $k$ colors induces a subgraph of treedepth at most $k$. For every positive integers $p$ and $k$, a $(p,k)$-treedepth coloring of a graph $G$ is a $k$-coloring of $G$ satisfying that each set $U\subseteq[k]$ of at most $p$ colors induces a graph of treedepth at most $|U|$.
Definition 2.8.
Given a function $f:\mathbb{N}\rightarrow\mathbb{N}$, a class of graphs $\mathcal{G}$ has expansion $f$ if, for every $p\in\mathbb{N}_{>0}$, every $G\in\mathcal{G}$ has a $(p,f(p))$-treedepth coloring. A class of graphs has bounded expansion if it has expansion $f$ for some function $f:\mathbb{N}\rightarrow\mathbb{N}$.
A class of graphs has effective bounded expansion if it has expansion $f$ for some Turing-computable function $f$. In the following, we always assume that the considered functions $f$ are Turing-computable. Interestingly, treedepth colorings can be efficiently computed in the CONGEST model.
Proposition 2.9 ([55]).
Let $\mathcal{G}$ be a class of graphs of bounded expansion $f$. There exists a function $g:\mathbb{N}\to\mathbb{N}$ such that, for any graph $G\in\mathcal{G}$, and any constant positive integer $p$, a $(p,g(p))$-treedepth coloring of $G$ can be computed in $\mathcal{O}(\log n)$ rounds in the CONGEST model.
The algorithm of [55] takes as input the function $f$, and an integer $p$, and outputs a $(p,g(p))$-treedepth coloring of the actual graph $G$ in a distributed manner. The function $g$ may however be different from the given bound $f$ on the expansion of the class $\mathcal{G}$, which may itself be different from the “best” expansion function for the class $\mathcal{G}$.
3 Distributed Model Checking of Local Formulas
We have now all the ingredients to solve the open problem in [55], as established in Theorem 1.1 restated below.
See 1.1
Our algorithm runs on labeled graphs $(G,\ell)\in\mathcal{G}[\Lambda]$. Initially, every node $v\in V$ of an $n$-node labeled graph $(G,\ell)\in\mathcal{G}[\Lambda]$ is only given its identifier $\textsf{id}(v)$ on $\mathcal{O}(\log n)$ bits, which is unique in the graph, and its labels $\ell(u)\in 2^{\Lambda}$ (in addition to a polynomial upper bound on the number $n$ of vertices, as specified in the CONGEST model). At the end of the execution of our algorithm, every node $v$ outputs $\top$ if $(G,\ell)\models\varphi(v)$, and $\bot$ otherwise.
We prove Theorem 1.1 by adapting the (centralized) quantifier elimination technique of [59] (see also [18, 38]) to the distributed setting. Let $(G,\ell)$ be a labeled graph with $G\in\mathcal{G}$ and $\ell:V\to 2^{\Lambda}$, on which we aim at checking a local formula $\varphi(x)$ at every node. Quantifier elimination proceeds by induction on the depth of the formula. The formula considered at any stage of the induction may thus have several free variables $\overline{x}$ with $x_{1}=x$. We therefore describe a distributed algorithm that, in $\mathcal{O}(\log n)$ rounds, transforms both the formula $\varphi(\overline{x})$ and the node-labeling $\ell$ into a new formula $\widehat{\varphi}(\overline{x})$ and a new node labeling $\widehat{\ell}$ such that:
  • $\widehat{\varphi}(\overline{x})$ is quantifier-free, but uses new unary and binary predicates,
  • $\widehat{\ell}$ uses a new set of labels $\widehat{\Lambda}\supseteq\Lambda$ for the nodes of $G$, using additional labels involved in the new unary and binary predicates, and
  • the transformation preserves locality, and, for any tuple $\overline{v}$ of $|\overline{x}|$ vertices of $G$, we have:
    $(G,\ell)\models\varphi(\overline{v})\iff(G,\widehat{\ell})\models\widehat{\varphi}(\overline{v}).$
Note that, while the transformation preserves locality, it might be the case that $\varphi$ is $r$-local for some $r$, where as $\widehat{\varphi}$ is $r^{\prime}$-local for some $r^{\prime}\geq r$. Quantifiers are eliminated one by one, and the main difficulty consists in the elimination of the existential quantifiers. As in [59], the proof works in three steps.
  • First, the transformation from $(\varphi,\ell)$ to $(\widehat{\varphi},\widehat{\ell})$ is performed on rooted forests of constant depth (cf. Subsection 3.1). This is the most technical part of the quantifier elimination process.
  • Second, the quantifier elimination is performed on graphs of bounded treedepth (cf. Subsection 3.2), by reducing the analysis of graphs of bounded treedepth to the analysis of rooted forests. This is where the set of labels is enriched.
  • Third, the analysis of graphs of bounded expansion is reduced to the analysis of graphs of bounded treedepth thanks to a $(p,f(p))$-treedepth coloring (cf. Subsection 3.3), for an appropriate $p$. This is the steps for which the preservation of the locality property needs particular care.
We emphasis that, at each step of the transformation of the original formula $\varphi$, and the original labeled graph $(G,\ell)$, the computations must be efficiently performed in the CONGEST model. Typically, one needs to make sure that adding new labels to the graph can be done efficiently in CONGEST. The rooted forests involved in Subsection 3.1, as well as the graphs of bounded treedepth involved in Subsection 3.2, are subgraphs of the communication graph $G$, and thus all communications passing through the edges of the former can be implemented along the edges of the latter.
3.1 Rooted Forests of Bounded Depth
In this section we adapt the quantifier elimination procedure from [59] for rooted forests of constant depth (i.e., every tree in the forest is rooted and has constant depth) to the case of local formulas, and show how to implement this procedure in CONGEST. The basic idea for eliminating the quantifiers is to encode information about the satisfiability of the formula in node-labels. This approach does increase the size of the label set. Nevertheless, by taking advantage of the simple structure of FO formulas on forests of constant depth, it is possible to bound the number of labels by a constant. This constant depends solely on the size of the formula, and on the depth of the trees.
As the locality property focuses on distances between nodes, the diameter of each tree plays an important role. Therefore, instead of considering forests with depth at most $d$ for some $d$, we consider forests in which each tree has diameter at most $d-1$ for some $d$. For every $d\in\mathbb{N}$, let us denote by $\mathcal{F}_{d}$ the set of forests in which each tree has diameter at most $d-1$. Note that, in particular, all forests of $\mathcal{F}_{d}$ have depth at most $d$. For a set of labels $\Lambda$, let us denote by $\mathcal{F}_{d}[\Lambda]$ the set of all $\Lambda$-labeled graphs $(G,\ell)$ such that $G\in\mathcal{F}_{d}$.
3.1.1 First-Order Formulas on Forests of Bounded Depth
Let $\Lambda$ be a (finite) set of labels, and let $d$ be a positive integer. It is convenient to consider a specific way of encoding formulas, suitable for $\mathcal{F}_{d}[\Lambda]$. For each $i\in\{-1,0,1,\dots,d-1\}$, we define the predicate $\textsf{lca}_{i}(x,y)$ (for least common ancestor) as the predicate with free variables $x$ and $y$ that holds if the path $P_{x}$ from $x$ to the root of the tree containing $x$, and the path $P_{y}$ from $y$ to the root of the tree containing $y$ share exactly $i+1$ nodes. Moreover, we set $\textsf{lca}_{-1}(x,y)$ to be true whenever $x$ and $y$ are in different trees. Note that $\textsf{lca}_{i}(x,x)$ holds if and only if $x$ is a node at depth $i$ in its tree. In particular, $\textsf{lca}_{0}(x,x)$ holds if and only if $x$ is a root. The predicates $\textsf{lca}_{i}$ are called lca-predicates. As opposed to the label predicates, for which it may be the case that $\textsf{lab}_{\lambda}(x)$ and $\textsf{lab}_{\lambda^{\prime}}(x)$ both hold for two different labels $\lambda\neq\lambda^{\prime}$, we have that, for every $x$ and $y$, $\textsf{lca}_{i}(x,y)$ holds for exactly one index $i\in\{-1,0,1,\dots,d-1\}$.
A formula $\varphi\in FO[\Lambda]$ is said to be lca-expressed if it contains solely predicates of the form $\textsf{lca}_{i}(x,y)$ (lca predicates), and $\textsf{lab}_{\lambda}(x)$ (label predicates). In other words, in a lca-expressed formula, there are no predicates of type $\textsf{adj}(x,y)$ or $x=y$. The following is a direct consequence of the definitions.
Lemma 3.1.
Every formula $\varphi\in\mathrm{FO}[\Lambda]$ to be evaluated in $\mathcal{F}_{d}[\Lambda]$ can be lca-expressed by replacing every equality predicate $x=y$ by
$\bigvee_{i\in[0,d-1]}\Big{(}\textsf{lca}_{i}(x,x)\land\textsf{lca}_{i}(x,y)\land\textsf{lca}_{i}(y,y)\Big{)},$
and every adjacency predicate $\textsf{adj}(x,y)$ by
$\bigvee_{i\in[0,d-2]}\Big{(}\textsf{lca}_{i}(x,y)\land\big{(}(\textsf{lca}_{i}(x,x)\land\textsf{lca}_{i+1}(y,y))\lor(\textsf{lca}_{i+1}(x,x)\land\textsf{lca}_{i}(y,y))\big{)}\Big{)}.$
Given a fixed set $\Lambda$ of labels, a formula $\varphi\in\mathrm{FO}[\Lambda]$ is called lca-reduced if (1) it is lca-expressed, and (2) it is quantifier-free. In other words, an lca-reduced formula is a Boolean combination of label and lca predicates on free variables.
Definition 3.2.
Let $\overline{x}=(x_{1},\dots,x_{k})$ be a $k$-tuple of variables. For any two functions $\gamma:[k]\rightarrow 2^{\Lambda}$, and $\delta:[k]\times[k]\rightarrow[-1,d-1]$, let us define the formula
$\textsf{type}_{\gamma,\delta}(\overline{x})\in\mathrm{FO}[\Lambda]$
as the lca-reduced formula defined by the conjunction of the following two predicates:
[…]
and
[…]
We call such predicate lca-type, and we denote by $\textsf{Type}(k,d,\Lambda)$ the set of all lca-type predicates with $k$ variables.
Observe that the size of $\textsf{Type}(k,d,\Lambda)$ is at most $2^{k|\Lambda|}(d+1)^{k^{2}}$, and that it can be computed from $k,d$ and $\Lambda$. Note also that, given $\gamma$ and $\delta$, it may be the case that the formula $\textsf{type}_{\gamma,\delta}(\overline{x})$ is always false. This is typically the case when the function $\delta$ yields a conjunction of incompatible lca predicates (e.g., $\delta(i,j)=5$, $\delta(i,i)=2$, and $\delta(j,j)=1$, or merely any function $\delta$ such that $\delta(i,j)\neq\delta(j,i)$). In such cases, i.e., when $\textsf{type}_{\gamma,\delta}(\overline{x})$ is always false, we say that the formula is trivial.
For any non-trivial lca-type formulas $\varphi(\overline{x})=\textsf{type}_{\gamma,\delta}(\overline{x})$, we have that $\delta$ and $\gamma$ induce a labeled forest $(F_{\varphi},\ell_{\varphi})\in\mathcal{F}_{d}[\Lambda]$, as follows. Each $i\in[k]$ is identified with a node $u_{i}$ in $F_{\varphi}$, and each leaf of $F_{\varphi}$ belongs to the $k$-tuple $\overline{u}_{\varphi}=(u_{1},\dots,u_{k})$. Furthermore, for every $(i,j)\in[k]\times[k]$, $\textsf{lca}_{\delta(i,j)}(u_{i},u_{j})$ holds in $F_{\varphi}$. Finally, for every $i\in[k]$, $\ell_{\varphi}(u_{i})=\gamma(i)$, and $\ell_{\varphi}(v)=\emptyset$ whenever $v\notin\overline{u}_{\varphi}$.
For any labeled forest $(F,\ell)$ with $F=(V,E)$, any integer $k$, and any set of vertices $\overline{v}\in V^{k}$, let us denote by $F[\overline{v}]$ the subgraph of $F$ induced by all the nodes in $\overline{v}$, and their ancestors.
Lemma 3.3.
Let $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$ with $F=(V,E)$, and let $\varphi(\overline{x})=\textsf{type}_{\gamma,\delta}(\overline{x})$ be a non-trivial lca-type formula with $k$ free variables. For every $\overline{v}=(v_{1},\dots,v_{k})\in V^{k}$, $(F,\ell)\models\varphi(\overline{v})$ if and only if there exists a graph isomorphism $f$ between $F[\overline{v}]$ and $F_{\varphi}$ such that, for every $i\in[k]$, $\ell_{\varphi}(f(v_{i}))=\ell(v_{i})$.
Proof.
Recall that $\overline{u}_{\varphi}=(u_{1},\dots,u_{k})$ includes the leaves of $F_{\varphi}$. Let $\overline{v}=(v_{1},\dots,v_{k})$ be a set of $k$ vertices of $F$. Suppose first that $(F,\ell)\models\varphi(\overline{v})$. Then, for every $(i,j)\in[k]\times[k]$,
$\gamma(i)=\ell(v_{i}),\text{ and }\textsf{lca}(v_{i},v_{j})=\delta(i,j).$
Let $f$ be the function that, for every $i\in[k]$, $f(v_{i})=u_{i}$, and, for every $r\in\{1,\dots,\delta(i,i)-1\}$, $f$ maps the ancestor of $v_{i}$ in $F$ at distance $r$ from $v_{i}$ to the corresponding ancestor of $u_{i}$ in $F_{\varphi}$. By construction $f$ is an isomorphism between $F[\overline{v}]$ and $F_{\varphi}$. Moreover, for every $i\in[k]$, we do have $\ell_{\varphi}(f(v_{i}))=\ell_{\varphi}(u_{i})=\gamma(i)=\ell(v_{i})$, as claimed.
Conversely, let us assume that there exists a graph isomorphism $f$ between $F[\overline{v}]$ and $F_{\varphi}$ such that, for every $i\in[k]$, $\ell_{\varphi}(f(v_{i}))=\ell(v_{i})$. Then we directly infer that, for every $(i,j)\in[k]\times[k]$, $\gamma(i)=\ell(v_{i})$ and $\textsf{lca}(v_{i},v_{j})=\delta(i,j)$. This implies that $(F,\ell)\models\varphi(\overline{v})$, as claimed. ∎
The following statement establishes that every lca-reduced formula can be expressed as a disjunction of lca-types.
Lemma 3.4.
For every lca-reduced formula $\varphi(\overline{x})\in\mathrm{FO}[\Lambda]$, there exists a set $I\subseteq\textsf{Type}(|\overline{x}|,d,\Lambda)$, and a formula
$\varphi^{\prime}(\overline{x})=\bigvee_{\psi\in I}\psi(\overline{x})$
such that, for every $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$, $\textsf{true}(\varphi,F,\ell)=\textsf{true}(\varphi^{\prime},F,\ell)$. Moreover, $\varphi^{\prime}(\overline{x})$ can be computed given only $d$, $\Lambda$, and $\varphi$. The formula $\varphi^{\prime}(\overline{x})$ is called the basic normal form of $\varphi$.
Proof.
Let $\varphi(\overline{x})$ be an lca-reduced formula with $k$ free variables, and let $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$. Let $\overline{v}=(v_{1},\dots,v_{k})$ be a tuple of nodes such that $(F,\ell)\models\varphi(\overline{v})$. Let us define the function
$\delta_{\overline{v}}:[k]\times[k]\rightarrow[-1,d-1]$
by $\delta_{\overline{v}}(i,j)=h$ if the path from $v_{i}$ to the root of the tree of $F[\overline{v}]$ containing $v_{i}$, and the path from $v_{j}$ to the root of the tree of $F[\overline{v}]$ containing $v_{j}$ share exactly $h+1$ nodes. We also define the function $\gamma_{\overline{v}}$ by
$\gamma_{\overline{v}}(i)=\ell(v_{i})$
for every $i\in[k]$. We obtain that
$(F,\ell)\models\textsf{type}_{\gamma_{\overline{v}},\delta_{\overline{v}}}(\overline{v}).$
Now let $I\subseteq\textsf{Type}(k,d,\Lambda)$ be the set of all possible lca-types $\textsf{type}_{\gamma_{\overline{v}},\delta_{\overline{v}}}$ that can be defined from a labeled forest $(F,\ell)\in\mathcal{F}_{d}$, and a tuple of vertives $\overline{v}\in\textsf{true}(\varphi,F,\ell)$. By construction, the function
$\varphi^{\prime}(\overline{x})=\bigvee_{\psi\in I}\psi(\overline{x})$
satisfies $\textsf{true}(\varphi,F,\ell)=\textsf{true}(\varphi^{\prime},F,\ell)$ for every $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$. ∎
Example.
Let us consider the following formula using labels in $\{0,1\}$, hence the two predicates $\textsf{lab}_{0}$ and $\textsf{lab}_{1}$:
$\varphi(x_{1},x_{2},x_{3})=\textsf{adj}(x_{1},x_{2})\wedge\textsf{adj}(x_{1},x_{3})\wedge\textsf{lab}_{1}(x_{1})\wedge\textsf{lab}_{0}(x_{2}))\wedge\textsf{lab}_{0}(x_{3}),$
which holds when $x_{1}$ is adjacent to $x_{2}$, $x_{1}$ is labeled $1$, and $x_{2}$ and $x_{3}$ are both labeled 00. This formula can be rewritten as the lca-reduced formula
$\tilde{\varphi}(x_{1},x_{2},x_{3})=\left(\bigvee_{i\in[0,d-2]}\xi^{1}_{i}(x_{1},x_{2},x_{3})\right)\vee\left(\bigvee_{i\in[0,d-3]}\xi_{i}^{2}(x_{1},x_{2},x_{3})\right)\vee\left(\bigvee_{i\in[0,d-3]}\xi_{i}^{3}(x_{1},x_{2},x_{3})\right)$
where
$\xi^{1}_{i}(x_{1},x_{2},x_{3})=\bigg{(}$ $\textsf{lca}_{i}(x_{1},x_{1})\wedge\textsf{lca}_{i+1}(x_{2},x_{2})\wedge\textsf{lca}_{i+1}(x_{3},x_{3})\wedge$
$\textsf{lca}_{i}(x_{1},x_{2})\wedge\textsf{lca}_{i}(x_{1},x_{3})\wedge\textsf{lca}_{i}(x_{2},x_{3})\wedge$
$\textsf{lab}_{1}(x_{1})\wedge\textsf{lab}_{0}(x_{2})\wedge\textsf{lab}_{0}(x_{3})\bigg{)},$
$\xi^{2}_{i}(x_{1},x_{2},x_{3})=\bigg{(}$ $\textsf{lca}_{i+1}(x_{1},x_{1})\wedge\textsf{lca}_{i}(x_{2},x_{2})\wedge\textsf{lca}_{i+2}(x_{3},x_{3})\wedge$
$\textsf{lca}_{i}(x_{1},x_{2})\wedge\textsf{lca}_{i+1}(x_{1},x_{3})\wedge\textsf{lca}_{i}(x_{2},x_{3})\wedge$
$\textsf{lab}_{1}(x_{1})\wedge\textsf{lab}_{0}(x_{2})\wedge\textsf{lab}_{0}(x_{3})\bigg{)},$
and
$\xi^{3}_{i}(x_{1},x_{2},x_{3})=\bigg{(}$ $\textsf{lca}_{i+1}(x_{1},x_{1})\wedge\textsf{lca}_{i+2}(x_{2},x_{2})\wedge\textsf{lca}_{i}(x_{3},x_{3})\wedge$
$\textsf{lca}_{i+1}(x_{1},x_{2})\wedge\textsf{lca}_{i}(x_{1},x_{3})\wedge\textsf{lca}_{i}(x_{2},x_{3})\wedge$
$\textsf{lab}_{1}(x_{1})\wedge\textsf{lab}_{0}(x_{2})\wedge\textsf{lab}_{0}(x_{3})\bigg{)}.$
Observe that $\tilde{\varphi}$ is the basic normal form of $\varphi$. Indeed, for every $i\in[0,d-2]$, $\xi_{i}^{1}$ is an lca-type formula, and, for every $i\in[0,d-3]$, $\xi_{i}^{2}$ and $\xi_{i}^{3}$ are lca-type formulas. For instance, $\xi_{i}^{1}=\textsf{type}_{\gamma,\delta}$ with
$\gamma(1)=\{1\},~{}\gamma(2)=\{0\},~{}\gamma(3)=\{0\},$
and with
$\delta(1,1)=i,~{}\delta(2,2)=i+1,~{}\delta(3,3)=i+1,~{}\delta(1,2)=i,~{}\delta(1,3)=i\text{ and }\delta(2,3)=i.$
Figure 2 depicts $F_{\xi^{1}_{i}}$, $F_{\xi^{2}_{i}}$, and $F_{\xi^{3}_{i}}$. If a labeled forest $(F,\ell)$ with label set $\{0,1\}$ satisfies $\varphi$, then $F$ contains a forest isomorphic to $F_{\xi^{1}_{i}}$, $F_{\xi^{2}_{i}}$ or $F_{\xi^{3}_{i}}$, for some $i\in[0,d-2]$.
$x_{1}$$x_{2}$$x_{3}$$i$$x_{2}$$x_{1}$$x_{3}$$i$$x_{3}$$x_{1}$$x_{2}$$i$
Figure 2: $F_{\xi^{1}_{i}}$ (left), $F_{\xi_{i}^{2}}$ (middle), and $F_{\xi^{3}_{i}}$ (right).
3.1.2 Quantifier Elimination of lca-Expressed Formulas
In this section, we show that it is possible to eliminate existential quantifiers on lca-expressed formulas, by augmenting the set of labels.
Definition 3.5.
Let $\Lambda$ be a finite set of labels, and let $d$ be a positive integer. A formula $\varphi(\overline{x})\in\mathrm{FO}[\Lambda]$ is said reducible on $\mathcal{F}_{d}[\Lambda]$ if there exists a set $\widehat{\Lambda}$, and an lca-reduced formula $\widehat{\varphi}(\overline{x})\in\mathrm{FO}[\widehat{\Lambda}]$ such that (1) $\widehat{\Lambda}$ and $\widehat{\varphi}$ depend only on $\Lambda$, $\varphi$, and $d$, and (2) for every $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$, there exists a $\widehat{\Lambda}$-labeling $\widehat{\ell}$ of $F$ such that
$\textsf{true}(\varphi,F,\ell)=\textsf{true}(\widehat{\varphi},F,\widehat{\ell}).$
Each element in the triplet $(\widehat{\Lambda},\widehat{\varphi},\widehat{\ell})$ of Definition 3.5 is called the reduction on $\mathcal{F}_{d}[\Lambda]$ of $\Lambda$, $\varphi$, and $\ell$, respectively. We define a similar notion for local formulas.
Definition 3.6.
Let $\Lambda$ be a finite set of labels, and let $d$ be a positive integer. A formula $\varphi(\overline{x})\in FO[\Lambda]$ is said $r$-local reducible on $\mathcal{F}_{d}[\Lambda]$ if there exists a set $\widehat{\Lambda}$, and an lca-reduced formula $\widehat{\varphi}(\overline{x})\in\mathrm{FO}[\widehat{\Lambda}]$ such that (1) $\widehat{\Lambda}$ and $\widehat{\varphi}$ depend only on $\Lambda$, $\varphi$, and $d$, and (2) for every $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$, there exists a $\widehat{\Lambda}$-labeling $\widehat{\ell}$ of $F$ such that
$\textsf{true}(\varphi,F,\ell)=\textsf{true}(\widehat{\varphi},F,\widehat{\ell})=\textsf{true}(\widehat{\varphi}\wedge\mathrm{ball}(\overline{x},x_{1},r),F,\widehat{\ell}).$
In Definition 3.6, one insists on the fact that the reduction $\widehat{\varphi}$ of $\varphi$ must be $r$-local. This is called a $r$-local reduction.
Remark.
The fact that both Definitions 3.5 and 3.6 require $\widehat{\Lambda}$ and $\widehat{\varphi}$ to depend only on $\Lambda$, $\varphi$ and $d$ is not crucial for deciding (local) formulas, but will be used for the design of our certification scheme later in the paper (the prover gives $\widehat{\Lambda}$ and $\widehat{\varphi}$ to the nodes as constant-size certificates).
The following lemma is at the core of quantifier elimination.
Lemma 3.7.
Every formula $\varphi(\overline{x})=\exists\overline{y}\;\zeta(\overline{x},\overline{y})$ in $\mathrm{FO}[\Lambda]$, where $\zeta$ is an lca-reduced formula, is reducible on $\mathcal{F}_{d}[\Lambda]$.
Proof.
Without loss of generality, let us assume that $|x|=k\geq 1$. (In the case where $\varphi$ has no free-variables, we can create a formula $\varphi(z)=\varphi$ for a dummy variable $z$.) We only deal with the case where $|y|=1$, i.e., with the elimination of a single quantified variable. Indeed, for the case where $|y|>1$ we can iteratively eliminate the quantifiers one by one, until we obtain a quantifier-free (actually an lca-reduced) formula. Using Lemma 3.1 we can assume without loss of generality that $\zeta$ is lca-reduced. Moreover, thanks to Lemma 3.4, we can also assume that $\zeta(\overline{x},y)$ is expressed in the basic normal form. That is, there exist $I\subseteq\textsf{Type}(k+1,d,\Lambda)$ such that $\zeta(\overline{x},y)=\bigvee_{\psi\in I}\psi(\overline{x},y)$. Since the existential quantifier commutes with the disjunction, we get that
$\varphi(\overline{x})=\bigvee_{\psi\in I}\exists y\;\psi(\overline{x},y).$
Let us fix an arbitrary labeled graph $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$. Let us first assume that each formula $\psi\in I$ is reducible on $\mathcal{F}_{d}[\Lambda]$ (we show how to do this reduction later in the proof). Under this assumption, let us denote by $\widehat{\Lambda}^{\psi}$, $\widehat{\psi}$, and ^$\widehat{\ell}^{\psi}$ the resulting reductions. We can relabel every set $\Lambda^{\psi}$, for $\psi\in I$, so that the sets $\widehat{\Lambda}^{\psi}\smallsetminus\Lambda$ are pairwise disjoint. Let us then define:
$\widehat{\Lambda}=\bigcup_{\psi\in I}\widehat{\Lambda}^{\psi},\;\widehat{\varphi}=\bigvee_{\psi\in I}\widehat{\psi},\text{ and, for every node $u\in F$, }\;\widehat{\ell}(u)=\bigcup_{\psi\in I}\widehat{\ell}^{\psi}(u).$
We get that $\widehat{\Lambda}$, $\widehat{\varphi}$, and $\widehat{\ell}$ are reductions on $\mathcal{F}_{d}[\Lambda]$ of $\Lambda$, $\varphi$, and $\ell$, respectively. Therefore, it just remains to prove that every $\psi\in\textsf{Type}(k+1,d,\Lambda)$ is reducible.
Let $\gamma$ and $\delta$ be such that $\psi(\overline{x},y)=\textsf{type}_{\gamma,\delta}(\overline{x},y)\in\textsf{Type}(k+1,d,\Lambda)$, where $y$ is identified with variable numbered $k+1$ in the definitions of $\gamma$ and $\delta$. Let $s\in[k]$ such that $\delta(s,k+1)$ is maximum (this value exists because $\varphi$ has at least one free variable). We define
$h=\delta(s,k+1),\;h_{s}=\delta(s,s),\text{ and }h_{y}=\delta(k+1,k+1).$
In other words, $x_{s}$ is at depth $h_{s}$, $y$ is of depth $h_{y}$, and the least common ancestor of $x_{s}$ and $y$ (if exists) is at depth $h$ in $F_{\psi}$. We consider three cases, depending on the value of $\delta(s,k+1)$, and on the relative values of $h_{y}$ and $h$.
Case 1: $\delta(s,k+1)\geq 0$ and $h_{y}=h$.
In this case $y$ is an ancestor of $x_{s}$ in $F_{\psi}$. Let $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$. A node $w$ at depth $h$ is called a candidate if $\ell(w)=\gamma(k+1)$. A node at depth $h_{s}$ having an ancestor that is a candidate is called a good node. Let $\widehat{\Lambda}^{\psi}=\Lambda\cup\{\textsf{good}\}$. We define a $\widehat{\Lambda}^{\psi}$-labeling ^$\widehat{\ell}^{\psi}$ of $F$ as follows:
^$\widehat{\ell}^{\psi}(u)=\left\{\begin{array}[]{ll}\ell(u)\cup\{\textsf{good}\}&\mbox{if $u$ is a good node, }\\ \ell(u)&\mbox{otherwise.}\end{array}\right.$
Let us now construct an lca-reduced formula $\widehat{\psi}$ that is equivalent to $\psi$. Let us denote by $\widehat{\gamma}:[k]\rightarrow 2^{\Lambda}$ and $\widehat{\delta}:[k]\times[k]\rightarrow[-1,d-1]$ the restrictions of $\gamma$ and $\delta$ to $[k]$ (i.e., for each $i,j\in[k]$, $\widehat{\gamma}(i)=\gamma(i)$ and $\widehat{\delta}(i,j)=\delta(i,j)$). Let $\xi=\textsf{type}_{\widehat{\gamma},\widehat{\delta}}$. Since $h_{y}=h$, we have that $F_{\xi}=F_{\psi}$, and $\ell_{\xi}(\overline{u}_{\xi})=\ell_{\psi}(\overline{u}_{\psi})$. Now, given a tuple of vertices $\overline{v}=(v_{1},\dots,v_{k})$, we have that $(F,\ell)\models\exists y,\psi(\overline{v},y)$ if and only if the following conditions hold:
  • $F[\overline{v}]$ is isomorphic to $F_{\psi}$ (which equals $F_{\xi}$),
  • for every $i\in[k]$, $\ell(v_{i})=\gamma(i)=\widehat{\gamma}(i)$, and
  • the ancestor of $v_{s}$ at level $h$ in $F[\overline{v}]$ is labeled $\gamma(k+1)$, i.e., $v_{s}$ is a good node.
From Lemma 3.3 we obtain that the first two conditions imply that $(F,\ell)\models\psi(\overline{v})$. The third condition can be expressed as $\textsf{good}\in\ell(v_{1})$. Therefore, $\widehat{\psi}$ is defined as the conjunction of the two previous expressions, that is,
$\widehat{\psi}(\overline{x})=\textsf{type}_{\widehat{\gamma},\widehat{\delta}}(\overline{x})\land\textsf{lab}_{\textsf{good}}(x_{s}).$
Therefore, $\textsf{true}(\psi,F,\ell)=\textsf{true}(\widehat{\psi},F,\widehat{\ell}^{\psi})$ for every $\Lambda$-labeled forest $(F,\ell)$ of depth at most $d$.
Case 2: $\delta(s,k+1)\geq 0$ and $h_{y}>h$.
Let $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$. In this case, a node $v$ at depth $h_{y}$ is called a candidate if $\ell(v)=\gamma(k+1)$. A node $v$ at depth greater than $h$ is called a good node if at least one of its descendants is a candidate node. Finally, a node $v$ at depth $h$ is called a pivot, and we denote by $\kappa(v)$ the number of children of $v$ that are good nodes.
We define $\widehat{\gamma}$, $\widehat{\delta}$, and $\xi=\textsf{type}_{\widehat{\gamma},\widehat{\delta}}$ the same way than in Case 1. Under the assumptions of Case 2, $F_{\xi}$ is however not necessarily equal to $F_{\psi}$. In fact $F_{\xi}$ is a sub-forest of $F_{\psi}$. Indeed, $F_{\psi}$ can be obtained from $F_{\xi}$ by adding a path that connects a pivot node in $F_{\psi}$ to a candidate node not in $F_{\psi}$.
Given a tuple $\overline{v}$ of $k$ vertices, checking whether $F[\overline{v}]$ is isomorphic to $F_{\xi}$ can be achieved by merely using formula $\xi$. Let $z$ be the pivot in $F[\overline{v}]$, that is, $z$ is the ancestor of $v_{s}$ at level $h$. Let $\kappa(z,\overline{v})$ be the set of good children of $z$ contained in $F[\overline{v}]$. To eliminate the existential quantifier in $\varphi(\overline{x})$, it is sufficient to check whether one can extend $F[\overline{v}]$ by plugging a path from $z$ to a candidate node not contained in $F[\overline{v}]$. This condition holds if $\kappa(z,\overline{v})<\kappa(z)$, and it is exactly what is checked when considering the formula $\widehat{\psi}$.
Let us define $\widehat{\Lambda}^{\psi}=\Lambda\cup[0,k+1]\cup\{\textsf{good}\}$, and let us consider the $\widehat{\Lambda}^{\psi}$-labeling ^$\widehat{\ell}^{\psi}$ of $F$ defined as follows. Initially, for every node $u$ of $F$, ^$\widehat{\ell}^{\psi}(u)=\ell(u)$. We then update the label ^$\widehat{\ell}^{\psi}(u)$ of every node $u$ as follows.
  • If $u$ is a node at depth $h_{s}$, then let us denote $v$ the ancestor of $u$ at depth $h$. The label $\lambda$ is added to ^$\widehat{\ell}^{\psi}(u)$ where
    $\lambda=\begin{cases}\kappa(v)&\text{ if }\kappa(v)\leq k\\ k+1&\text{otherwise}.\end{cases}$
  • If $u$ is a good node, the label good is added to ^$\widehat{\ell}^{\psi}(u)$.
Observe that the two conditions above are not necessarily exclusive, and it may be the case that both labels $\lambda$ and good are added to the set of labels of a the same node $u$.
Let us suppose that there exists an lca-reduced $\widehat{\Lambda}^{\psi}$-formula $\alpha(\overline{x})$ that checks whether $\kappa(z)>\kappa(z,\overline{x})$ where $z$ is the ancestor at level $h$ of $x_{s}$. We can then define $\widehat{\psi}$ as the lca-reduced $\widehat{\Lambda}^{\psi}$-formula
$\widehat{\psi}(\overline{x})=\textsf{type}_{\widehat{\gamma},\widehat{\delta}}(\overline{x})~{}\wedge~{}\alpha(\overline{x}),$
where
$\alpha(\overline{x})=\bigvee_{q\in[p]}\bigg{[}\bigg{(}\bigvee_{q\leq m\leq k+1}\textsf{lab}_{m}(x_{s})\bigg{)}\wedge\bigvee_{S\in K(p,q-1)}\bigg{(}\bigwedge_{i\in S}\bigvee_{j\in X_{i}}\textsf{lab}_{\textsf{good}}(x_{j})\wedge\bigwedge_{i\notin S}\bigwedge_{j\in X_{i}}\neg\textsf{lab}_{\textsf{good}}(x_{j})\bigg{)}\bigg{]}.$
We claim that $\textsf{true}(\psi,F,\ell)=\textsf{true}(\widehat{\psi},F,\widehat{\ell}^{\psi})$ for every $\Lambda$-labeled forest $(F,\ell)\in\mathcal{F}_{d}$. Indeed, observe first that $\kappa(z,\overline{x})\leq k$ as the forest $F_{\xi}$ has degree at most $k$ (because it has at most $k$ leaves). Therefore, $\alpha(\overline{x})$ holds whenever $\kappa(z)>k$, which is equivalent to the predicate $\textsf{lab}_{k+1}(x_{s})$. If $\kappa(z)\leq k$ then the value of $\kappa(z)$ is stored in the label of $x_{s}$. Let us denote by $z_{1},\dots,z_{p}$ the $p\leq k$ children of $z$ in $F_{\psi}$. For every $i\in[p]$, let us denote by $X_{i}$ the set of variables in $\overline{x}$ that are descendants of $z_{i}$. For every $q\in[p]$, we also denote by $K(p,q)$ the set of all subsets of $[p]$ of at most $q$ elements. If $q=\kappa(z)$ and $t=\kappa(z,\overline{x})$, then checking $q>t$ can be done by checking the existence of set $S\in K(p,q-1)$ satisfying that, for every $i\in S$, $X_{i}$ contains a good node, and, for every $i\notin S$, $X_{i}$ does not contain a good node.
Case 3: $\delta(s,k+1)=-1$.
In this case, the node $y$ in $F_{\psi}$ has no common ancestors with any of the nodes $x_{1},\dots,x_{k}$. In other words, $y$ is in a tree of $F_{\psi}$ that is different from any of the trees containing $x_{1},\dots,x_{k}$. Let $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$. Like in previous cases, a node $u$ at depth $h_{y}$ is called a candidate if $\ell(u)=\gamma(k+1)$. A root $r$ of $F$ is said to be active if some descendant of $r$ is a candidate node. A node is called good if it is a descendant of an active root.
Let $\widehat{\gamma}$, $\widehat{\delta}$ and $\xi=\textsf{type}_{\widehat{\gamma},\widehat{\delta}}$ be defined as in Case 1. In Case 3, $F_{\xi}$ is a sub-forest of $F_{\psi}$, where $F_{\psi}$ can be obtained from $F_{\xi}$ by adding a path from an active root $r$ not contained in $F_{\xi}$ to a candidate node.
Let $\overline{v}=(v_{1},\dots,v_{k})$ be a set of vertices of $F$, and let $\rho(F)$ be the number of active roots in $F$. Similarly, let $\rho(F,\overline{v})$ be the number of active roots in $F[\overline{v}]$. It follows that $(F,\ell)\models\exists y,\psi(\overline{v},y)$ if and only if $F[\overline{v}]$ is isomorphic to $F_{\xi}$, $\ell(v_{i})=\gamma(i)$ for each $i\in[k]$, and $F$ contains an active root not contained in $F[\overline{v}]$. The first two conditions can be checked using $\xi$, and the latter is equivalent to $\rho(F)>\rho(F,\overline{v})$.
Let $\widehat{\Lambda}=\Lambda\cup[0,k+1]\cup\{\textsf{good}\}$ be a set of labels, and let us consider the $\widehat{\Lambda}^{\psi}$-labeling ^$\widehat{\ell}^{\psi}$ of $F$ defined as follows. Initially, for every node $u$ in $F$, let us set ^$\widehat{\ell}^{\psi}(u)=\ell(u)$. We then update the label ^$\widehat{\ell}^{\psi}(u)$ of every node $u$ as follows.
  • For every node $u$, the label $\lambda$ is added to ^$\widehat{\ell}^{\psi}(u)$, where
    $\lambda=\begin{cases}\rho(F)&\text{ if }\rho(F)\leq k\\ k+1&\text{ if }\rho(F)>k.\end{cases}$
  • If $u$ is a good node, then the label good is added ^$\widehat{\ell}^{\psi}(u)$.
Let us suppose that there exists an lca-reduced $\widehat{\Lambda}^{\psi}$-formula $\alpha(\overline{x})$ that checks whether $\rho(F)>\rho(F,\overline{x})$ where $z$ is the ancestor at level $h$ of $x_{1}$. We can then define $\widehat{\psi}$ as the lca-reduced $\widehat{\Lambda}^{\psi}$-formula
$\widehat{\psi}(\overline{x})=\textsf{type}_{\widehat{\gamma},\widehat{\delta}}(\overline{x})~{}\wedge~{}\xi(\overline{x}),$
where
$\xi(\overline{x})=\bigvee_{q\in[p]}\bigg{[}\bigg{(}\bigvee_{q\leq m\leq k+1}\textsf{lab}_{m}(x_{1})\bigg{)}\wedge\bigvee_{S\in K(p,q-1)}\bigg{(}\bigwedge_{i\in S}\bigvee_{j\in X_{i}}\textsf{lab}_{\textsf{good}}(x_{j})\wedge\bigwedge_{i\notin S}\bigwedge_{j\in X_{i}}\neg\textsf{lab}_{\textsf{good}}(x_{j})\bigg{)}\bigg{]}$ (1)
We claim that $\textsf{true}(\psi,F,\ell)=\textsf{true}(\widehat{\psi},F,\widehat{\ell}^{\psi})$ for every $\Lambda$-labeled forest $(F,\ell)\in\mathcal{F}_{d}$. Indeed, observe that $F_{\xi}$ can contain at most $k$ roots. Therefore $\alpha$ is true if $\rho(F)>k$, which is equivalent the predicate $\textsf{lab}_{k+1}(x_{1})$. If $\rho(F)\leq k$ then the value of $\rho(F)$ is stored in the label of $x_{1}$. Let $r_{1},\dots,r_{p}$ denote the $p\leq k$ roots of $F_{\xi}$. For every $i\in[p]$, let $X_{i}$ be the set of variables in $\overline{x}$ that are descendants of $r_{i}$. We obtain that checking whether $q=\rho(F)>\rho(F,\overline{x})$ is equivalent to check the existence of a set $S\in K(p,q-1)$ such that, for every $i\in[p]$, $X_{i}$ contains a good node if and only if $i\in S$, from which the expression of $\xi$ follows. ∎
Finally, the following lemma is a refined version of Lemma 3.7 enforcing locality.
Lemma 3.8.
Let $d\in\mathbb{N}$, $r\geq d$, and $\varphi(\overline{x})=\exists\overline{y}\,\big{(}\zeta(\overline{x},\overline{y})\wedge\mathrm{ball}(\{\overline{x},\overline{y}\},x_{1},r)\big{)}$ be an $r$-local lca-expressed existential formula in $\mathrm{FO}[\Lambda]$, i.e., $\zeta$ is a lca-reduced formula in $\mathrm{FO}[\Lambda]$. It holds that $\varphi$ is $d$-local reducible on $\mathcal{F}_{d}[\Lambda]$.
Proof.
Since all trees in the forests $F\in\mathcal{F}_{d}$ have diameter at most $d$, and since $r\geq d$, we get that $\mathrm{ball}(z,x_{1},r)$ holds if and only if $x_{1}$ and $z$ are in the same tree, which is equivalent to $\neg\textsf{lca}_{-1}(x_{1},z)$. We define:
$\varphi_{1}(\overline{x})=\exists\overline{y},\zeta(\overline{x},\overline{y})\wedge\bigwedge_{z\in\overline{x},\overline{y}}\neg\textsf{lca}_{-1}(x_{1},z)$
We have $\textsf{true}(\varphi,F,\ell)=\textsf{true}(\varphi_{1},F,\ell)$. We can now apply Lemma 3.7 on $\varphi_{1}$. (Note that there were three cases the proof of Lemma 3.7, but the third case cannot happen here as this case corresponds to the case where two variables are not in the same connected components of $F$, which is forbidden by the terms of the form $\neg\textsf{lca}_{-1}(x_{1},z)$.) By Lemma 3.7, we obtain an lca-reduced formula $\widehat{\varphi}$ and a labeling $\widehat{\ell}$ such that $\textsf{true}(\varphi_{1},F,\ell)=\textsf{true}(\widehat{\varphi},F,\widehat{\ell})$. Now, as mentioned above, each $x_{i}$ is in the same component as $x_{1}$. Since each tree has diameter at most $d$, this mere fact implies that $\mathrm{ball}(x_{i},x_{1},d)$ holds for every $i$. As a consequence, $\textsf{true}(\widehat{\varphi},F,\widehat{\ell})=\textsf{true}(\widehat{\varphi}\wedge\mathrm{ball}(\overline{x},x_{1},d),F,\widehat{\ell})$. ∎
3.1.3 Quantifier Elimination of lca-Reduced Formulas in CONGEST
We now show that the quantifier elimination on local formulas described in the proof of Lemma 3.8 can be computed in $\mathcal{O}(1)$ rounds in the CONGEST model.
Lemma 3.9.
Let $\Lambda$ be a finite set of labels, $d\in\mathbb{N}$ and $r\geq d$. For every $r$-local lca-expressed existential formula $\varphi(\overline{x})=\exists\overline{y}\;\big{(}\zeta(\overline{x},\overline{y}\wedge\mathrm{ball}(\{\overline{x},\overline{y}\},x_{1},r))\big{)}$ in $\mathrm{FO}[\Lambda]$, there exists a set of labels $\widehat{\Lambda}$, a lca-reduced formula $\widehat{\varphi}$ and a distributed algorithm performing in $\mathcal{O}(1)$ round in the CONGEST model satisfying the following. For every $\Lambda$-labeled graph $(G,\ell)$, every subgraph $F\in\mathcal{F}_{d}$ of $G$:
  • every node $v\in V(G)$ is given as input: a set of labels $\ell(v)\subseteq\Lambda$, whether it belongs to $F$, and, if applicable, its depth and parent in $F$;
  • every node $v\in V(G)$ outputs a set of labels ^$\widehat{\ell}(v)\subseteq\widehat{\Lambda}$ such that $\widehat{\Lambda},\widehat{\varphi},\widehat{\ell}$ form a $d$-local reduction of $\Lambda,\varphi,\ell$.
Proof.
Let $k=|\overline{x}|$. As for the proof of Lemma 3.7, we assume w.l.o.g. that $k\geq 1$, and we treat only the case where $|\overline{y}|=1$ (as the general case $|\overline{y}|\geq 1$ can be obtained by iteratively eliminating the $|\overline{y}|$ quantifiers one by one). In the same way they are defined in Lemma 3.7, let us define the set $I\subseteq~{}\textsf{Type}(k+1,D,\Lambda)$, and, for each $\psi\in I$, the label sets $\widehat{\Lambda}^{\psi}$, the lca-reduced formula $\widehat{\psi}$, and the $\hat{\Lambda}^{\psi}$-labeling ^$\widehat{\ell}^{\psi}$ of $F$. Furthermore, we define
$\widehat{\Lambda}=\bigcup_{\psi\in I}\widehat{\Lambda}^{\psi},\;\widehat{\varphi}=\bigvee_{\psi\in I}\widehat{\psi},\text{ and, for every node $u\in F$, }\widehat{\ell}(u)=\bigcup_{\psi\in I}\widehat{\ell}^{\psi}(u).$
We then extend $\widehat{\ell}$ to the nodes of $G\smallsetminus F$ by merely defining ^$\widehat{\ell}(v)=\ell(v)$ for all $v\notin V(F)$. Since $\textsf{Type}(k+1,d,\Lambda)$ is a finite set, it is enough to show that, for every $\psi\in I$, the label set $\widehat{\Lambda}^{\psi}$, and the formulas $\widehat{\psi}$ and ^$\widehat{\ell}^{\psi}$ can be computed in $\mathcal{O}(1)$ rounds. In fact, the definitions of the set $I$, $\widehat{\Lambda}$, and $\widehat{\varphi}$ do not depend in the sub-forest $F$ of $G$, so $I$, $\widehat{\Lambda}$, and $\widehat{\varphi}$ can be computed at every node without any communication. It remains to show how to compute, for any given $\psi\in I$, the value of ^$\widehat{\ell}^{\psi}(v)$ at all nodes $v$ in a constant number of rounds.
Initially the algorithm sets ^$\widehat{\ell}^{\psi}(v)=\ell(v)$ at every node. Let $\gamma$ and $\delta$ be the functions such that $\psi=\textsf{type}_{\gamma,\delta}$. Let $s$, $h_{s}$, $h_{y}$, and $h$ be as defined in the proof of Lemma 3.7. Observe that every node knows $\varphi$, and thus every node can compute $\psi$, $h_{s}$, $h_{y}$, and $h$ without any communication. Furthermore, every node knows whether its depth is equal to $h$, $h_{s}$, or $h_{y}$, or none of these three values.
We now revisit the three cases treated in the proof of Lemma 3.7.
Case 1: $\delta(s,k+1)\geq 0$ and $h_{y}=h$.
In this case, the algorithm just needs to identify the good nodes in $F$. To do so, for every $i\in[p]$, every node $v$ at depth $h_{y}$ in $T_{i}$ checks whether it is a candidate node. Every such nodes does not need any communication for this, as it knows its depth in $T_{i}$, as well as $\ell(v)$ and $\gamma(v)$. Then, the algorithm broadcasts ping messages (on a single bit) downward every tree, from every candidate node. These ping messages are forwarded by their descendants downward every $T_{i}$, for $h_{s}-h$ rounds. More precisely, each candidate node sends a ping as a single bit 1 to all its children, and, for $h-h_{s}-1$ rounds, each node that receives that ping from its parent forwards it to all its children. Finally, every node $u$ at depth $h_{s}$ that receives the ping know that their ancestor at level $h_{y}$ are candidates, and thus they mark themselves as good nodes. Moreover, every such node $u$ sets ^$\widehat{\ell}^{\psi}(u)=\ell(u)\cup\{\textsf{good}\}$. Every other node $u$ in $F$ sets ^$\widehat{\ell}^{\psi}(u)=\ell(u)$. All these operations consume $h_{s}-h=\mathcal{O}(d)$ rounds.
Case 2: $\delta(1,k+1)\geq 0$ and $h_{y}>h$.
[…]
Case 3: $\delta(1,k+1)=-1$.
As mentioned in the proof of Lemma 3.8, this case cannot happen for local formulas.
We conclude that for every $\psi\in I$, every node $u$ can compute $\ell^{\psi}(u)$, in $\mathcal{O}(d)$ rounds. Therefore, the total number of rounds sufficient for computing $\widehat{\ell}$ is $\mathcal{O}(d\cdot|I|)=\mathcal{O}(d\cdot 2^{k|\Lambda|}(d+1)^{k^{2}})$. ∎
3.2 Quantifier Elimination on Graphs of Bounded Treedepth
Given a graph $G=(V,E)$, a decomposition forest $F$ of $G$ is a rooted spanning forest of $G$ satisfying that for every edge $e\in E$, one of its endpoints is an ancestor of the other in $F$. Observe that if $G$ is connected then any decomposition forest of $G$ is actually a tree, and every DFS tree of $G$ is a decomposition tree of $G$. For every $t>0$, any induced subgraph $H$ of $G$ with treedepth at most $t$ is called a treedepth-$t$ induced subgraph.
We stress the fact that a decomposition forest $F$ of $G$ must be a subgraph of $G$, which is very convenient for distributed algorithms because this enables communication along the edges of $F$. Note that there exist graphs of treedepth $t$ in which any decomposition forest is of diameter $\Omega(2^{t})$. This is for example the case of paths on $12^{t}-$ 1 vertices. Nevertheless, graphs of treedepth at most $t$ have paths of at most $12^{t}-$ 1 vertices. In particular, for any class of graphs of bounded treedepth, all possible decomposition forests also have trees of constant (although exponentially larger) diameter.
Proposition 3.10 ([54]).
Any decomposition forest of a graph of treedepth at most $t$ belongs to $\mathcal{F}_{2^{t}}$.
In the following lemma, we show that any distributed algorithm aiming at checking a quantifier-free formula in a treedepth-$t$ induced subgraph $H$ of a labeled graph $G$, may rather check an equivalent formula on a decomposition forest of $H$.
Lemma 3.11.
Let $\Lambda$ be a finite set of labels, and let $t>0$. For every quantifier free formula $\varphi(\overline{x})\in\mathrm{FO}[\Lambda]$, there exists a finite set $\widehat{\Lambda}$ depending only on $t$ and $\Lambda$, and an lca-reduced formula $\widehat{\varphi}(\overline{x})\in\mathrm{FO}[\widehat{\Lambda}]$ depending on $\varphi$, $\Lambda$, and $t$, such that, for every $\Lambda$-labeled treedepth-$t$ induced subgraph $(H,\ell)$ of a $\Lambda$-labeled graph $(G,\ell)$, and for every decomposition forest $F$ of $H$, there exists a $\widehat{\Lambda}$-labeling $\widehat{\ell}$ of $G$ such that
$\textsf{true}(\varphi,H,\ell)=\textsf{true}(\widehat{\varphi},F,\widehat{\ell}).$
Moreover, there is a 1-round CONGEST algorithm that computes $\widehat{\ell}$, assuming each node $v$ is given as input $\ell(v)$, whether it belongs to $F$ and, if applicable, its depth and parent in $F$.
Proof.
Each node $u$ initially knows $\Lambda$, $\varphi$, $\ell(u)$, $t$, whether it belongs to $H$, and, if applicable, its depth and parent in $F$. Node $u$ returns $\widehat{\Lambda}$, $\widehat{\varphi}$, and ^$\widehat{\ell}(u)$. Let $d=2^{t}$, and let us consider a $\Lambda$-labeled graph $(G,\ell)$, a treedepth-$t$ induced subgraph $H$ of $G$, and a decomposition forest $F$ of $H$. The idea of the construction is the following. Each node $u$ of $F$ can be labeled by a subset of $[0,d-1]$ indicating the set of levels of $F$ containing a neighbor of $u$ in $H$. Then, the $\textsf{adj}(x,y)$ predicate can be transformed into label and lca predicates on $x$ and $y$. More precisely, for every node $v\in V(H)$, let $\textsf{Levels}(v)\subseteq[d]$ be the set of levels of $F$ containing ancestors $u$ of $v$ such that $\{u,v\}\in E(H)$. Then, let $\widehat{\Lambda}=\Lambda\cup[0,d-1]$ (we assume w.l.o.g. that $\Lambda\cap[0,d-1]=\emptyset$), and let us consider the $\widehat{\Lambda}$-labeling $\widehat{\ell}$ of $G$ defined by
^$\widehat{\ell}(v)=\begin{cases}\ell(v)\cup\textsf{Levels}(v)&\text{if }v\in V(F),\\ \ell(v)&\text{otherwise.}\end{cases}$
For each $i\in[0,d-1]$, we denote by $\textsf{level}_{i}(x)$ the predicate $\textsf{lab}_{i}(x)$, which is true when $x$ has a neighbor at depth $i$. Finally, we define $\widehat{\varphi}$ by replacing each $\textsf{adj}(x,y)$ predicates by
$\bigvee_{1\leq i<j\leq d-1}\Bigg{[}\bigg{(}\textsf{lca}_{j}(x,x)\wedge\textsf{lca}_{i}(y,y)\wedge\textsf{level}_{i}(x)\bigg{)}\vee\bigg{(}\textsf{lca}_{i}(x,x)\wedge\textsf{lca}_{j}(y,y)\wedge\textsf{level}_{i}(y)\bigg{)}\Bigg{]},$
and every equality predicate by lca predicates as we explained in Lemma 3.1. In this way, we obtain a lca-reduced formula $\widehat{\varphi}$ such that $\textsf{true}(\varphi,H,\ell)=\textsf{true}(\widehat{\varphi},F,\ell)$.
Finally, observe that every node $u$ can compute $\widehat{\Lambda},\widehat{\varphi}$, and ^$\widehat{\ell}(u)$ in a single communication round. In fact $\widehat{\Lambda},\widehat{\varphi}$, and ^$\widehat{\ell}(u)$ for every $u\in V(G)\smallsetminus V(H)$ are computed without any communication, as these values do not depend on the instance. To compute ^$\widehat{\ell}(u)$ for $u\in V(F)$, each node broadcasts its depth in $F$ to all its neighbors. After this communication round, $\textsf{Levels}(u)$ can be set as the set of neighbors with a smaller depth in $F$, and the computation of ^$\widehat{\ell}(u)=\ell(u)\cup\textsf{Levels}(u)$ follows. ∎
We now prove that existential quantifier in formulas on graphs of bounded treedepth can be eliminated efficiently. The reduction results in a lca-reduced formula to be checked on a decomposition forest of the considered graph.
Lemma 3.12.
Let $\Lambda$ be a set of labels, $\varphi(\overline{x})=\exists\overline{y},\zeta(\overline{x},\overline{y})\wedge\mathrm{ball}((\overline{x},\overline{y}),x_{1},r)$ be a $r$-local existential formula in $\mathrm{FO}[\Lambda]$, where $r\in\mathbb{N}$ and $\zeta$ is a quantifier free formula. There exists a set of labels $\widehat{\Lambda}$, and an lca-reduced formula $\widehat{\varphi}(\overline{x})$ such that, for every treedepth-$t$ induced subgraph $(H,\ell)$ of a $\Lambda$-labeled graph $(G,\ell)$, and for every decomposition forest $F$ of $H$, there exists a $\widehat{\Lambda}$-labeling $\widehat{\ell}$ such that:
$\textsf{true}(\varphi,H,\ell)=\textsf{true}(\widehat{\varphi},F,\widehat{\ell})=\textsf{true}(\widehat{\varphi}\wedge\mathrm{ball}(\overline{x},x_{1},2^{t}),F,\widehat{\ell}).$
Moreover, there exists an algorithm computing $\widehat{\ell}$ in $\mathcal{O}(1)$ CONGEST rounds, assuming each node $v$ is given as input $\ell(v)$, whether it belongs to $F$ and, if applicable, its depth and parent in $F$.
The lemma above considers both a graph $H$ of bounded treedepth, and a decomposition forest $F$. The distances in both structures are not the same, and two vertices may be further apart in $F$ than they are in $H$. However, the connected components remain the same in both graphs.
Proof.
We are assuming that each node $u$ initially knows $\Lambda,\varphi,\ell(u)$, whether it belongs to $H$, and, if applicable, its depth and parent in $F$. We consider graphs of treedepth at most $t$, thus their connected components have diameter at most $2^{t}$. If $r<2^{t}$, then, by Lemma 2.5 we can rewrite $\mathrm{ball}(z,x_{1},r)$ as an existential formula of the form $\exists\overline{w},\psi(x_{1},z,\overline{w})\wedge\mathrm{ball}((z,\overline{w}),x_{1},2^{t})$. Thus, without loss of generality, we consider the case $r\geq 2^{t}$. Therefore, the balls of radius $r$ are then exactly the connected components of $H$.
By Lemma 3.11, there exists an lca-reduced formula $\zeta_{1}(\overline{x},\overline{y})$ such that, for every induced labeled subgraph $(H,\ell)$ of treedepth at most $t$, for every decomposition forest $F$ of $H$, there exists a labeling $\ell_{1}$ such that $\textsf{true}(\zeta,H,\ell)=\textsf{true}(\zeta_{1},F,\ell_{1})$, and $\ell_{1}$ can be computed in $\mathcal{O}(1)$ rounds. Moreover, since connected components are the same in $H$ and $F$, we also have
$\textsf{true}(\zeta\wedge\mathrm{ball}((\overline{x},\overline{y}),x_{1},r),H,\ell)=\textsf{true}(\zeta_{1}\wedge\mathrm{ball}((\overline{x},\overline{y}),x_{1},r),F,\ell_{1}).$
The proof is completed by applying Lemma 3.9 to the formula $\exists\overline{y},\zeta_{1}(\overline{x},\overline{y})\wedge\mathrm{ball}((\overline{x},\overline{y}),x_{1},r)$. ∎
3.3 Graphs of Bounded Expansion
Let us fix a graph class $\mathcal{G}$ of expansion $f$. Given a set of labels $\Lambda$, let $\mathcal{G}[\Lambda]$ be the set of $\Lambda$-labeled graphs $(G,\ell)$ such that $G\in\mathcal{G}$. Let $(G,\ell)\in\mathcal{G}[\Lambda]$ be the input instance. For every integer $p$, let $\textsf{Col}(p)$ be the collection of sets $U\subseteq[f(p)]$ with cardinality at most $p$.
3.3.1 Skeletons
We introduce a central notion for our proof of Theorem 1.1.
Definition 3.13.
Let $p\in\mathbb{N}$, and let $G$ be a graph of expansion $f$. A $p$-skeleton of $G$ is a triple $S=(G,c,\delta)$ where $c$ is $(p,f(p))$-treedepth coloring of $G$, and $\delta$ is a function mapping every $U\in\textsf{Col}(p)$ to a decomposition forest $F^{U}$ of the subgraph $G[U]$ induced by nodes with colors in $U$. Labeled $p$-skeletons are pairs $(S,\ell)$, where $S=(G,c,\delta)$ is a $p$-skeleton, and $\ell$ is a labeling of $G$.
Note that, by definition of a $(p,f(p))$-treedepth coloring $c$, the graph $G[U]$ is of treedepth at most $p$ for every $U\in\textsf{Col}(p)$. Hence, by Proposition 3.10, we have $F^{U}\in\mathcal{F}_{2^{p}}$. The following lemma states that a $p$-skeleton can be computed in $\mathcal{O}(\log n)$ rounds in the CONGEST model, where the big-$\mathcal{O}$ notation hides constants depending only of $p$, and of the expansion of the class.
Lemma 3.14.
For every $p\in\mathbb{N}$, there exists an algorithm that constructs a $p$-skeleton of $G$ in $\mathcal{O}(\log n)$ rounds in CONGEST. Specifically, at the end of the algorithm, each node knows its own color, and, for every set $U\in\textsf{Col}(p)$ containing its color, its parent and depth in the forest $F^{U}$ corresponding to the skeleton.
Remark.
In the remaining of the paper, when an algorithm takes a skeleton as input, or outputs a skeleton, it is assumed that the skeleton is represented in a distributed manner, in the way the algorithm in Lemma 3.14 computes it.
Proof.
The algorithm of Proposition 2.9 enables to compute a $(p,f(p))$-treedepth coloring $c$ of $G$, in $\mathcal{O}(\log n)$ rounds. After running this algorithm, each node $u$ knows $c(u)$. Let us order the sets in $\textsf{Col}(p)$ lexicographically. The rest of the algorithm runs in $N=|\textsf{Col}(p)|\leq 2^{f(p)}$ phases, where each phase takes at most $2^{p+1}$ rounds. The $i$-th phase consists of dealing with the $i$th set $U\in\textsf{Col}(p)$. Let us fix one such set $U$. For every $u\in V(G)$, if $c(u)\notin U$, then $u$ does not participate in that phase (it remains idle, and, in particular, it does not forward any messages). Only the nodes $u$ such that $c(u)\in U$ exchange messages during phase $U$. The phase is divided in two steps, each consisting of $2^{p}$ rounds.
[…]
  • In the second step, the algorithm runs in parallel in all components $V_{i}$, $i\in\{1,\dots,h\}$ for computing a decomposition tree $T_{i}$ of $G[U]$ rooted at $r_{i}$. For this purpose, we use the algorithm from [25] (cf. Lemma 11 therein). This algorithm performs in $2^{2p}$ rounds.
At the end of these two steps, every node $u$ knows its depth in its tree $T_{i}$, as well as the identifier of its parent in $T_{i}$. Observe that the collection of trees $\{T_{i}\mid i\in[h]\}$ is a decomposition forest $F^{U}$ of $G[U]$. This completes the description of the phase for set $U$. The whole algorithm runs in $\mathcal{O}(\log n)+|\textsf{Col}(p)|\cdot 2^{2p+1}=\mathcal{O}(\log n)$ rounds. ∎
3.3.2 Reducibility on Graphs of Bounded Expansion
We now define a specific notion of quantifier elimination, denoted reducibility, for graphs of bounded expansion. For this purpose, similarly to how FO formulas on treedepth-$t$ graphs are transformed into FO formulas on their decomposition forests, we will transform FO formulas on graphs of $\mathcal{G}$ into FO formulas on their $p$-skeletons. For this purpose, we define additional predicates for describing the structure of $p$-skeletons.
  • For each $k\in[f(p)]$, the predicate $\textsf{color}_{k}(x)$ checks whether the considered vertex $x$ has color $k$ in the skeleton.
  • For every $U\in\textsf{Col}(p)$, and every $j\in[-1,d-1]$, the predicate $\textsf{lca}^{U}_{j}(x,y)$ is equivalent to the $\textsf{lca}_{j}$ predicates, but on $F^{U}$. That is, $\textsf{lca}^{U}_{j}(x,y)$ is true whenever $x$ and $y$ are colored by colors in $U$, and $\textsf{lca}_{j}(x,y)$ is true on $F^{U}$. Such predicates are called $\textsf{lca}^{U}$ predicates.
A formula $\varphi(\overline{x})\in\mathrm{FO}[\Lambda]$ is called $p$-lca-reduced if it is a quantifier-free, and it contains only label, color and $\textsf{lca}^{U}$ predicates, for some collection of sets $U\in\textsf{Col}(p)$. For such a formula, one can ask whether a labeled $p$-skeleton $(S,\ell)$ satisfies $\varphi$ on some vertices $\overline{v}\in S$, i.e., one can ask whether
$(S,\ell)\models\varphi(\overline{v}),$
where $S=(G,s,\delta)$ is a labeled $p$-skeleton. The set $\textsf{true}(\varphi,S,\ell)$ is then defined accordingly. Note that we abusedly write $\overline{v}\in S$ instead of $\overline{v}\in G$, which itself is an abuse of notation for $\overline{v}\in V$, where $V$ is the set of vertices of $G$.
Definition 3.15.
A formula $\varphi\in\mathrm{FO}[\Lambda]$ is called reducible on $\mathcal{G}$ if there exists a positive integer $p$, a set of labels $\widehat{\Lambda}$, and a $p$-lca-reduced formula $\widehat{\varphi}$ such that, for every labeled graph $(G,\ell)\in\mathcal{G}[\Lambda]$, and every $p$-skeleton $S$ of $G$, there exists a $\widehat{\Lambda}$-labeling $\widehat{\ell}$ of $S$ such that
$\textsf{true}(\varphi,G,\ell)=\textsf{true}(\widehat{\varphi},S,\widehat{\ell}).$
Given a graph $G$ and a $p$-skeleton $S$, the tuple $(p,\widehat{\Lambda},\widehat{\varphi},\widehat{\ell})$ is called the reduction of $(\Lambda,\varphi,\ell)$ on $S$.
Importantly, when considering local formulas, it is also required that the reduction $\widehat{\varphi}$ be a local formula too.
Definition 3.16.
An $r$-local formula $\varphi(\bar{x})\in\mathrm{FO}[\Lambda]$ is said locally reducible on $\mathcal{G}$ if there exists $r^{\prime}\in\mathbb{N}$, such that, for every $(G,\ell)\in\mathcal{G}[\Lambda]$, and every $p$-skeleton $S$ of $G$, there exists a reduction $(p,\widehat{\Lambda},\widehat{\varphi},\widehat{\ell})$ of $(\Lambda,\varphi,\ell)$ on $S$ such that:
$\textsf{true}(\varphi,G,\ell)=\textsf{true}(\widehat{\varphi},S,\widehat{\ell})=\textsf{true}(\widehat{\varphi}\wedge\mathrm{ball}(\overline{x},x_{1},r^{\prime}),S,\widehat{\ell}).$
Note that, when considering distance predicates on a skeleton $S=(G,c,\delta)$, we use the distances on the underlying graph $G$. Our goal is now to show that every (local) formula can be efficiently reduced in $\mathcal{G}$ by a distributed algorithm. We first establish this result for existential formulas.
Lemma 3.17.
Let $\Lambda$ be a set of labels, and let $\varphi(\overline{x})=\exists\overline{y},\psi(\overline{x},\overline{y})\wedge\mathrm{ball}((\overline{x},\overline{y}),x_{1},r)$ be an $r$-local existential formula in $\mathrm{FO}[\Lambda]$, where $\psi$ is a quantifier free formula. For every integer $p\geq r\cdot(|\overline{x}|+|\overline{y}|)$, there exists a set of labels $\widehat{\Lambda}$, and a $p$-lca-reduced formula $\widehat{\varphi}(\overline{x})$ satisfying that, for every $(G,\ell)\in\mathcal{G}[\Lambda]$, and for every $p$-skeleton $S$ of $G$, there exists a $\widehat{\Lambda}$-labeling of $\widehat{\ell}$ such that
$\textsf{true}(\varphi,G,\ell)=\textsf{true}(\widehat{\varphi},S,\widehat{\ell})=\textsf{true}(\widehat{\varphi}\wedge\mathrm{ball}(\overline{x},x_{1},2^{p}),S,\widehat{\ell}).$
Moreover, there exists an algorithm that computes $\widehat{\ell}$ in $\mathcal{O}(1)$ rounds in CONGEST, assuming $S$ and $\ell$ are given as input.
Proof.
Let $p\geq r(|\overline{x}|+|\overline{y}|)$. We apply Lemma 3.12 to $\varphi$, which says that there exists a set of labels $\Lambda_{1}$, and a $lca$-reduced formula $\varphi_{1}(\overline{x})$ such that, for every labeled graph $(H,\ell)$ of treedepth at most $p$, and for every decomposition forest $F$ of $H$, there is a $\Lambda_{1}$ labeling $\ell_{1}$ of $F$ such that
$\textsf{true}(\varphi,H,\ell)=\textsf{true}(\varphi_{1},F,\ell_{1})=\textsf{true}(\varphi_{1}\wedge\mathrm{ball}(\overline{x},x_{1},2^{p}),F,\ell_{1}).$
Let us now consider now a $p$-skeleton $S=(G,s,\delta)$. For each $U\in\textsf{Col}(p)$, the induced subgraph $G[U]$ has treedepth at most $p$, and $F^{U}=\delta(U)$ is a decomposition forest of $G[U]$. Thus, we can compute the corresponding labeling $\ell_{1}^{U}$ in $\mathcal{O}(1)$ rounds. Let us denote by $\widehat{\varphi}^{U}(\overline{x})$ the $p$-lca-reduced formula $\varphi_{1}^{U}(\overline{x})\wedge\textsf{color}_{U}(\overline{x})$. That is, $\varphi_{1}^{U}$ is the formula $\varphi_{1}$ in which all lca predicates of were replaced by $\textsf{lca}^{U}$ predicates. Finally, let us define
$\widehat{\varphi}(\overline{x})=\bigvee_{U\in\textsf{Col}(p)}\widehat{\varphi}^{U}(\overline{x}),$
and, for every node $u$,
^$\widehat{\ell}(u)=\bigcup_{U\in\textsf{Col}(p)}\ell_{1}^{U}(u),$
where, w.l.o.g., we assume that the formulas $\varphi^{U}$ use disjoint sets of labels. Let us prove that this formula has the desired property. Let $(G,\ell)\in\mathcal{G}[\Lambda]$, and let $S$ be a $p$-skeleton of $G$. We prove the equivalence of the following three statements:
  1. (i)
    $(G,\ell)\models\varphi(\overline{v})$,
  2. (ii)
    $(S,\widehat{\ell})\models\widehat{\varphi}(\overline{v})$, and
  3. (iii)
    $(S,\widehat{\ell})\models\widehat{\varphi}(\overline{v})\wedge\mathrm{ball}(\overline{v},v_{1},2^{p})$.
The implication $(iii)\Rightarrow(ii)$ is obvious. Let us prove $(i)\Rightarrow(iii)$. For this purpose, let us assume that
$(G,\ell)\models\psi(\overline{v},\overline{w})\wedge\mathrm{ball}((\overline{v},\overline{w}),v_{1},r)$
for some $\overline{v},\overline{w}\subseteq V(G)$. Since $p\geq r\cdot(|v|+|w|)$, there exists $U\in\textsf{Col}(p)$ such that $G[U]$ contains both $\overline{v}$ and $\overline{w}$, and also, for every $u\in\overline{v},\overline{w}$, a path of length at most $r$ from $v_{1}$ to $u$. We get that
$(G[U],\ell)\models\psi(\overline{v},\overline{w})\wedge\mathrm{ball}((\overline{v},\overline{w}),v_{1},r).$
The first of the following series of implications is a direct consequence of the latter, and the second implication follows from Lemma 3.12:
$(G,\ell)\models\varphi(\overline{v})$ $\implies\exists U\in\textsf{Col}(p),\overline{v}\subseteq G[U]\wedge(G[U],\ell)\models\varphi(\overline{v})$
$\implies\exists U\in\textsf{Col}(p),\overline{v}\subseteq G[U]\wedge(F^{U},\ell_{1}^{U})\models\varphi_{1}(\overline{v})\wedge\mathrm{ball}(\overline{v},v_{1},2^{p})$
$\implies\left[\exists U\in\textsf{Col}(p),(S,\ell_{1}^{U})\models\widehat{\varphi}^{U}(\overline{v})\right]\wedge\left[\exists U\in\textsf{Col}(p),F^{U}\models\mathrm{ball}(\overline{v},v_{1},2^{p})\right]$
$\implies\left[(S,\widehat{\ell})\models\widehat{\varphi}(\overline{v})\right]\wedge\left[S\models\mathrm{ball}(\overline{v},v_{1},2^{p})\right]$
$\implies(S,\widehat{\ell})\models\widehat{\varphi}(\overline{v})\wedge\mathrm{ball}(\overline{v},v_{1},2^{p})$
Finally we prove $(ii)\Rightarrow(i)$. We have
$(S,\widehat{\ell})\models\widehat{\varphi}(\overline{v})$ $\implies\exists U\in\textsf{Col}(p),(S,\ell_{1}^{U})\models\widehat{\varphi}^{U}(\overline{v})$
$\implies\exists U\in\textsf{Col}(p),\overline{v}\subseteq G[U]\wedge(F^{U},\ell_{1}^{U})\models\varphi_{1}(\overline{v})$
$\implies\exists U\in\textsf{Col}(p),\overline{v}\subseteq G[U]\wedge(G[U],\ell)\models\varphi(\overline{v})$
$\implies(G,\ell)\models\varphi(\overline{v})$
where the penultimate implication follows from Lemma 3.12, and the last from the fact that $\varphi$ is existential (i.e., if the quantified variables exist in a subgraph, then they exist in $G$). This completes the proof. ∎
Lemma 3.18.
Let $p\in\mathbb{N}$, let $\Lambda$ be a set of labels, and let $\varphi(\overline{x})$ be a $p$-$lca$-reduced formula. There exists a set of labels $\widehat{\Lambda}$, and there exists a quantifier free formula $\psi(\overline{x},\overline{y})$ such that, for every $G\in\mathcal{G}$, for every $\Lambda$-labeling $\ell$ of $G$, and for every $p$-skeleton $S$ of $G$, there exists a labeling $\widehat{\ell}$ of $G$ for which
$\textsf{true}\big{(}(\varphi\wedge\mathrm{ball}(\overline{x},x_{1},r)),S,\ell\big{)}=\textsf{true}(\widehat{\varphi},G,\widehat{\ell})$
where
$\widehat{\varphi}(\overline{x})=\exists\overline{y}\,\big{(}\psi(\overline{x},\overline{y})\wedge\mathrm{ball}((\overline{x},\overline{y}),x_{1},r+2^{p})\big{)}.$
Moreover, $\widehat{\ell}$ can be computed without any communication, assuming $S$ and $\ell$ are given as inputs.
The main idea of the proof is to replace each $\textsf{lca}^{U}$-predicate of $\varphi$ by an existential FO formula on the graph $G$ underlying the (labeled) $p$-skeleton $(S,\ell)$. To that end, we augment the labeling $\ell$ with additional labels, which encode the structure of $S$. Finally, we add distance predicates to enforce the locality of the resulting formula.
Proof.
Let us assume that $\varphi$ is written in conjunctive normal form (CNF). Let $d=2^{p}$. We have that
$\neg\textsf{lca}_{i}(x,y)\equiv\neg\textsf{color}_{U}(x)\vee\neg\textsf{color}_{U}(y)\vee\bigvee_{j\in[-1,d]\setminus\{i\}}\textsf{lca}_{j}(x,y).$
Therefore, we can assume that $\varphi$ does not have any predicate of the form $\neg\textsf{lca}_{i}(x,y)$, as it suffices to replace each occurrence of such predicates by the corresponding disjunction of predicates. Note that the formula we obtain is still in conjunctive normal form. It only remains to replace each predicate $\textsf{lca}_{i}^{U}(x,y)$ by an equivalent existential formula: since $\varphi$ is in conjunctive normal form and no $\textsf{lca}^{U}$ predicate is negated, the formula we obtain is existential. To that end, we add to $\Lambda$ new labels representing the structure of the $p$-skeleton: let
$\widehat{\Lambda}=\Lambda\cup\{\textsf{color}_{k},k\in[f(p)]\}\cup\{\textsf{depth}_{k}^{U},U\in\textsf{Col}(p),k\in[2^{p}]\}.$
We assume w.l.o.g. that $\Lambda$ does not already contain these labels. We construct the labeling $\widehat{\ell}$ as follows, without any communication. Each node $v$ adds its own color to its labeling, and, for each $U\in\textsf{Col}(p)$ that contains $v$’s color, node $v$ adds its depth in $F^{U}$ to its label.
Consider now a predicate $\textsf{lca}_{i}^{U}(y,z)$. We define a formula $\zeta_{i}^{U}\in\mathrm{FO}[\widehat{\Lambda}]$ such that $(S,\ell)\models\textsf{lca}_{i}^{U}(u,v)\iff(G,\widehat{\ell})\models\zeta_{i}^{U}(u,v)$.
$\zeta_{i}^{U}(y,z)=$ $\exists(y_{0},\dots,y_{d})\;\exists(z_{0},\dots,z_{d})\;\bigvee_{(a,b)\in[\max\{0,i\},d-1]^{2}}\Biggl{(}(y_{a}=y)\wedge(z_{b}=z)$
$\wedge\Big{(}\bigwedge_{s\in[0,i]}y_{s}=z_{s}\Big{)}\wedge\Big{(}\bigwedge_{s\in[i+1,\min(a,b)]}y_{s}\neq z_{s}\Big{)}\wedge\textsf{depth}_{0}^{U}(y_{0})\wedge\textsf{depth}_{0}^{U}(z_{0})$
$\wedge\Big{(}\bigwedge_{s\in[1,a]}\textsf{depth}_{s}^{U}(y_{s})\wedge\textsf{adj}(y_{s-1},y_{s})\Big{)}\wedge\Big{(}\bigwedge_{s\in[1,b]}\textsf{depth}_{s}^{U}(z_{s})\wedge\textsf{adj}(z_{s-1},z_{s})\Big{)}\Biggr{)}.$
Indeed, $a$ (resp. $b$) is the depth of $y$ (resp. $z$) in $F^{U}$, and $y_{0},\dots,y_{a}$ (resp. $z_{0},\dots,z_{b}$) correspond to the path from the root of $y$’s tree to $y$ (resp. from the root of $z$’s tree to $z$). The equality predicates on the first line, and the depth predicates on the second line guarantee that the endpoints of the path are as expected. The equality predicates on the second line guarantee that both paths share exactly $i+1$ vertices. The depth and adjacency predicates on the third line guarantee that, for every $s\in[1,a]$ (resp., every $s\in[1,b]$), $y_{s-1}$ (resp., $z_{s-1}$) is the parent of $y_{s}$ (resp., $z_{s}$). By construction, we can assume that each of the $y_{i}$ (resp., $z_{i}$) we introduced are at distance at most $d$ from $y$ (resp., $z$).
We would like to replace each occurrence of $\textsf{lca}_{i}^{U}$ in $\varphi$ by $\zeta_{i}^{U}$, while preserving locality. To that end, we add a distance predicate to the formula $\zeta_{i}^{U}$, which enforces locality. Recall that we are interested by the formula $\varphi\wedge\mathrm{ball}(\overline{x},x_{1},r)$. Whenever the formula is satisfied, every $x_{i}$ is at distance at most $r$ from $x_{1}$. This implies that every $y_{s}$ and $z_{s}$ introduced in $\zeta_{i}^{U}(y,z)$ is at distance at most $r+d$ from $x_{1}$. We thus replace each $\textsf{lca}_{i}^{U}(y,z)$ predicate in $\varphi$ by the following formula:
$\exists(y_{0},\dots,y_{d})\;$ $\exists(z_{0},\dots,z_{d})\;\Bigg{(}\mathrm{ball}((\overline{y},\overline{z}),x_{1},r+d)\wedge\bigvee_{(a,b)\in[\max(0,i),2^{p}]^{2}}\Biggl{[}(y_{a}=y)\wedge(z_{b}=z)$
$\wedge$ $\Big{(}\bigwedge_{s\in[0,i]}y_{s}=z_{s}\Big{)}\wedge\Big{(}\bigwedge_{s\in[i+1,\min\{a,b\}]}y_{s}\neq z_{s}\Big{)}\wedge\textsf{depth}_{0}^{U}(y_{0})\wedge\textsf{depth}_{0}^{U}(z_{0})$
$\wedge$ $\Big{(}\bigwedge_{s\in[1,a]}\big{(}\textsf{depth}_{s}^{U}(y_{s})\wedge\textsf{adj}(y_{s-1},y_{s})\big{)}\Big{)}\wedge\Big{(}\bigwedge_{s\in[1,b]}\big{(}\textsf{depth}_{s}^{U}(z_{s})\wedge\textsf{adj}(z_{s-1},z-s)\big{)}\Big{)}\Biggr{]}\Bigg{)}$
Let us denote by $\tilde{\varphi}(\overline{x})$ this formula. By the definition of $\beta_{r,r^{\prime}}$ in Proposition 2.5, we can then define
$\widehat{\varphi}(\overline{x})=\tilde{\varphi}(\overline{x})\wedge\beta_{r,r+d}(\overline{x})\wedge\mathrm{ball}(\overline{x},x_{1},r+d).$
By construction, $\widehat{\varphi}$ is existential, and thus
$\textsf{true}(\widehat{\varphi},G,\widehat{\ell})=\textsf{true}(\widehat{\varphi},G,\widehat{\ell})\cap\textsf{true}(\mathrm{ball}(\overline{x},x_{1},r),G,\widehat{\ell})=\textsf{true}(\varphi\wedge\mathrm{ball}(\overline{x},x_{1},r),S,\ell),$
which completes the proof. ∎
We now tackle the general case.
Lemma 3.19.
Let $\Lambda$ be a set of labels, and let $\varphi(\overline{x})$ be an FO formula in $r$-local form. There exists $p\in\mathbb{N}$, a set of labels $\widehat{\Lambda}$, and a $p$-$lca$-reduced formula $\widehat{\varphi}(\overline{x})$ such that, for every labeled graph $(G,\ell)\in\mathcal{G}[\Lambda]$, and for every $p$-skeleton $S$ of $G$, there exists a $\widehat{\Lambda}$-labeling $\widehat{\ell}$ for which
$\textsf{true}(\varphi,G,\ell)=\textsf{true}(\widehat{\varphi},S,\widehat{\ell})=\textsf{true}(\widehat{\varphi}\wedge\mathrm{ball}(\overline{x},x_{1},2^{p}),S,\widehat{\ell})$
Moreover $\widehat{\ell}$ can be computed in $\mathcal{O}(\log n)$ rounds in the CONGEST model, assuming $S$ and $\ell$ are given as inputs.
Proof.
The proof proceeds by induction on the structure of the formulas written in local form as defined (in an inductive way) in Definition 2.3. We consider the three cases occurring in Definition 2.3.
Case 1 (base case).
$\varphi(\overline{x})=\psi(\overline{x})\wedge\mathrm{ball}(\overline{x},x_{1},r)$, where $\psi$ is quantifier free. We can directly apply Lemma 3.17 with $|\overline{y}|=0$, to compute $\widehat{\ell}$ from the skeleton $S$ and labeling $\ell$ given as inputs.
Case 2.
$\varphi(\overline{x})=\exists y\;\psi(\overline{x},y)$, where $\psi$ is in local form. By induction, there exists $p^{\prime}$, $\Lambda_{1}$, and $\psi_{1}(\overline{x},y)$ such that, for every $(G,\ell)\in\mathcal{G}[\Lambda]$, and every $p^{\prime}$-skeleton $S^{\prime}$ of $G$, there exists a $\Lambda_{1}$-labeling $\ell_{1}$ for which
$\textsf{true}(\psi,G,\ell)=\textsf{true}(\psi_{1},S^{\prime},\ell_{1})=\textsf{true}(\psi_{1}\wedge\mathrm{ball}((\overline{x},y),x_{1},2^{p^{\prime}}),S^{\prime},\ell_{1})$
By Lemma 3.14, a $p^{\prime}$-skeleton $S^{\prime}$ of $G$ can be constructed in $\mathcal{O}(\log n)$ rounds in CONGEST. By induction, the corresponding labeling $\ell_{1}$ can be computed in $\mathcal{O}(\log n)$ rounds. For eventually applying Lemma 3.17 on $\exists y\,\psi_{1}(\overline{x},y)$, we first apply Lemma 3.18 on $\psi_{1}$, with $r=2^{p^{\prime}}$. Thanks to this lemma, there exists a set of labels $\Lambda_{2}$, a quantifier-free formula $\rho_{2}(\overline{x},y,\overline{z})$, and a labeling $\ell_{2}$ that can be constructed from $\ell_{1}$ such that
$\textsf{true}(\psi_{1}\wedge\mathrm{ball}((\overline{x},y),x_{1},2^{p^{\prime}}),S^{\prime},\ell_{1})=\textsf{true}(\psi_{2}(\overline{x},y),G,\ell_{2}),$
where $\psi_{2}(\overline{x},y)=\exists\overline{z}\,\rho_{2}(\overline{x},y,\overline{z})\wedge\mathrm{ball}((\overline{x},y,\overline{z}),x_{1},2^{p^{\prime}+1})$. Let us then define $\varphi_{2}(\overline{x})=\exists y\,\psi_{2}(\overline{x},y)$. This is formula is existential since $\psi_{2}$ is existential. We can thus apply Lemma 3.17 to $\varphi_{2}$, from which the induction step follows. In particular, we can compute $\widehat{\ell}$ from the skeleton $S$ given as input.
Case 3.
$\varphi(\overline{x})=\mathrm{ball}(\overline{x},x_{1},r)\wedge\forall y\,\Big{(}\neg\mathrm{ball}(\overline{x},y),x_{1},r)\vee\psi(\overline{x},y)\Big{)}$ where $\psi$ is in $r$-local form. We can rewrite $\varphi$ as
$\varphi(\overline{x})=\mathrm{ball}(\overline{x},x_{1},r)\wedge\neg\exists y,\mathrm{ball}((\overline{x},y),x_{1},r)\wedge\neg\psi(\overline{x},y).$
By induction on $\psi$, there exist $p^{\prime}\in\mathbb{N}$, a set of labels $\Lambda_{1}$, and a $p^{\prime}$-lca-reduced formula $\psi_{1}(\overline{x},y)$ such that, for every labeled graph $(G,\ell)\in\mathcal{G}[\Lambda]$, and for every $p^{\prime}$-skeleton $S^{\prime}$ of $G$, there exists a $\Lambda_{1}$-labeling $\ell_{1}$ satisfying
$\textsf{true}(\psi,G,\ell)=\textsf{true}(\psi_{1},S^{\prime},\ell_{1}).$
By Lemma 3.14, a $p^{\prime}$-skeleton $S^{\prime}$ of $G$ can be constructed in $\mathcal{O}(\log n)$ rounds in the CONGEST model. By induction, the labeling $\ell_{1}$ can thus be computed in $\mathcal{O}(\log n)$ rounds. Let us then take the negation of $\psi_{1}$. Note that the resulting formula $\neg\psi_{1}$ is still a $p^{\prime}$-lca-reduced formula. We then apply Lemma 3.18 to $\neg\psi_{1}$, which results in a set of labels $\Lambda_{2}$, an existential formula $\rho_{2}$ in $(r+2^{p})$-local form, and a $\Lambda_{2}$-labeling $\ell_{2}$ of $G$ such that
$\textsf{true}(\neg\psi_{1}\wedge\mathrm{ball}(\overline{x},x_{1},r),S^{\prime},\ell_{1})=\textsf{true}(\rho_{2},G,\ell_{2}).$
As a consequence, we get
$(G,\ell)\models\mathrm{ball}((\overline{v},u),v_{1},r)\wedge\neg\psi(\overline{v},u)$ $\iff\Bigl{(}G\models\mathrm{ball}((\overline{v},u),v_{1},r)\Bigr{)}\wedge\Bigl{(}(S^{\prime},\ell_{1})\models\neg\psi_{1}(\overline{v},u)\Bigr{)}$
$\iff(S^{\prime},\ell_{1})\models\mathrm{ball}((\overline{v},u),v_{1},r)\wedge\neg\psi_{1}(\overline{v},u)$
$\iff(G,\ell_{2})\models\rho_{2}(\overline{v},u)$
where the first equivalence is by induction, and the last by application of Lemma 3.18.
Let us now focus on the formula
$\eta_{2}(\overline{x})=\exists y\;\rho_{2}(\overline{x},y).$
Since $\rho_{2}$ is existential, and since it is in $(r+2^{p^{\prime}})$-local form, the formula $\eta_{2}$ is also existential and in $(r+2^{p^{\prime}})$-local form. By Lemma 3.17, there exists an integer $p\geq r\cdot|\overline{x}|$, a set of labels $\Lambda_{3}$, and a $p$-$lca$-reduced-formula $\eta_{3}(\overline{x})$ such that, for every $p$-skeleton $S$ of $G$ (and in particular the skeleton $S$ given as input), there exists a labeling $\ell_{3}$ for which
$\textsf{true}(\eta_{2},G,\ell_{2})=\textsf{true}(\eta_{3},S,\ell_{3}).$
Once again, the negation of $\eta_{3}$ is still $p$-$lca$-reduced.
We also need to transform the first half of $\varphi$, namely $\mathrm{ball}(\overline{x},x_{1},r)$, into a $p$-$lca$-reduced formula. The formula $\mathrm{ball}(\overline{x},x_{1},r)$ is in fact an existential formula in $r$-local form (with zero existential quantifiers). Therefore, we can apply Lemma 3.17. We can even apply this lemma with the same $p$ since $p$ satisfies the required condition $p\geq r|\overline{x}|$. Thus, there exists a set of label $\Lambda_{\gamma}$, and a $p$-lca-reduced formula $\gamma(\overline{x})$ such that, for every $G\in\mathcal{G}$, and every $p$-skeleton $S$ of $G$, there exists a $\Lambda_{\gamma}$-labeling $\ell_{\gamma}$ for which
$\textsf{true}(\mathrm{ball}(\overline{x},x_{1},r)),G)=\textsf{true}(\gamma,S,\ell_{\gamma})=\textsf{true}(\gamma\wedge\mathrm{ball}(\overline{x},x_{1},2^{p}),S,\ell_{\gamma}).$
Finally, let $\widehat{\Lambda}=\Lambda_{\gamma}\cup\Lambda_{3}$ where, w.l.o.g., we assume that $\Lambda_{\gamma}$ and $\Lambda_{3}$ are disjoint, let
$\widehat{\varphi}(\overline{x})=\gamma(\overline{x})\wedge\neg\eta_{3}(\overline{x}),$
and for every node $v$ let ^$\widehat{\ell}(v)=\ell_{\gamma}(v)\cup\ell_{3}(v)$. We get
$\textsf{true}(\varphi,G,\ell)$ $=\textsf{true}\big{(}\mathrm{ball}(\overline{x},x_{1},r),G,\ell\big{)}\cap\textsf{true}\Big{(}\neg\exists y\,\big{(}\mathrm{ball}((\overline{x},y),x_{1},r)\wedge\neg\psi(\overline{x},y)\big{)},G,\ell\Big{)}$
$=\textsf{true}\big{(}\mathrm{ball}(\overline{x},x_{1},r),G,\ell_{2}\big{)}\cap\textsf{true}(\neg\eta_{2},G,\ell_{2})$
$=\textsf{true}\big{(}\gamma(\overline{x}),S,\ell_{\gamma}\big{)}\cap\textsf{true}\big{(}\neg\eta_{3}(\overline{x}),S,\ell_{3}\big{)}$
$=\textsf{true}\big{(}\widehat{\varphi},S,\widehat{\ell}\big{)}.$
Moreover, since $\textsf{true}(\gamma,S,\ell_{\gamma})=\textsf{true}(\gamma\wedge\mathrm{ball}(\overline{x},x_{1},2^{p}),S,\ell)$, we also have
$\textsf{true}(\widehat{\varphi},S,\widehat{\ell})=\textsf{true}\big{(}\widehat{\varphi}\wedge\mathrm{ball}(\overline{x},x_{1},2^{p}),S,\widehat{\ell}\big{)}.$
We complete the proof by studying the round-complexity of computing the labeling $\widehat{\ell}$ in CONGEST. For each of the lemmas used in the proof, the construction of the labeling is performed in $\mathcal{O}(1)$ rounds. For cases 2 and 3, we also need to construct a skeleton, which can be done in $\mathcal{O}(\log n)$ rounds. Thus, the overall round complexity is $\mathcal{O}(\log n)$, as claimed. ∎
3.3.3 Proof of Theorem 1.1
Let $\varphi(x)$ be a $r$-local formula. By Lemma 2.4, it can be rewritten in $r$-local form. Then, by Lemma 3.19, for some $p\in\mathbb{N}$ $\varphi$ can be reduced in $\mathcal{O}(\log n)$ rounds to a $p$-lca-reduced formula $\widehat{\varphi}(x)$. Using Lemma 3.14, we can build a $p$-skeleton $S$ in $\mathcal{O}(\log n)$ rounds. Then, from an input labeling $\ell$, the reduced labeling $\widehat{\ell}$ can be computed in $\mathcal{O}(\log n)$ rounds. It satisfies:
$(G,\ell)\models\varphi(v)\iff(S,\widehat{\ell})\models\widehat{\varphi}(v)$
Since $\widehat{\varphi}(x)$ is quantifier free, each node $v$ can check locally without additional communication whether it satisfies $\widehat{\varphi}$, which completes the proof.
4 Distributed Model Checking of General Formulas
In this section, we establish Theorem 1.2. We thus consider an arbitrary FO formulas — not necessarily local. Actually, we show a result stronger than Theorem 1.2. Given a formula $\varphi(x)$ with one free variable, we provide a CONGEST algorithm that marks all nodes of $n$-node graphs with bounded expansion that satisfy the formula, in $\mathcal{O}(D+\log n)$ rounds, where $D$ is the diameter of the input graph [^4] This problem is more challenging than simply checking whether the graph satisfies a formula $\varphi$, i.e., stronger than the model checking problem as, given an FO formula $\varphi$ without free variables, one can define a logically equivalent formula $\varphi^{\prime}(x)$ by adding a dummy free variable, so that the graph satisfies $\varphi$ if and only if it satisfies $\varphi^{\prime}(v)$ at each of its nodes $v$.. To prove the result, we need to revisit the proof of Theorem 1.1, which involves quantifier elimination of existential formulas first over forests of bounded-depth, then over graphs of bounded tree-depth, and finally over graphs with bounded expansion. Fortunately, once we have resolved quantifier elimination over forests bounded-depth, the extensions to the other two cases will follow from the arguments already developed in the previous section for local formulas. Therefore, it will be sufficient to revisit quantifier elimination for existential formulas on forests of bounded-depth, and then proceed by induction to cover the general case.
4.1 Rooted Forests of Bounded Depth
As established in the proof of Lemma 3.7, an existential formula over bounded-depth forests corresponds to checking the presence of “subpatterns” (representing subforests) in which the free variables of the formula can be mapped onto. One of these free variables, quantified existentially, is to be eliminated. Depending on the position of this existentially quantified variable $y$, we distinguish three cases. In the first two case, $y$ shares common ancestors with the other free variables — these cases are the “connected cases”. In the third case, $y$ belongs to a connected component that is different from all the other variables in the formula. When dealing with local formulas, we only needed to handle the connected cases. However, for general formulas, we must also address the remaining case. This is for this “disconnected case” that our algorithm requires $\mathcal{O}(D)$ additional rounds.
Lemma 4.1.
Let $\varphi(\overline{x})=\exists\overline{y}\;\zeta(\overline{x},\overline{y})$ be a formula in $\mathrm{FO}[\Lambda]$. There exists $\widehat{\Lambda}$, a lca-reduced formula $\widehat{\varphi}(\overline{x})$ and a distributed algorithm performing in $\mathcal{O}(D)$ rounds in the CONGEST model satisfying the following. For every $\Lambda$-labeled graph $(G,\ell)$ of diameter $D$, and for every subgraph $F\in\mathcal{F}_{d}$ of $G$,
  • every node $v\in V(G)$ is given as input its labeling $\ell(v)$, whether it belongs to $F$, and whether it is a root of a tree in $F$ (if not, it also receives as input its depth and parent in $F$);
  • every node $v\in V(G)$ outputs a set of labels ^$\widehat{\ell}(v)$, such that $\widehat{\Lambda},\widehat{\varphi},\widehat{\ell}$ form a reduction of $\Lambda,\varphi,\ell$.
Proof.
We follow the same guidelines as in the proof of Lemma 3.9 for Cases 1 and 2. It is thus sufficient to consider Case 3, as described in Lemma 3.7. We use the same notations as in these lemmas. Recall that Case 3 assumes
$\delta(1,k+1)=-1.$
The algorithm consists in four phases.
  1. 1.
    During the first phase, the algorithm aims at identifying the active roots. First, the candidate nodes are identified in the same way as in the previous cases 1 and 2. By broadcasting a ping message upward each tree from the candidate nodes for $h_{y}$ rounds, each root becomes aware of the presence of a candidate node in its tree. The roots that receive a ping message mark themselves as active roots. This phase takes $\mathcal{O}(d)$ rounds.
  2. 2.
    During the second phase, the algorithm aims at identifying the good children, by broadcasting a ping message downward from every active root. Every node $u$ that receives such ping message on this phase marks itself as good, and adds the label good to ^$\widehat{\ell}^{\psi}(u)$. This phase takes $\mathcal{O}(d)$ rounds. At this point, the algorithm constructs a BFS tree $T$ of $G$, rooted at node $r_{1}$. This procedure takes $\mathcal{O}(D)$ rounds.
  3. 3.
    During the third phase, the root $r_{1}$ gets the number $\rho(F)$ of active roots in $F$, by counting upward $T$. More specifically, let us denote by $d_{T}=\mathcal{O}(D)$ the depth of $T$. The counting proceeds in $d_{T}+1$ round. For each node $u$ in $T$, let $a(u)$ be $1$ if $u$ is an active root, and 00 otherwise. In the first round, each node $u$ at depth $d_{T}$ set $\textsf{val}(u)=a(u)$, and sends $\textsf{val}(u)$ to its parent in $T$. Then, at round $i+1$ with $i\in[d_{t}]$, every node $u$ of $T$ at depth $d_{T}-i$ computes $\textsf{val}(u)=a(u)+\sum_{v\in C(u)}\textsf{val}(u)$, where $C(u)$ is the set of children of $u$ in $T$, and sends $\textsf{val}(u)$ to its parent. At the end of this phase, we have $\textsf{val}(r_{1})=\rho(F)$. Observe that, for every node $u$, $\textsf{val}(u)\leq n$, and thus this value can be communicated in one round under the CONGEST model. Therefore, the third phase runs in $\mathcal{O}(D)+d_{T}+1=\mathcal{O}(D)$ rounds.
  4. 4.
    Finally, during the fourth phase, the root $r_{1}$ broadcast $\rho(F)$ downward the tree $T$, to all nodes in $G$. Then, each node $u$ in $F$ adds the label $\min(\rho(F),k+1)$ to $\widehat{F}^{\psi}(u)$. This phase takes $d_{T}=O(D)$ rounds.
We conplete the proof by noticing that, for every $\psi\in I$, every node $u$ can compute $\ell^{\psi}(u)$ in $\mathcal{O}(D)$ rounds. Therefore, the total number of rounds sufficient for computing $\widehat{\ell}$ is
$|I|\cdot D\leq c\cdot 2^{k|\Lambda|}(d+1)^{k^{2}}\cdot D,$
where $c$ is a constant not depending on either $d$, $\Lambda$, $\psi$, or $D$. ∎
4.2 Quantifier Elimination on Graphs of Bounded Treedepth
We now deal with the quantifier elimination on graphs of bounded treedepth.
Lemma 4.2.
Let $\Lambda$ be a set of labels, and let $\varphi(\overline{x})=\exists\overline{y},\zeta(\overline{x},\overline{y})$ be an existential formula in $\mathrm{FO}[\Lambda]$, where $\zeta$ is a quantifier free formula. There exists a set of labels $\widehat{\Lambda}$, and a lca-reduced formula $\widehat{\varphi}(\overline{x})$ such that, for every induced subgraph $(H,\ell)$ of treedepth $t$ of a $\Lambda$-labeled graph $(G,\ell)$, and for every decomposition forest $F$ of $H$, there exists a $\widehat{\Lambda}$-labeling $\widehat{\ell}$ such that:
$\textsf{true}(\varphi,H,\ell)=\textsf{true}(\widehat{\varphi},F,\widehat{\ell})$
Moreover, there exists an algorithm computing $\widehat{\ell}$ in $\mathcal{O}(D)$ rounds in the CONGEST model, each node $u$ is given as input $\ell(u)$, whether it belongs to $H$ and, if applicable, its depth and parent in $F$.
Proof.
By Lemma 3.11, there exists an lca-reduced formula $\zeta_{1}(\overline{x},\overline{y})$ such that, for every induced labeled subgraph $(H,\ell)$ of treedepth at most $t$, for every decomposition forest $F$ of $H$, there exists a labeling $\ell_{1}$ such that $\textsf{true}(\zeta,H,\ell)=\textsf{true}(\zeta_{1},F,\ell_{1})$, and $\ell_{1}$ can be computed in $\mathcal{O}(1)$ rounds in the CONGEST  model. It is thus sufficient to merely apply Lemma 4.1 to the formula $\exists\overline{y},\zeta_{1}(\overline{x},\overline{y})$, which takes $\mathcal{O}(D)$ CONGEST rounds. ∎
4.3 Graphs of Bounded Expansion
We can now directly treat the case of graphs of bounded expansion. As we did in the local case, we consider first the case of existential formulas. Let $\mathcal{G}$ be a class of graphs, of bounded expansion.
Lemma 4.3.
Let $\Lambda$ be a set of labels, and let $\varphi(\overline{x})=\exists\overline{y}\,\psi(\overline{x},\overline{y})$ be an existential formula in $\mathrm{FO}[\Lambda]$, where $\psi$ is quantifier-free. For every $p\geq|\overline{x}|+|\overline{y}|$, there exists a set $\widehat{\Lambda}$ of labels, and a $p$-lca-reduced formula $\widehat{\varphi}$ such that, for every $(G,\ell)\in\mathcal{G}[\Lambda]$, and for every $p$-skeleton $S$ of $G$, there exists a $\widehat{\Lambda}$-labeling $\widehat{\ell}$ such that
$\textsf{true}(\varphi,G,\ell)=\textsf{true}(\widehat{\varphi},S,\widehat{\ell}).$
Moreover, there exists a distributed algorithm that computes $\widehat{\ell}$ in $\mathcal{O}(D)$ rounds in CONGEST, assuming $\ell$ and $S$ are given as input.
Proof.
Let us fix an integer $p\geq|\overline{x}|+|\overline{y}|$. By applying Lemma 4.2 to $\varphi$, we get that there exist a set of labels $\Lambda_{1}$, and an $lca$-reduced formula $\varphi_{1}(\overline{x})$ such that, for every labeled graph $(H,\ell)$ of treedepth at most $p$, and for every decomposition forest $F$ of $H$, there exists a $\Lambda_{1}$-labeling $\ell_{1}$ of $F$ such that
$\textsf{true}(\varphi,H,\ell)=\textsf{true}(\varphi_{1},F,\ell_{1}).$
Let us now consider a $p$-skeleton $S$ of $(G,\ell)$. For each $U\in\textsf{Col}(p)$, the induced subgraph $G[U]$ has treedepth at most $p$, and $F^{U}$ is a decomposition forest of $G[U]$. Thus, the corresponding labeling $\ell_{1}^{U}$ can be computed in $\mathcal{O}(D)$ rounds. Let us consider the $p$-lca-reduced formula
$\widehat{\varphi}^{U}(\overline{x})=\varphi_{1}^{U}(\overline{x})\wedge\textsf{color}_{U}(\overline{x}).$
Note that $\varphi_{1}^{U}$ is the formula $\varphi_{1}$ in which all lca predicates were replaced by $\textsf{lca}^{U}$ predicates. We also define
$\widehat{\varphi}(\overline{x})=\bigvee_{U\in\textsf{Col}(p)}\widehat{\varphi}^{U}(\overline{x}).$
Finally, for each node $u$, we set
^$\widehat{\ell}(u)=\bigcup_{U\in\textsf{Col}(p)}\ell_{1}^{U}(u),$
where it is assumed, w.l.o.g., that the formulas $\varphi^{U}$ use disjoint sets of labels. Let us now prove that the formula $\widehat{\varphi}(\overline{x})$ has the desired property. Let $(G,\ell)\in\mathcal{G}[\Lambda]$, let $S$ be a $p$-skeleton of $G$, and let us prove that
$(G,\ell)\models\varphi(\overline{v})\iff(S,\widehat{\ell})\models\widehat{\varphi}(\overline{v}).$
For this purpose, let us assume that $(G,\ell)\models\psi(\overline{v},\overline{w})$ for some $\overline{v},\overline{w}\subseteq V(G)$. Since $p\geq|v|+|w|$, there exists $U\in\textsf{Col}(p)$ such that $G[U]$ contains both $\overline{v}$ and $\overline{w}$. Therefore, $(G[U],\ell)\models\psi(\overline{v},\overline{w})$. Furthermore,
$(G,\ell)\models\varphi(\overline{v})$ $\iff\exists U\in\textsf{Col}(p),\overline{v}\subseteq G[U],(G[U],\ell)\models\varphi(\overline{v})$
$\iff\exists U\in\textsf{Col}(p),\overline{v}\subseteq G[U],(F^{U},\ell_{1}^{U})\models\varphi_{1}(\overline{v})$
$\iff\exists U\in\textsf{Col}(p),(S,\ell_{1}^{U})\models\widehat{\varphi}^{U}(\overline{v})$
$\iff(S,\widehat{\ell})\models\widehat{\varphi}(\overline{v})$
where the second equivalence is thanks to Lemma 4.2. ∎
Next, we adapt Lemma 3.18 to the case of general (i.e., non necessarily local) formulas.
Lemma 4.4.
Let $p\in\mathbb{N}$, let $\Lambda$ be a set of labels, and let $\varphi(\overline{x})$ be a $p$-$lca$-reduced formula. There exists a set of labels $\widehat{\Lambda}$, and a quantifier free formula $\psi(\overline{x},\overline{y})$ such that, for every $G\in\mathcal{G}$, every $p$-skeleton $S$ of $G$, and every $\Lambda$-labeling $\ell$ of $G$, there exists a labeling $\widehat{\ell}$ of $G$ for which
$\textsf{true}(\varphi,S,\ell)=\textsf{true}(\widehat{\varphi},G,\widehat{\ell}).$
Moreover, for every $p$, $\Lambda$ and $\varphi$, there exists a distributed algorithm that, given $S$ and $\ell$ as inputs, computes $\widehat{\ell}$ without any communication.
Proof.
This lemma is a direct consequence of the arguments in the proof of Lemma 3.18. Indeed, we follow the same steps as in that proof, until the predicates $\textsf{lca}_{i}^{U}(y,z)$ has to be replace by the similar existential formula
$\exists(y_{0},\dots,y_{d})\;$ $\exists(z_{0},\dots,z_{d})\;\bigvee_{(a,b)\in[i,2^{p}-1]^{2}}\Biggl{(}(y_{a}=y)\wedge(z_{b}=z)$
$\wedge\Big{(}\bigwedge_{s\in[0,i]}y_{s}=z_{s}\Big{)}\wedge\Big{(}\bigwedge_{s\in[i+1,\min(a,b)]}y_{s}\neq z_{s}\Big{)}\wedge\textsf{depth}_{0}^{U}(y_{0})\wedge\textsf{depth}_{0}^{U}(z_{0})$
$\wedge\Big{(}\bigwedge_{s\in[1,a]}\textsf{depth}_{s}^{U}(y_{s})\wedge\textsf{adj}(y_{s-1},y_{s})\Big{)}\wedge\Big{(}\bigwedge_{s\in[1,b]}\textsf{depth}_{s}^{U}(z_{s})\wedge\textsf{adj}(z_{s-1},z_{s})\Big{)}\Biggr{)}$
This completes the proof. ∎
We are now ready to deal with quantifier elimination of general first-order formulas.
Lemma 4.5.
Let $\Lambda$ be a set of labels, and let $\varphi(\overline{x})$ be a formula in $\mathrm{FO}[\Lambda]$. There exists $p\in\mathbb{N}$, a set of labels $\widehat{\ell}$, and a $p$-lca-reduced formula $\widehat{\varphi}(\overline{x})$ such that, for every $(G,\ell)\in\mathcal{G}[\Lambda]$, and every $p$-skeleton $S$ of $G$, there exists a $\widehat{\ell}$-labeling $\widehat{\ell}$ of $G$ such that
$\textsf{true}(\varphi,G,\ell)=\textsf{true}(\widehat{\varphi},S,\widehat{\ell}).$
Moreover, there exists a distributed algorithm that computes the labeling $\widehat{\ell}$ in $\mathcal{O}(D+\log n)$ rounds, assuming $\ell$ and $S$ are given as input.
Proof.
We prove the lemma by designing the algorithm by induction on the quantifier depth $t$ of the formula. For the base case, $\varphi(\overline{x})$ is quantifier-free. Let $p=|\overline{x}|$. Thanks to Lemma 3.14, the $p$-skeleton $S$ can be computed in $\mathcal{O}(\log n)$ rounds. We can then apply Lemma 4.3 with $|\overline{y}|=0$, and return the reduction of $(\Lambda,\varphi,\ell)$ on $S$.
Now let us suppose that the statement of the lemma holds for all formulas with quantifier depth $t\geq 0$. Let $\varphi(\overline{x})=Qy,\psi(\overline{x},y)$, where $Q\in\{\exists,\forall\}$, and $\psi$ is a formula of quantifier depth $t$. We study $Q=\exists$ and $Q=\forall$ separately.
Existential quantifier.
By induction hypothesis, an integer $p_{1}$, a set $\Lambda_{1}$, a $p_{1}$-lca-reduced formula $\psi_{1}(\overline{x},y)$, a set of labels $\ell_{2}$, and a $p_{1}$-skeleton $S$ such that
$\textsf{true}(\psi,G,\ell)=\textsf{true}(\psi_{1},S,\ell_{1})$
can be computed in $\mathcal{O}(D+\log n)$ rounds. One would then like to apply Lemma 4.3 on $\exists y\,\psi_{1}(\overline{x},y)$. However, $\psi_{1}$ is not defined over graphs of bounded expansion, but over $p_{1}$-skeletons. Therefore, we first apply Lemma 4.4 on $p_{1},\Lambda_{1},\psi_{1},G,S,\ell_{1}$, to compute $\Lambda_{2},\psi_{2}(\overline{x},y)$, and $\ell_{2}$ satisfying
$\textsf{true}(\psi_{1},S,\ell_{1})=\textsf{true}(\psi_{2},G,\ell_{2})$
in $\mathcal{O}(\log n)$ rounds, where $\psi_{2}$ is existential, of the form $\psi_{2}(\overline{x},y)=\exists\overline{z},\rho(\overline{x},y,\overline{z})$, where $\rho\in\mathrm{FO}[\Lambda_{2}]$ is quantifier-free. Now we can define $\varphi_{2}(\overline{x})=\exists y\,\psi_{2}(\overline{x},y)$. Note that $\varphi_{2}$ is also existential, and
$\textsf{true}(\varphi_{2},G,\ell_{2})=\textsf{true}(\varphi,G,\ell).$
We can finally apply Lemma 4.3 to compute a reduction $(p,\widehat{\Lambda},\widehat{\varphi},\widehat{\ell})$ of $(\Lambda_{2},\varphi_{2},\ell_{2})$.
Universal quantifier.
In this case, we merely rewrite $\varphi$ as:
$\varphi(\overline{x})=\neg\exists y\,\neg\psi(\overline{x},y).$
By induction hypothesis, an integer $p_{1}$, a set $\Lambda_{1}$, a $p_{1}$-lca-reduced formula $\psi_{1}(\overline{x},y)$, a $\Lambda_{1}$-labeling $\ell_{1}$ of $G$, and a $p_{1}$-skeleton $S_{1}$ of $G$ can be computed in $\mathcal{O}(D+\log n)$ rounds, such that
$\textsf{true}(\psi,G,\ell)=\textsf{true}(\psi_{1},S_{1},\ell_{1}).$
The negation of $\psi_{1}$ is still a $p_{1}$-lca-reduced formula. Let us apply Lemma 4.4 on $p_{1}$, $\Lambda_{1}$, $\neg\psi_{1}$, $G$, $S_{1}$, and $\ell_{1}$ to compute a set $\Lambda_{2}$, an existential formula $\psi_{2}$, and a $\Lambda_{2}$-labeling $\ell_{2}$ of $G$ in $\mathcal{O}(\log n)$ rounds, such that
$\textsf{true}(\neg\psi_{1},S_{1},\ell_{1})=\textsf{true}(\psi_{2},G,\ell_{2}).$
Now we can define $\rho(\overline{x})=\exists y\,\psi_{2}(\overline{x},y)$. Observe that this formula is existential, and it satisfies
$\textsf{true}(\varphi,G,\ell)=\textsf{true}(\neg\rho,G,\ell_{2}).$
We can eventually apply Lemma 4.3 to compute the reduction $(p,\widehat{\Lambda},\widehat{\rho},\widehat{\ell})$ of $(\Lambda_{2},\rho,\ell_{2})$ in $\mathcal{O}(D+\log n)$ additional rounds. In particular, $\widehat{\rho}$ is $p$-lca-reduced, as well as $\neg\widehat{\rho}$. Therefore $(p,\widehat{\Lambda},\neg\widehat{\rho},\widehat{\ell})$ is the reduction of $(\Lambda,\varphi,\ell)$.
We obtain that eliminating each quantifier of $\varphi$ can be done in $\mathcal{O}(D+\log n)$ rounds. ∎
4.4 Proof of Theorem 1.2
Let $\varphi$ be a formula in $\mathrm{FO}[\Lambda]$. We define the formula $\varphi^{\prime}(x)$ with one (dummy) free variable by
$\varphi^{\prime}(x)=\varphi\wedge(x=x).$
We do have that, for every labeled graph $(G,\ell)$, and for every node $v\in V(G)$,
$(G,\ell)\models\varphi^{\prime}(v)\iff(G,\ell)\models\varphi.$
By Lemma 4.5, $\varphi^{\prime}(x)$ can be reduced to a $p$-lca-reduced formula $\widehat{\varphi}(x)$ in $\mathcal{O}(D+\log n)$ rounds. Since $\widehat{\varphi}(x)$ is quantifier free, each node $v$ can check locally (without additional communication) whether it satisfies $\widehat{\varphi}$. By definition of the reduction, the truth of $\widehat{\varphi}$ is equivalent to that of $\varphi^{\prime}$, which completes the proof.∎
5 Counting and Optimization Problems
In this section, we prove Theorems 1.3 and 1.4. To ease the notations, let us adopt the following notations.
  • Let $\textsf{OPT}_{\varphi}(G,L)$ denote the optimum value for $\varphi$, i.e., the maximum value of $\sum_{i=1}^{k}\omega(v_{i})$ over all $k$-uples of vertices $(v_{1},\dots,v_{k})$ such that $G\models\varphi(v_{1},\dots,v_{k})$.
  • Let $\textsf{count}_{\varphi}(G,L)$ denote the number of $k$-uples satisfying the formula.
  • Given some $(p,f(p))$-treedepth coloring of $G$, and a set of colors $U$, let $\textsf{OPT}_{\varphi}(G[U],L)$ and $\textsf{count}_{\varphi}(G[U],L)$ be the same quantities as above, but assuming that the formula is restricted to the subgraph $G[U]$ of $G$ induced by the colors in $U$.
Let us prove the following lemma, which is restricted to quantifier-free formulas.
Lemma 5.1.
Let $\Lambda$ be a finite set. For every quantifier-free FO formula $\varphi(x_{1},\dots,x_{k})$ on $\Lambda$-labeled graphs with $k\geq 1$ free variables $x_{1},\dots,x_{k}$, and for every class of graphs $\mathcal{G}$ of bounded expansion, there exists a distributed algorithm that, for every $n$-node network $G=(V,E)\in\mathcal{G}$ of diameter $D$, and every $\ell:V\to 2^{\Lambda}$, performs global counting in ${O(D+\log n)}$ rounds under the CONGEST model. If the formula is local, then there exists an algorithm that performs local counting in $\mathcal{O}(\log n)$ rounds under the CONGEST model.
5.1 Proof of Lemma 5.1 and Theorem 1.4
We first consider Lemma 5.1, on counting. In this case, the formula $\varphi$ has exactly $k$ variables, namely $x_{1},\dots,x_{k}$. We now apply Proposition 2.9 to compute, in $\mathcal{O}(\log n)$ rounds, a $(k,f(k))$-treedepth coloring of $G$. Since formula $\varphi$ has exactly $k$ variables, a $k$-uple of vertices $(v_{1}\dots,v_{k})$ satisfies the formula in $(G,\ell)$ if and only if it satisfies it in some set $G[U]$. By definition, $\textsf{count}_{\varphi}(G[U],\ell)$ is the size of the set $\textsf{true}(\varphi,G[U],\ell)$ of $k$-uples $(v_{1},\dots,v_{k})$ such that $(G[U],\ell)\models\varphi(v_{1},\dots,v_{k})$. An easy but crucial observation is that, for any two sets of colors $U_{1},U_{2}$, we have that
$\textsf{true}(\varphi,G[U_{1}],\ell)\cap\textsf{true}(\varphi,G[U_{2}],\ell)=\textsf{true}(\varphi,G[U_{1}\cap U_{2}],\ell),$
because $\varphi(x_{1},\dots,x_{k})$ has no other variables than the free ones. The condition fails if one allows other variables, which explains that the technique does not extend directly to arbitrary FO formulas; this is the subject of Section 5.2. The equality above allows us to use the inclusion-exclusion principle in order to count the solutions of $\varphi$, assuming that we were able to count the partial solutions over subgraphs $G[U]$, where $U$ are sets of at most $k$ colors.
Lemma 5.2.
Let $\varphi(x_{1},\dots,x_{k})\in FO(\Lambda)$ be a quantifier-free formula. Let us consider a $(k,f(k))$-treedepth coloring of $(G,\ell)\in\mathcal{G}[\Lambda]$. Let $\kappa=\binom{f(k)}{k}$, and let $U_{1}\dots U_{\kappa}$ be all subsets of $[f(k)]$ of size $k$. Then
$\textsf{count}_{\varphi}(G,\ell)=\sum_{h=1}^{\kappa}(-1)^{h-1}\sum_{1\leq i_{1}\leq\dots\leq i_{h}\leq\kappa}\textsf{count}_{\varphi}(G[U_{i_{1}}\cap\dots\cap U_{i_{h}}],\ell).$
Proof.
Recall that for any $i\in[\kappa]$, $\textsf{count}_{\varphi}(G[U_{i}],\ell)$ is the size of the set $\textsf{true}(\varphi,G[U_{i}],\ell)$ of $k$-uples $(v_{1},\dots,v_{k})$ such that $(G[U_{i}],\ell)\models\varphi(v_{1},\dots,v_{k})$. Also, $\textsf{count}_{\varphi}(G[U_{i_{1}}\cap\dots\cap U_{i_{h}}],\ell)$ is the size of the set $\textsf{true}(\varphi,G[U_{i_{1}}\cap\dots\cap U_{i_{h}}],\ell)$, which is itself equal to $\textsf{true}(\varphi,G[U_{i_{1}}])\cap\dots\cap\textsf{true}(\varphi,G[U_{i_{h}}],\ell)$. Eventually, $\textsf{true}(\varphi,G,\ell)=\textsf{true}(\varphi,G[U_{1}],\ell)\cup\dots\cup\textsf{true}(\varphi,G[U_{\kappa}],\ell)$ and $\textsf{count}_{\varphi}(G,\ell)=|\textsf{true}(\varphi,G,\ell)|$, so the formula is a straightforward application of the inclusion-exclusion principle. ∎
Before explaining (in Lemma 5.5) how to compute the values $\textsf{count}_{\varphi}(G[U],\ell)$, we address optimization as stated in Theorem 1.4. Recall that, in general, the formula $\varphi$ may have other (quantified) variables than its free valiables. We first transform $\varphi(\overline{x})$ into an equivalent existential formula $\tilde{\varphi}(\overline{x})$ at the cost of enriching the set of labels, and the labeling of the input graph.
Lemma 5.3.
Given a formula $\varphi(\overline{x})\in FO[\Lambda]$, and a labeled graph $(G,\ell)\in\mathcal{G}[\Lambda]$, a tuple $(\tilde{\Lambda},\tilde{\varphi},\tilde{\ell})$ such that
  1. 1.
    $\tilde{\ell}$ is a $\tilde{\Lambda}$-labeling of $G$,
  2. 2.
    $\tilde{\varphi}(\overline{x})$ is an existential formula in $FO[\tilde{\Lambda}]$, i.e., $\tilde{\varphi}(\overline{x})=\exists\overline{y}\,\psi(\overline{x},\overline{y})$ where $\psi\in\tilde{\Lambda}$ is quantifier-free, and
  3. 3.
    $\textsf{true}(\varphi,G,\ell)=\textsf{true}(\tilde{\varphi},G,\tilde{\ell})$,
can be computed in $\mathcal{O}(D+\log n)$ rounds in CONGEST.
Proof.
Thanks to Lemma 4.5 (see also Definition 3.15), there exists some constant $p$ such that, for any $p$-skeleton $S$ of $G$ (obtained by Lemma 3.14), the tuple $(\Lambda,\varphi,\tau_{p},\ell)$ can be reduced to a tuple $(p,\widehat{\Lambda},\widehat{\varphi},\widehat{\ell})$ in $\mathcal{O}(D+\log n)$ rounds, where $\widehat{\varphi}(\overline{x})$ is a $p$-lca-reduced (thus quantifier-free) formula in $FO(\widehat{\Lambda})$, $\widehat{\ell}$ is a $\widehat{\Lambda}$ labeling of $S$, and $\textsf{true}(\varphi,G,\ell)=\textsf{true}(\widehat{\varphi},S,\widehat{\ell})$. Now, given the tuple $(p,\widehat{\Lambda},\widehat{\varphi},\widehat{\ell})$, Lemma 4.4 states that the desired tuple $(\tilde{\Lambda},\tilde{\varphi},\tilde{\ell})$ can be computed with no extra communication. ∎
Thanks to Lemma 5.3, we may now assume, w.l.o.g., that, in the proof of Theorem 1.4, the formula $\varphi(\overline{x})=\exists\overline{y}\psi(\overline{x},\overline{y})$ is given in an existential form, where $\psi\in FO[\Lambda]$ is quantifier-free.
Let $p=|\overline{x}|+|\overline{y}|$. We use again Proposition 2.9 to compute a $(p,f(p))$-treedepth coloring of $G$ in $\mathcal{O}(\log n)$ rounds. Since $\varphi$ has exactly $p$ variables, a $k$-uple of vertices $(v_{1}\dots,v_{k})$ satisfies the formula in $(G,\ell)$ if and only if it satisfies it in some subgraph $G[U]$, where $U$ can be viewed as obtained by “guessing” the colors of $(v_{1},\dots,v_{k})$ as well as the colors of the quantified variables $(y_{1},\dots,y_{p-k})=\overline{y}$. This is formally stated as follows, for further references.
Lemma 5.4.
Let $\varphi\in FO(\Lambda)$ be an existential formula, i.e., $\varphi(\overline{x})=\exists\overline{y},\psi(\overline{x},\overline{y})$, where $\psi\in FO[\Lambda]$ is quantifier-free. Let $p=|\overline{x}|+|\overline{y}|$, and let us assume a given a $(p,f(p))$ coloring of $(G,\ell)\in\mathcal{G}[\Lambda]$. We have
$\textsf{OPT}_{\varphi}(G,\ell)=\max_{U\subseteq[f(p)],\;|U|\leq p}\textsf{OPT}_{\varphi}(G[U],\ell).$
It remains to compute $\textsf{OPT}_{\varphi}(G[U],\ell)$ and $\textsf{count}_{\varphi}(G[U],\ell)$ for every set $U$ of at most $p$ colors, given some $(p,f(p))$-treedepth coloring of $G$. This is partially done in Proposition 2.7 thanks a previous result in [25] in which it is shown that $\textsf{OPT}_{\varphi}(G,\ell)$ and $\textsf{count}_{\varphi}(G,\ell)$ can be computed in a constant number of rounds in connected graphs of bounded treedepth. We however need to extend their techniques to fit our purposes, i.e., to deal with disconnected graphs of bounded treedepth (the non-necessarily connected subgraphs of $G$ induced by a set of colors $U$). For global counting, the cost of communication among different components adds an additive term $\mathcal{O}(D)$ to the round complexity.
Lemma 5.5.
Let $\mathcal{G}[\Lambda]$ be a class of labeled graphs of bounded expansion, let $\varphi(\overline{x})$ be a $FO(\Lambda)$ formula, and $p\in\mathbb{N}$. Let us assume given a $(p,f(p))$-treedepth coloring of $(G,\ell)\in\mathcal{G}[\Lambda]$, and let $U\subseteq[f(p)]$ be any set of at most $p$ colors. Then $\textsf{OPT}_{\varphi}(G[U],\ell)$ and $\textsf{count}_{\varphi}(G[U],\ell)$ can be computed in $\mathcal{O}(D)$ rounds in CONGEST. Moreover, if $\varphi$ is local, then local counting in $G[U]$ can be solved in $\mathcal{O}(1)$ rounds.
Proof.
Let us first consider the case of local counting, that is, we assume $\varphi$ is local. In this case, $\textsf{count}_{\varphi}(G[U],\ell)$ is equal to the sum of $\textsf{count}_{\varphi}$ over all connected components of $G[U]$, i.e.,
$\textsf{count}_{\varphi}(G[U],\ell)=\sum_{H\in\mbox{CC}(G[U])}\textsf{count}_{\varphi}(H,\ell),$
where $\mbox{CC}(G[U])$ denotes the set of connected components of $G[U]$. Each component $H\in CC(G[U])$ has treedepth at most $|U|$, and thus, by Proposition 2.7, $\textsf{count}_{\varphi}(H,\ell)$ can be computed in $\mathcal{O}(1)$ rounds. One arbitrary node $v_{H}\in H$ outputs $\nu(v_{H})=\textsf{count}_{\varphi}(H,\ell)$, whereas every other node $u$ outputs $\nu(u)=0$. Since $H$ has bounded treedepth and thus bounded diameter, we can choose $v_{H}$ to be for example the node in $H$ with maximal identifier. It follows that
$\sum_{v\in V}\nu(v)=\sum_{H\in\mbox{CC}(G[U])}\nu(v_{H})=\sum_{H\in\mbox{CC}(G[U])}\textsf{count}_{\varphi}(H,\ell)=\textsf{count}_{\varphi}(G[U],\ell)$
Let us now focus on counting and optimization for an arbitrary (non-necessarily local) formula.
To prove the lemma, we generalize the techniques of [25] for proving Proposition 2.7. We will only restate here the main features of the proof in [25], based on tree decomposition and treewidth-related techniques. Recall that the tree decomposition of a graph $H=(V,E)$ is a pair
$(T=(I,F),\{B_{i}\}_{i\in I})$
where $T$ is a tree, and, for every $i\in I$, $B_{i}$ is a subset of vertices of $V$, referred to as the bag of node $i$ in $T$, or merely “bag $i$”, such that
  • every vertex of $H$ is contained in at least one bag,
  • for every edge of $H$, at least one bag contains its two endpoints, and
  • for every vertex $v$ of $H$, the nodes of $T$ whose bags contain $v$ form a connected subgraph of $T$.
The width of the decomposition is the maximum size of a bag, minus one. The treewidth of $H$ is the minimum width over all its tree decompositions. Note that graphs of treedepth at most $p$ have treewidth at most $2^{p}$ [54]. The proof of Proposition 2.7 is based on three ingredients.
  1. 1.
    The first ingredient is a dynamic programming centralized algorithm for decision, optimization and counting solutions of a given formula $\varphi(x)\in FO(\Lambda)$, over labeled graphs $(H,\ell)$ of bounded treewidth. Here we will assume that a tree decomposition of constant width is part of the input, and moreover the tree is rooted. The distributed algorithm used in [25] is based on the centralized algorithm by Borie, Parker and Tovey [6]. The nodes of the tree decomposition $(T=(I,F),\{B_{i}\}_{i\in I})$ are processed bottom-up, from the leaves to the root of $T$. When processing node $i$, let $j_{1},\dots,j_{q}$ be its children in $T$. The essential observation due to [6] is that the dynamic programming tables at node $i$ are of constant size for decision, and of $\mathcal{O}(\log n)$ size for counting and optimization. Moreover, for computing the information (encoding partial solutions) at node $i$, one only needs to know the subgraphs $H[B_{i}]$ and $H[B_{j_{1}}],\dots,H[B_{j_{q}}]$ induced by the bags at node $i$ and its children, together with the information already computed at the children nodes. In particular, when $i$ is a leaf, $H[B_{i}]$ suffices for computing the dynamic programming table at node $i$.
  2. 2.
    The second ingredient of [25] is an implementation of the above centralized algorithm in the CONGEST model, in a number of rounds equal to the depth of the tree decomposition $T$, under the condition that one can associate to each node $i$ of $T$ a vertex $h(i)$ of $H$, such that vertices $h(i)$ satisfy several properties in order to “simulate the behavior” of node $i$ in the dynamic programming. Vertex $h(i)$ must know the depth of tree $T$ and the depth of node $i$ in $T$, the bag $B_{i}$ and the subgraph $H[B_{i}]$, and moreover $h(i)$ knows the node $h(p_{T}(i))$ associated to the parent $p_{T}(i)$ of $i$ in $T$, and vertices $h(i),h(p_{T}(i))$ are equal or adjacent in $H$. Informally, the distributed algorithm implements the centralized one by making vertex $h(i)$ to help it computing the right moment (i.e., at the round corresponding to the depth of $T$ minus the depth of $i$ in $T$, plus one) the dynamic programming table at node $i$ is processed, and the result transmitted to $h(p_{T}(i))$. Observe that $h(i)$ has collected all the information from the children $j_{1},\dots,j_{q}$ of $i$, thanks to the fact that $h(j_{1}),\dots,h(j_{q})$ have already computed and communicated it at the previous round. (To be complete, in [25], one actually has $h(i)=i$ and $T$ is a subgraph of $H$, but our generalization is a simple observation.) This explains that the number of rounds of the distributed algorithm is the depth of the tree decomposition $T$.
  3. 3.
    The third ingredient in [25] is an algorithm computing the desired tree decomposition $(T=(I,F),\{B_{i}\}_{i\in I})$ of graph $H$ of treedepth at most $p$ in a constant number of rounds, together with the information requested by the previous item. The width of the tree-decomposition is at most $p$, and the depth of $T$ is at most $2^{p}$.
Let us now come back to our framework, that is when $H=G[U]$ is the subgraph of $G$ induced by some set $U$ of at most $p$ colors of some $(p,f(p))$-treedepth decomposition of $G$. In particular $H=G[U]$ has treedepth at most $p$. Our goal is to replace the third item above by a distributed algorithm running in $\mathcal{O}(D)$ rounds in CONGEST for computing a suitable tree decomposition of $G[U]$. We compute a $p$-skeleton using Lemma 3.14. Note that this can be done in $\mathcal{O}(2^{2p})$ rounds, hence in a constant number of rounds whenever the coloring is provided. Let $F^{U}$ be the resulting decomposition forest spanning $G[U]$, of depth at most $2^{d}$. Recall that each node knows its parent and depth in the forest.
Generalizing [25], we transform the decomposition forest $F^{U}$ into a tree decomposition of $G[U]$. This tree decomposition $(T=(I,F),\{B_{i}\}_{i\in I})$ has to be computed in time $\mathcal{O}(D)$ and to satisfy all the conditions of the second ingredient above. The only difference is that, for each node $i$ of $T$, its “leader” $h(i)$ has to be a vertex of $G$ (not necessarily of $G[U]$), and, for each pair of nodes $i,j$ of $T$ such that $j=p_{T}(i)$, vertices $h(i)$ and $h(j)$ must be adjacent in $G$. In our construction, there may be up to two different nodes of $T$ having the same leader in $G$.
Let us choose an arbitrary vertex $r$ of $G$, and compute a breadth-first search tree of $G$. From this rooted tree, remove every node $v$ that does not have a descendant which is a root in $F^{U}$. Note that the construction of this tree can be performed in time $\mathcal{O}(D)$ rounds. We call $TS$ the resulting tree, rooted at $r$. $TS$ can be viewed as a Steiner tree in $G$ spanning $r$ and all the vertices labeled $\textsf{root}^{U}$. The depth of $TS$ is, by construction, at most $D$, the diameter of $G$.
From now on, let us focus on $TS$ alone, independently from the the fact that it is a subgraph of $G$. Let us then construct a “virtual” tree $T$ by augmenting $TS$ a follows. For every vertex labeled $\textsf{root}^{U}$ in $TS$, we root the corresponding tree $T^{U}$ to that vertex. So, such a vertex has two types of children in $T$, those originally in $TS$, and those in the tree $T^{U}$ rooted at that vertex. Note that two different vertices $i$ and $j$ in $T$ may correspond to the same node $v$ in $G$, as node $i$ may be $v$ as a node in $TS$, and $j$ may be $v$ again, but as a node of $T^{U}$. For every node $i$ of $T$, let $h(i)$ be the corresponding vertex of $G$. Note that, by a simple distributed algorithm working in $\mathcal{O}(D)$ rounds, each node $h(i)$ of $G$ can learn the depth of $T$, the depth of $i$ in $T$, and the identifier of $h(p_{T}(i))$. The depth of $T$ is at most $D+2^{p}$ because the forest $F$ has depth at most $2^{p}$. By construction, $h(i)$ and $h(p_{T}(i))$ are either adjacent in $G$ or equal, whenever $i$ is labeled $\textsf{root}^{U}$.
It remains to construct the bag $B_{i}$ for each node $i$ of $T$, and to ensure that the node $h(i)$ of $G$ knows the subgraph induced by $B_{i}$ in $H=G[U]$.
  • For every node $i$ of $T$ belonging to $TS$, we set $B_{i}=\emptyset$.
  • For every node $i$ belonging to $F^{U}$, we act as in  [25], i.e., we set $h(i)=i$, and $B_{i}$ is the set of ancestors of $i$ in $F^{U}$. As observed in [25], each node $i$ can learn the subgraph induced by bag $B_{i}$ in a constant number of rounds. Moreover, for each tree $T^{U}$ of the forest $F^{U}$, the restriction of the decomposition $(T=(I,F),\{B_{i}\}_{i\in F})$ to $T^{U}$ is a tree decomposition of the subgraph of $G$ induced by the vertices of $T^{U}$, of width at most $2^{p}$.
It is a matter of exercise to check that, by attaching to $TS$ the tree-decompositions corresponding to the trees $T^{U}$ in the forest $F^{U}$, the resulting tree is a tree decomposition of $G[U]$ with width at most $2^{p}$. Therefore we can apply the same dynamic programming scheme as above for computing values $\textsf{OPT}_{\varphi}(G[U],\ell)$ and $\textsf{count}_{\varphi}(G[U],\ell)$ in $D+2^{p}$ rounds. Together with the construction of the tree decomposition, the overall algorithm takes $\mathcal{O}(D)$ rounds. ∎
By plugging the result of Lemma 5.5 into Lemma 5.4 and Lemma 5.2 over all sets $U$ of at most $p$ colors for a suitable $p$ depending on $\varphi$ and a $(p,f(p))$-treedepth coloring of $G$, we achieve the proof of Lemma 5.1 and Theorem 1.4. Note that in each case, we need to build a $(p,f(p))$-treedepth coloring, which can be done in $\mathcal{O}(\log n)$ CONGEST rounds by Proposition 2.9. ∎
5.2 Proof of Theorem 1.3
Lemma 5.1 shows how to solve the counting problem for quantifier-free FO formulas on bounded expansion classes. In order to prove the more general Theorem 1.3, we apply the quantifier elimination techniques developed in Sections 3 and 4. Lemma 4.5 transforms a general FO formula into a $p$-lca-reduced formula. However, Lemma 5.1 does not handle such formulas, as it only deals with usual adjacency and equality predicates. In order to solve that issue, it is tempting to use Lemma 4.4, which replaces lca predicates by FO formulas. However, this would result in an existential formula rather than a quantifier free formula as needed for applying Lemma 5.1. The main idea of the proof of Theorem 1.3 is to maintain that the quantified variables introduced by Lemma 4.4 are quantified uniquely, in the following sense.
Definition 5.6.
Let $p\in\mathbb{N}$, let $\Lambda$ a set of labels, and let $\varphi(\overline{x})\in\mathrm{FO}[\Lambda]$ be a $p$-lca-reduced formula. $\varphi$ is lca-freeable on $\mathcal{G}$ if there exists $\widehat{\Lambda}$, and a quantifier free formula $\psi(\overline{x},\overline{y})\in\mathrm{FO}[\widehat{\Lambda}]$ such that, for every $G=(V,E)\in\mathcal{G}$, for every $\Lambda$-labeled $p$-skeleton $(S,\ell)$ of $G$, there exists a $\widehat{\Lambda}$-labeling $\widehat{\ell}$ of $G$ satisfying
  1. 1.
    $\textsf{true}(\varphi,S,\ell)=\textsf{true}\Big{(}\big{(}\exists\overline{y}\,\psi(\overline{x},\overline{y})\big{)},G,\widehat{\ell}\Big{)}$, and
  2. 2.
    for every $\overline{v}\subseteq V$ satisfying $(S,\ell)\models\varphi(\overline{v})$, there exists a unique $\overline{w}\subseteq V$ such that $(G,\widehat{\ell})\models\psi(\overline{v},\overline{w})$.
The following lemma outlines the usefulness of lca-freeability for counting.
Lemma 5.7.
Let $p\in\mathbb{N}$, and let $\varphi(\overline{x})\in\mathrm{FO}[\Lambda]$ be a $p$-lca-reduced, and lca-freeable formula. For every $\widehat{\Lambda},\psi(\overline{x},\overline{y}),G,S,\ell,\widehat{\ell}$ as in Definition 5.6,
$\textsf{count}_{\varphi}(S,\ell)=\textsf{count}_{\psi}(G,\widehat{\ell}).$
Proof.
Definition 5.6 specifies the existence of a one-to-one correspondence between $\textsf{true}(\varphi,S,\ell)$ and $\textsf{true}(\psi,G,\widehat{\ell})$. These two set have thus the same cardinality. ∎
The following lemma is an analog of Lemma 4.4.
Lemma 5.8.
Every formula $\varphi(y,z)=\textsf{lca}_{i}^{U}(y,z)$ is lca-freeable on $\mathcal{G}$. Moreover, assuming $(S,\ell)$ is given as input, every node $u$ can compute $\widehat{\Lambda},\psi,\widehat{\ell}(u)$ without any communication.
Proof.
The proof follows the same guidelines as the one of Lemma 4.4. Recall labels $\textsf{color}_{k}$ were added for each color $k$, and the label $\textsf{depth}^{U}_{j}$ was added to mark the depth of a node in each forest $F^{U}$. Each lca predicates was then replaced by an existential formula. The following formula is very similar, except that we removed the quantifiers on $y_{0},\dots,y_{d-1},z_{0},\dots,z_{d-1}$, which are now free variables, and we added another constraint, specified on the last line of the formula below.
$\psi(y,z,y_{0},\dots,y_{d-1},$ $z_{0},\dots,z_{d-1})=\bigvee_{(a,b)\in[\max(0,i),d-1]^{2}}\Biggl{(}(y_{a}=y)\wedge(z_{b}=z)$
$\wedge\Big{(}\bigwedge_{s\in[0,i]}y_{s}=z_{s}\Big{)}\wedge\Big{(}\bigwedge_{s\in[i+1,\min(a,b)]}y_{s}\neq z_{s}\Big{)}\wedge\textsf{depth}_{0}^{U}(y_{0})\wedge\textsf{depth}_{0}^{U}(z_{0})$
$\wedge\Big{(}\bigwedge_{s\in[1,a]}\textsf{depth}_{s}^{U}(y_{s})\wedge\textsf{adj}(y_{s-1},y_{s})\Big{)}\wedge\Big{(}\bigwedge_{s\in[1,b]}\textsf{depth}_{s}^{U}(z_{s})\wedge\textsf{adj}(z_{s-1},z_{s})\Big{)}$
$\wedge\Big{(}\bigwedge_{s\in[a+1,d-1]}y_{s}=y\Big{)}\wedge\Big{(}\bigwedge_{s\in[b+1,d-1]}z_{s}=z\Big{)}\Biggr{)}.$
Recall that $a$ (resp., $b$) represents here the depth of $y$ (resp., $z$) in $F^{U}$, and $y_{0},\dots,y_{a}$ (resp., $z_{0},\dots,z_{b}$) represent the path from the root of $y$’s tree to $y$ (resp., from the root of $z$’s tree to $z$). Since, in general, variables $y_{a+1},\dots,y_{d-1},z_{b+1},\dots,z_{d-1}$ can take any value, the (new) last line of the formula above is needed to enforce their uniqueness. For the same reasons as in Lemma 4.4, $\psi$ satisfies
$\textsf{true}(\varphi,S,\ell)=\textsf{true}(\exists\overline{y},\exists\overline{z},\psi(y,z,\overline{y},\overline{z}),G,\widehat{\ell}).$
Let us now prove that the quantified variables are unique. Fix $(y,z)\in\textsf{true}(\varphi,S,\ell)$ and let $\overline{y},\overline{z}$ such that $(G,\widehat{\ell})\models\psi(y,z,\overline{y},\overline{z})$. For some $(a,b)\in[i,d-1]^{2}$, the inside of the disjunction of $\psi$ is true. This implies that:
  • $y$ (resp. $z$) has depth $a$ (resp. $b$),
  • for $s\leq a$ (resp. $s\leq b$), $y_{s}$ (resp. $z_{s}$) is the (unique) ancestor of $y$ (resp. $z$) at depth $s$, and
  • for $s>a$ (resp. $s>b$), $y_{s}=y$ (resp. $z_{s}=z$), thus it is also unique.
This completes the proof. ∎
Lemma 5.9.
A conjunction of lca-freeable formulas is lca-freeable.
Proof.
Let $\varphi_{1}(\overline{x}),\varphi_{2}(\overline{x})$ be two lca-freeable formulas and $\varphi(\overline{x})=\varphi_{1}(\overline{x})\wedge\varphi_{2}(\overline{x})$. Let $\widehat{\Lambda}_{1},\widehat{\Lambda}_{2}$, $\psi_{1}(\overline{x},\overline{y}),\psi_{2}(\overline{x},\overline{z})$ as in Definition 5.6. We set $\widehat{\Lambda}=\widehat{\Lambda}_{1}\cup\widehat{\Lambda}_{2}$, where we assume, w.l.o.g., that the sets of labels $\widehat{\Lambda}_{1},\widehat{\Lambda}_{2}$ are disjoint, and we set $\psi(\overline{x},\overline{y},\overline{z})=\psi_{1}(\overline{x},\overline{y})\wedge\psi_{2}(\overline{x},\overline{z})$. For $G,S,\ell_{1},\ell_{2},\widehat{\ell}_{1},\widehat{\ell}_{2}$ as in Definition 5.6, we define ^$\widehat{\ell}(u)=\widehat{\ell}_{1}(u)\cup\widehat{\ell}_{2}(u)$. Let $\overline{v}\in\textsf{true}(\varphi,S,\ell)$. Since $\psi_{1}$ (resp. $\psi_{2}$) is lca-freeable, there exists a unique $\overline{w}$ (resp. $\overline{u}$) such that $(\overline{v},\overline{w})\in\textsf{true}(\psi_{1},G,\widehat{\ell}_{1})$ (resp. $(\overline{v},\overline{u})\in\textsf{true}(\psi_{2},G,\widehat{\ell}_{2})$). Then, $(G,\widehat{\ell})\models\psi(\overline{v},\overline{w},\overline{u})$. Conversely, let $\overline{w^{\prime}},\overline{u^{\prime}}$ such that $(G,\widehat{\ell})\models\psi(\overline{v},\overline{w^{\prime}},\overline{u^{\prime}})$. Then, $(G,\widehat{\ell})\models\psi_{1}(\overline{v},\overline{w^{\prime}}$ (resp. $\psi_{2}(\overline{v},\overline{u^{\prime}})$), which implies $\overline{w^{\prime}}=\overline{w}$ (resp. $\overline{u^{\prime}}=\overline{u}$). ∎
At this stage, we can now prove the global counting part of Theorem 1.3.
Proof of Theorem 1.3 (global counting).
Let $\Lambda$ be a (finite) set of labels and $\varphi(\overline{x})\in\mathrm{FO}[\Lambda]$. By Lemma 4.5, there exists $p\in\mathbb{N}$, $\Lambda_{1}$ and $\varphi_{1}(\overline{x})\in\mathrm{FO}[\Lambda_{1}]$ $p$-lca-reduced such that, for every $(G,\ell)\in\mathcal{G}[\Lambda]$, and for every $S$ $p$-skeleton of $G$, there exists a $\Lambda_{1}$-labeling $\ell_{1}$ of $S$ such that $\textsf{true}(\varphi,G,\ell)=\textsf{true}(\varphi_{1},S,\ell_{1})$ (and thus $\textsf{count}_{\varphi}(G,\ell)=\textsf{count}_{\varphi_{1}}(S,\ell_{1})$. We compute a $p$-skeleton using Lemma 3.14, and the corresponding labeling $\ell_{1}$, in $\mathcal{O}(D+\log n)$ rounds. Let us write $\varphi_{1}$ in disjunctive normal form:
$\varphi_{1}(\overline{x})=\bigvee_{j=1}^{k}\psi_{j}(\overline{x})$
where $\psi_{j}$ is a conjunction of $\textsf{lca}^{U}_{i}$ predicates, label predicates, color predicates, or their negations. As in Lemma 3.18, note that $\neg\textsf{lca}^{U}_{i}(x,y)$ is equivalent to:
$\neg\textsf{color}_{U}(x)\vee\neg\textsf{color}_{U}(y)\vee\bigvee_{j\in[-1,d]\setminus\{i\}}\textsf{lca}^{U}_{j}(x,y).$
We may thus assume that $\varphi_{1}$ does not contain any negated lca predicate. We apply the inclusion/exclusion principle to $\varphi_{1}$, that is,
$\textsf{count}_{\varphi_{1}}(S,\ell_{1})=\sum_{h=1}^{k}(-1)^{h-1}\sum_{1\leq j_{1}<\dots<j_{h}\leq k}\textsf{count}_{\bigwedge_{m=1}^{h}\psi_{j_{m}}}(S,\ell_{1}).$
It only remains to compute the values of
$\textsf{count}_{\bigwedge_{m=1}^{h}\psi_{j_{m}}}(S,\ell_{1}),$
for some fixed indexes $j_{1},\dots,j_{h}$. Since each $\psi_{j_{m}}$ are conjunctions of $\textsf{lca}^{U}$-predicates and (possibly negated) label and color predicates, it is also true of $\bigwedge_{m=1}^{h}\psi_{j_{m}}$. By Lemmas 5.8 and 5.9, this formula is lca-freeable. By Lemma 5.7 there exists $\widehat{\Lambda}$, $\widehat{\psi}\in\mathrm{FO}[\Lambda]$ quantifier-free, and $\widehat{\ell}$ such that
$\textsf{count}_{\bigwedge_{m=1}^{h}\psi_{j_{m}}}(S,\ell_{1})=\textsf{count}_{\widehat{\psi}}(G,\widehat{\ell}).$
Moreover, $\widehat{\Lambda},\widehat{\psi},\widehat{\ell}(u)$ can be computed by any node $u$ without communication (cf. Lemmas 5.8 and 5.9, and their proofs). Finally, since $\widehat{\psi}$ is quantifier free, it only remains to apply Lemma 5.1 to complete the proof. ∎
In order to prove our result on local counting, we need to adapt the notion of lca-freeable formulas so that it preserves locality.
Definition 5.10.
Let $p\in\mathbb{N}$, let $\Lambda$ be a (finite) set of labels, and let $\varphi(\overline{x})\in\mathrm{FO}[\Lambda]$ be a local $p$-lca-reduced formula. $\varphi$ is locally lca-freeable on $\mathcal{G}$ if the following two conditions hold:
  • $\varphi$ is lca-freeable, and
  • the quantifier-free formula $\psi$ from Definition 5.6 can be chosen to be local.
The following lemma is the “locality preserving version” of Lemmas 5.8 and 5.9.
Lemma 5.11.
Let $\varphi(\overline{x})\in\mathrm{FO}[\Lambda]$ be a $p$-lca-reduced formula that is supposed to be a conjunction of $\textsf{lca}^{U}$ predicates, and (possibly negated) label and color predicates. If $\varphi$ is local then it is locally lca-freeable. Moreover, assuming a $p$-skeleton $(S,\ell)$ is given as input, each node $u$ can compute $\widehat{\Lambda},\psi,\widehat{\ell}(u)$ matching the requirement of Definition 5.6 without any communication.
Before proving Lemma 5.11, we first prove a technical lemma.
Lemma 5.12.
Let $\varphi(\overline{x})$ be a local $p$-lca-reduced formula that is assumed to be a conjunction of $\textsf{lca}^{U}$ predicates, and (possibly negated) label and color predicates. Let $k=|x|$, and let $H(\varphi)$ be the graph with vertex set equal to the set of variables $x_{1},\dots,x_{k}$, and where there is an edge between $x_{i}$ and $x_{j}$ if and only if the formula $\varphi$ contains a predicate of the form $\textsf{lca}^{U}_{m}(x_{i},x_{j})$ with $m\geq 0$. If there exists a labeled skeleton $(S,\ell)$, and a set of vertices $\overline{v}$ such that $(S,\ell)\models\varphi(\overline{v})$, then, $H(\varphi)$ is connected.
Proof.
Let us fix a formula $\varphi$, a labeled skeleton $(S,\ell)$, and a set of vertices $\overline{v}$ as in the statement of the lemma. We proceed by contradiction. Let us assume that $H(\varphi)$ is not connected. Up to a reordering of the variables $x_{2},\dots,x_{k}$, we may assume that $x_{1},\dots,x_{s}$ form a single connected component, $1\leq s<k$, and that the vertices $x_{s+1},\dots,x_{k}$ are in other connected components. Let us consider the labeled skeleton $(S^{\prime},\ell^{\prime})$ consisting of two connected components $S^{1}$ and $S^{2}$, each of which being a copy of $(S,\ell)$. For any vertex $u\in S$, let us denote by $u^{i}$ the copy of $u$ in $S^{i}$. Let $\overline{w}$ where, for every $i=1,\dots,k$,
$w_{i}=\begin{cases}v^{1}_{i}&\textrm{if }i\leq s\\ v^{2}_{i}&\textrm{if }i>s\end{cases}$
We have that $(S^{\prime},\ell^{\prime})\models\varphi(\overline{w})$. However $w_{s+1},\dots,w_{k}$ are not in the connected component of $w_{1}$. In particular, they are not at bounded distance from $w_{1}$. This contradicts the fact that $\varphi$ is local (see Definition 2.2). ∎
When $H(\varphi)$ is connected, we say that $\varphi$ is lca-connected.
Proof of Lemma 5.11.
Since $\varphi$ is local, we get from Lemma 5.12 that it is either lca-connected, or always false. Let us assume that it is not lca-connected. Then one can trivially output the formula $\bot$ that is always false. Thus, we now assume that $\varphi$ is lca-connected. We then merely adapt the proofs of Lemmas 5.8 and 5.9. Let us consider a labeled $p$-skeleton $(S,\ell)$. As in the proof of Lemma 5.8, we add new labels $\textsf{depth}^{U}_{i}$, and $\textsf{color}_{k}$, and we denote by $\widehat{\ell}$ the resulting labeling. We then replace each predicate $\textsf{lca}^{U}_{i}(y,z)$ of $\varphi$ by
$\bigvee_{(a,b)\in[\max(0,i),d-1]^{2}}$ $\Biggl{(}(y_{a}=y)\wedge(z_{b}=z)\wedge\Big{(}\bigwedge_{s\in[0,i]}y_{s}=z_{s}\Big{)}\wedge\Big{(}\bigwedge_{s\in[i+1,\min(a,b)]}y_{s}\neq z_{s}\Big{)}$
$\wedge\textsf{depth}_{0}^{U}(y_{0})\wedge\textsf{depth}_{0}^{U}(z_{0})$
$\wedge\Big{(}\bigwedge_{s\in[1,a]}\textsf{depth}_{s}^{U}(y_{s})\wedge\textsf{adj}(y_{s-1},y_{s})\Big{)}\wedge\Big{(}\bigwedge_{s\in[1,b]}\textsf{depth}_{s}^{U}(z_{s})\wedge\textsf{adj}(z_{s-1},z_{s})\Big{)}$
$\wedge\Big{(}\bigwedge_{s\in[a+1,d-1]}y_{s}=y\Big{)}\wedge\Big{(}\bigwedge_{s\in[b+1,d-1]}z_{s}=z\Big{)}\Biggr{)},$
where $y_{1},\dots,y_{d},z_{1},\dots,z_{d}$ are new free variables (which are unique, see the proof of Lemma 5.8). Note that if the formula above is satisfied, then each newly introduced variable $y_{i}$ (resp. $z_{i}$) is at distance at most $d$ from $y$ (resp. $z$). Moreover, if $i\geq 0$, then $y$ and $z$ are at distance at most $2d$, as they are both at distance at most $d$ from $y_{0}=z_{0}$. Let us denote by $\psi(\overline{x},\overline{t})$ the formula obtained by replacing each $\textsf{lca}^{U}$-predicate by the formula above ($\overline{t}$ is the collection of newly introduced variables). By the same argument as in Lemma 5.8, we have
$\textsf{true}(\varphi,S,\ell)=\textsf{true}(\exists\overline{t}\;\psi(\overline{x},\overline{t}),G,\widehat{\ell}),$
and
$\textsf{count}_{\varphi}(S,\ell)=\textsf{count}_{\psi}(G,\widehat{\ell}).$
It only remains to show that $\psi$ is local. Since it is quantifier-free, this is equivalent to showing that, for some $r^{\prime}$, every labeled graph $(G^{\prime},\ell^{\prime})$ satisfies
$(\overline{v},\overline{w})\in\textsf{true}(\psi,G^{\prime},\ell^{\prime})\implies\overline{v},\overline{w}\subseteq B_{G^{\prime}}(v_{1},r^{\prime}).$
Since, are underlined above, each newly introduced variable in $t_{i}\in\bar{t}$ is at distance at most $d$ from at least one of the variables $x_{j}$, it is sufficient to show that for some $r^{\prime}$
$(\overline{v},\overline{w})\in\textsf{true}(\psi,G^{\prime},\ell^{\prime})\implies\overline{v}\subseteq B_{G^{\prime}}(v_{1},r^{\prime}),$
as $\overline{v}\subseteq B_{G^{\prime}}(v_{1},r^{\prime})$ implies that $\overline{w}\subseteq B_{G^{\prime}}(v_{1},r^{\prime}+d)$. Recall that $\varphi(\overline{x})$ is lca-connected, and so let $H(\varphi)$ be defined as in Lemma 5.12. For each index $1\leq j\leq|x|$, since $H(\varphi)$ is connected there is a path $x_{i_{1}},\dots,x_{i_{k}}$ from $x_{1}$ to $x_{j}$ in $H(\varphi)$. Since $H(\varphi)$ has $|x|$ nodes, we may assume this path has length at most $|x|-1$. The formula $\psi$ guarantees that, for every $l<k$, $x_{i_{l}}$ and $x_{i_{l+1}}$ are at distance at most $2d$ in $G^{\prime}$. Finally, $x_{1}$ and $x_{j}$ are at distance at most $2d(|x|-1)$. This concludes the proof that $\varphi$ is local. ∎
We are now ready to establish Theorem 1.3.
Proof of Theorem 1.3 (local counting).
We follow the same guidelines as in the proof of the global counting result, but we make sure that locality is preserved throughout the process. Let us consider a labeled graph $(G,\ell)$. Let us apply the locality preserving elimination technique (cf. Lemma 3.19) for reducing $\varphi$ to a local $p$-lca-reduced formula $\varphi_{1}$. We then construct a $p$-skeleton $S$ (cf. Lemma 3.14), and a new labeling $\ell_{1}$, in $\mathcal{O}(\log n)$ rounds. Let us rewrite $\varphi_{1}$ in disjunctive normal form:
$\varphi_{1}(\overline{x})=\bigvee_{j=1}^{k}\psi_{j}(\overline{x})$
where $\psi_{j}$ is a conjunction of $\textsf{lca}^{U}$-predicates, and (possibly negated) label and color predicates — recall that we can assume that there are no negated $\textsf{lca}^{U}$-predicates. The formula $\varphi_{1}$ is quantifier-free so, thanks to Lemma 2.6, each term of the disjunction is local. Let us apply the inclusion/exclusion principle to $\varphi_{1}$, i.e.,
$\textsf{count}_{\varphi_{1}}(S,\ell_{1})=\sum_{h=1}^{k}(-1)^{h-1}\sum_{1\leq j_{1}<\dots<j_{h}\leq k}\textsf{count}_{\bigwedge_{m=1}^{h}\psi_{j_{m}}}(S,\ell_{1}).$
It thus remains solely to perform the local counting of
$\textsf{count}_{\bigwedge_{m=1}^{h}\psi_{j_{m}}}(S,\ell_{1})$
[…]
$\textsf{count}_{\bigwedge_{m=1}^{h}\psi_{j_{m}}}(S,\ell_{1})=\textsf{count}_{\widehat{\psi}}(G,\widehat{\ell}).$
Since $\widehat{\psi}$ is local and quantifier-free, it suffices to apply (the local case of) Lemma 5.1 for completing the proof. ∎
6 Distributed Certification
In this section we prove that every graph property expressible in FO can be certified with certificates on $\mathcal{O}(\log n)$ bits in graphs of bounded expansion. For establishing Theorem 1.5, we revisit the logical structures introduced and analyzed in the previous section, by mainly focusing on the points where theses structures must be certified instead of being computed.
6.1 Certifying Reduction of Trees with Bounded Depth
We begin with a basic primitive, that we use to count values, to identify connected components, and for several other applications.
Proposition 6.1 ([48]).
Let us suppose that each node $u$ in a graph $G=(V,E)$ is given as an input a triple $(\textsf{root}(u),\textsf{parent}(u),\textsf{depth}(u))$, where $\textsf{root}(u)$ and $\textsf{parent}(u)$ are node identifiers, and $\textsf{depth}(u)$ is a positive integer. Let us consider the algorithm where each node $u$ checks:
  • For each $v\in N(u)$, $\textsf{root}(v)=\textsf{root}(u)=\rho$.
  • If $\textsf{id}(u)=\rho$ then $\textsf{parent}(u)=\textsf{id}(u)$ and $\textsf{depth}(u)=0$.
  • If $\textsf{id}(u)\neq\rho$, then there exists $v\in N(u)$ s.t. $\textsf{id}(v)=\textsf{parent}(u)$ and $\textsf{depth}(v)=\textsf{depth}(u)-1$.
If all checks are passed at every node, then there exists a spanning tree $T$ of $G$ rooted at $\rho$ such that the function parent defines the parent relation in $T$, and $\textsf{depth}(u)$ is the depth of $u$ in $T$.
In the above proposition, if all checks are passed at every node, then the set
$\{(\textsf{root}(u),\textsf{parent}(u),\textsf{depth}(u))\mid u\in V\}$
is called the encoding of the tree $T$. Given a graph $G$ and a forest $F\subseteq G$ with components $T_{1},\dots,T_{k}$, we say that the nodes of $G$ are given the encoding of $F$ if, for every $i\in[k]$, each node $u\in T_{i}$ is given the encoding of $T_{i}$.
Lemma 6.2.
Let $\mathcal{A}$ be an algorithm such that, in every $n$-node network $G=(V,E)$, every node $u\in V$ is given an $\mathcal{O}(\log n)$-bit input $\textsf{In}(u)$, and must produce an $\mathcal{O}(\log n)$-bit output $\textsf{Out}(u)$. Let $R(n)$ be the round-complexity of $\mathcal{A}$ in the BROADCAST CONGEST model. Then, there is a certification scheme using certificates on $\mathcal{O}(R(n)\log n)$ bits that certifies that a pair of functions $f:V\to\{0,1\}^{*}$ and $g:V\to\{0,1\}^{*}$ form an input-output pair for $\mathcal{A}$ (i.e., if, for every node $u\in V$, $\textsf{In}(u)=f(u)$, then, for every node $u\in V$, $\textsf{Out}(u)=g(u)$).
Proof.
On legal instances $(f,g)$ for network $G=(V,E)$, the prover provides each vertex $u\in V$ with the sequence $M(u)=(m_{1}(u),\dots,m_{R(n)}(u))$ where $m_{i}(u)$ is the message that $u$ broadcasts to its neighbors on input $f(u)$ during the $i$-th round of $\mathcal{A}$.
The verification algorithm proceeds as follows. Every node collects the certificates of its neighbors, and simulates the $R(n)$ rounds of $\mathcal{A}$. That is, $u$ checks first whether $m_{1}(u)$ is indeed the message that it would send in the first round of $\mathcal{A}$. For every $i\in\{2,\dots,R(n)\}$, node $u$ assumes that the messages $m_{1}(v),\dots,m_{i-1}(v)$ of each of its neighbors $v\in N(u)$ were correct, and checks whether $m_{i}(u)$ is indeed the message that it would broadcast at round $i$. If some of these tests are not passed then $u$ rejects. If all tests are passed, then $u$ finally checks whether $g(u)$ is the output that $u$ would produce by executing $\mathcal{A}$, and accepts or rejects accordingly.
The completeness and soundness of the algorithm follow by construction. ∎
Let us now prove a result analogous to Lemma 4.1, but for distributed certification.
Lemma 6.3.
Let $(G,\ell)$ be a connected $\Lambda$-labeled graph, and let $F\in\mathcal{F}_{d}$ be a subgraph of $G$. Let $\varphi(\overline{x})=\exists\overline{y}\;\zeta(\overline{x},\overline{y})$ be a formula in $\mathrm{FO}[\Lambda]$, where $\zeta$ is an lca-reduced formula. Let us assume that every node $u$ of $G$ is provided with the following values:
  • Two sets $\Lambda$ and $\widehat{\Lambda}$;
  • The formula $\varphi$, and a lca-reduced formula $\widehat{\varphi}$;
  • A set of labels $\ell(u)\subseteq\Lambda$, and a set of labels ^$\widehat{\ell}(u)\subseteq\hat{\Lambda}$;
  • The encoding $(\textsf{root}(u),\textsf{parent}(u),\textsf{depth}(u))$ of $F$;
  • The encoding $(\textsf{root}^{T}(u),\textsf{parent}^{T}(u),\textsf{depth}^{T}(u))$ of a spanning tree $T$ of $G$.
There is a certification scheme using $\mathcal{O}(\log n)$-bit certificates that certifies that $\textsf{true}(\varphi,F,\ell)=\textsf{true}(\widehat{\varphi},F,\tilde{\ell})$.
Proof.
Let $k=|\overline{x}|$. As in the proof of Lemma 3.7, let us assume w.l.o.g. that $k\geq 1$ (as otherwise we can create a dummy free variable $x$ for $\varphi$). Moreover, let us assume that $\widehat{\Lambda}$ and $\widehat{\varphi}$ are the set and the formula defined in the proof of Lemma 3.7. Since $\widehat{\Lambda}$ and $\widehat{\varphi}$ can be computed by every node without any communication, each node can reject in case they are not appropriately defined. Similarly, we can assume that $\ell$ and $\hat{\ell}$ are, respectively, a $\Lambda$-labeling and a $\hat{\Lambda}$-labeling. Thanks to Proposition 6.1, $\mathcal{O}(\log n)$-bits certificates are sufficient to certify that the encodings of $F$ and $T$ are correctly. The remaining of the proof consists in checking that $\widehat{\ell}$ is correctly set as specified the proof of Lemma 3.7.
Let us first suppose that $|\overline{y}|=1$. We shall deal with the case when $|y|>1$ later in the proof. We start by proving the result for the case where $\zeta$ is a lca-type $\textsf{type}_{\gamma,\delta}\in\textsf{Type}(k+1,d,\Lambda)$, where the variable $y$ is identified with variable $k+1$ in the definition of $\gamma$ and $\delta$. In a second stage, we’ll define $s$, $h$, $h_{s}$ and $h_{y}$ in the same way as we did in Lemma 3.7, and revisit the cases of Lemma 3.7. asdasdasd In Case 1 and 2, observe that the algorithms provided in Lemma 3.9 are algorithms that fit with the broadcast constraint of the BROADCAST CONGEST model, running in $\mathcal{O}(d)$ rounds. On input $\ell$, these algorithms output $\widehat{\ell}$. By Lemma 6.2, we can certify whether $\widehat{\ell}$ is well defined in $d\cdot\mathcal{O}(\log n)=\mathcal{O}(\log n)$ rounds.
Now let us deal with Case 3. The algorithm $\mathcal{A}$ given in Lemma 4.1 for this case has three phases. The first two phases can be implemented by an algorithm $\mathcal{B}$ under the BROADCAST CONGEST model, running in $\mathcal{O}(d)$ rounds. The output of $\mathcal{B}$ consists in a marking of all active roots, as well as a marking of all good nodes. We can simulate these two first phases by giving a certificate to each node, interpreted as the markings given by $\mathcal{B}$, as well as a certificate of the execution of $\mathcal{B}$. According to Lemma 6.2, this can be implemented using certificates on $\mathcal{O}(\log n)$ bits. The last phase of Algorithm $\mathcal{A}$ aims at providing each node with the number of active roots. Each node $u$ receives a certificate consisting of:
  • an integer $\rho(u)$ equal the number of active roots in $F$;
  • a bit $a(u)$ indicating whether or not $u$ is an active root;
  • an integer $\textsf{val}(u)$ equal to the sum of all bits $a(v)$ in the subtree of $T$ rooted at $u$.
Observe that these certificates can be encoded in $\mathcal{O}(\log n)$ bits. For a non-negative integer $\nu$, let
$h(\nu)=\begin{cases}\nu&\textsf{ if }\nu\leq k,\\ \infty&\textsf{ otherwise.}\end{cases}$
The verifier checks the following conditions at every node $u$:
  • (1)
    For every $v\in N(u)$, $\rho(u)=\rho(v)$;
  • (2)
    $a(u)=1$ if and only if $u$ is marked as active by $\mathcal{B}$;
  • (3)
    The following condition holds
    $\textsf{val}(u)=a(u)+\sum_{v\in\textsf{Children}(u)}\textsf{val}(v),$
    where $\textsf{Children}(u)$ denotes the set of nodes $v$ such that $\textsf{parent}^{T}(v)=\textsf{id}(u)$;
  • (4)
    If $\textsf{id}(u)=\textsf{root}^{T}(u)$, then $\textsf{val}(u)=\rho(u)$;
  • (5)
    If $u\in F$ is marked as good by $\mathcal{B}$, then ^$\widehat{\ell}(u)=\ell(u)\cup\{\textsf{good},h(\rho(u))\}$;
  • (6)
    If $u\in F$ is not marked as good by $\mathcal{B}$, then ^$\widehat{\ell}(u)=\ell(u)\cup\{h(\rho(u))\}$;
  • (7)
    If $u\notin F$ then ^$\widehat{\ell}(u)=\ell(u)$.
We now check completeness and soundness of this certification scheme. Condition (1) is satisfied if and only if every node received the same value $\rho$ for the number of active roots. Conditions (2) and (3) are both satisfied if and only if $\textsf{val}(u)$ correspond to the number of active roots in the subgraph of $T$ rooted at $u$. In particular $\textsf{val}(r)$ is equal to the number of active roots. Then, conditions (1) to (4) are satisfied if and only if $\rho$ corresponds to the number of active roots in $F$. Finally, conditions (1) to (7) are satisfied if and only if $\widehat{\ell}$ is well defined according to the proof of Lemma 3.7.
Let us denote by $\mathcal{A}^{\textsf{type}}$ the verification algorithm, and by $\textsf{Cert}^{\zeta}$ the certificates described for each of the three cases when $\zeta$ is an lca-type formula. Let us now assume that $\zeta$ is an arbitrary lca-reduced formula. From Lemma 3.4, there exists a set $I\subseteq~{}\textsf{Type}(k+1,D,\Lambda)$ such that,
$\zeta(\overline{x},y)=\bigvee_{\psi\in I}\psi(\overline{x},y).$
For each $\psi\in I$, let us denote by $\widehat{\Lambda}^{\psi}$, $\widehat{\psi}$ and ^$\widehat{\ell}^{\psi}$ the reduction on $\mathcal{F}_{d}$ of $\Lambda$, $\psi$ and $\ell$, respectively. The prover provides each node $u$ of $G$ with the certificate $c^{\psi}(u)=\widehat{\ell}^{\psi}(u)$ for each $\psi$, in addition to the certificate $\textsf{Cert}^{\psi}$.
The verifier proceeds as follows. Each node $u$ checks that it accepts when running algorithm $\mathcal{A}^{\textsf{type}}$ with certificates $c^{\psi}(u)$, for every $\psi\in I$. Then, each node $u$ relabels the sets ^$\widehat{\ell}^{\psi}$ to make them disjoint, as we did in the proof of Lemma 3.7), and checks whether ^$\widehat{\ell}(u)=\bigcup_{\psi\in I}\widehat{\ell}^{\psi}(u)$. The soundness and completeness of the whole scheme follows from the soundness and completeness of $\mathcal{A}^{\textsf{type}}$.
It remains to deal with the case where $q=|\overline{y}|>1$. Let us denote by $\mathcal{A}$ the verification algorithm that we described for the case when $|\overline{y}|=1$. The prover provides the nodes with a set of triplets $(\zeta_{i},\ell_{i},\textsf{Cert}_{i})_{i\in[q]}$ such that, for each $i\in[q]$, $\text{Cert}_{i}$ is a certificate used to check whether
$\textsf{true}(\exists y_{i},\zeta_{i}(\overline{x},y_{1},\dots,y_{i}),G,\ell_{i})=\textsf{true}(\zeta_{i-1}(\overline{x},y_{1},\dots,y_{i-1}),G,\ell_{i-1}),$
Where $\zeta_{q}=\zeta$, $\ell_{q}=\ell$, $\zeta_{0}=\widehat{\varphi}$, and $\ell_{0}=\widehat{\ell}$. Every node $u$ accepts if $\mathcal{A}$ accept all certificates at $u$. From the soundness and completeness of $\mathcal{A}$, we obtain that every node accepts if and only if
$\textsf{true}(\varphi(\overline{x}),G,\ell)$ $=\textsf{true}(\exists(y_{1},\dots,y_{q})\,\zeta(\overline{x},(y_{1},\dots,y_{q})),G,\ell)$
$=\textsf{true}(\exists(y_{1},\dots,y_{q})\,\zeta_{q}(\overline{x},(y_{1},\dots,y_{q})),G,\ell_{q})$
$=\textsf{true}(\exists(y_{1},\dots,y_{q-1})\,\zeta_{q-1}(\overline{x},(y_{1},\dots,y_{q-1})),G,\ell_{q-1})$
$\dots$
$=\textsf{true}(\exists y_{1}\,\zeta_{1}(\overline{x},y_{1}),G,\ell_{1})$
$=\textsf{true}(\zeta_{0}(\overline{x}),G,\ell_{0})$
$=\textsf{true}(\widehat{\varphi}(\overline{x}),G,\widehat{\ell}).$
Which completes the proof. ∎
6.2 Certifying Low-Treedepth Colorings and Skeletons
Bousquet et al. [22] have shown that, for every MSO property $\mathcal{P}$, there is a certification scheme using $\mathcal{O}(\log n)$-bit certificates for certifying $\mathcal{P}$ on graphs of bounded treedepth. Since the mere property “treedepth at most $k$” is expressible in MSO, we obtain the following.
Proposition 6.4 ([22]).
For every integer $p\geq 1$, there exists a certification scheme using $\mathcal{O}(\log n)$-bit certificates for certifying “treedepth at most $p$”.
We show how to use this result for certifying a $(p,f(p))$-treedepth coloring. Moreover, we also extend this certification scheme to an encoding of a $p$-skeleton. Given a $(p,f(p))$-treedepth coloring $C_{p}$, and $p$-skeleton $S$, the encoding of $S$ is defined as the encoding of all decomposition forests $F^{U}$ of $G[U]$, for all $U\in\textsf{Col}(p)$. More precisely, in the encoding of $S$, for every $U\in\textsf{Col}(p)$, every node $u$ has a triple $\{(\textsf{root}^{U}(u),\textsf{parent}^{U}(u),\textsf{depth}^{U}(u))\}$ corresponding to the subtree $T_{u}^{U}$ containing $u$ of the decomposition forest $F^{U}$ of $G[U]$.
Lemma 6.5.
Let $f:\mathbb{N}\to\mathbb{N}$, and let $\mathcal{G}$ be a class of graphs of expansion $f$. Let $p>0$, and let us suppose that each node $u$ is given (1) a value $C_{p}(u)\in[f(p)]$, and (2) an encoding of $S$. There is a certification scheme using $\mathcal{O}(\log n)$-bit certificates for certifying that $C_{p}$ is a $(p,f(p))$-treedepth coloring of $G$, and that the encoding of $S$ is indeed defining a $p$-skeleton of $G$.
Proof.
Let us first certify that $C_{p}$ is a $(p,f(p))$-treedepth coloring of $G$. For every set $U\in\textsf{Col}(p)$, every node $u$ receives as a certificate: (1) The encoding of a rooted spanning tree $T$ of the component of $G[U]$ containing $u$ (where each component is identified with the node identifier of the root of $T$); (2) The certification of the property “treedepth at most $p$” for all the nodes in the component of $G[U]$ containing $u$. The soundness and completeness of (1) and (2) follow from Proposition 6.1 and Proposition 6.4.
Next, let us certify that $S$ is a $p$-skeleton of $G$. First, we use the verifier of Proposition 6.1 to check that the encoding of each tree $T_{u}^{U}$ is correct, for all $u\in V$ and all $U\in\textsf{Col}(p)$. If no nodes reject at that point, we can safely deduce that if $u$ and $v$ are in the same connected component of $G[U]$, then the encoding of $T_{u}^{U}$ and of $T_{v}^{U}$ must induce the same tree.
It remains to certify that each $T^{U}_{u}$ forms a decomposition tree of the component of $G[U]$ containing $u$. More precisely, we must certify that all neighbors of $u$ in $G[U]$ are either descendants or ancestors of $u$ in $T_{u}^{U}$. Let us fix an arbitrary set $U\in\textsf{Col}(p)$. Let $(\textsf{root}^{U}(u),\textsf{parent}^{U}(u),\textsf{depth}^{U}(u))$ be the encoding of $T_{u}^{U}$ that has $u$ in its input (given in the encoding of $S$). Then node $u$ receives the set $P^{U}_{u}=\{w^{u}_{1},\dots,w^{u}_{d}\}$ as certificate, corresponding to a sequence of at most $d\leq 2^{p}$ node identifiers representing the path from $\textsf{root}(u)$ to $u$. Then node $u$ accepts $P^{U}_{u}$ if the following holds:
  • (i)
    $d=\textsf{depth}^{U}(u)$;
  • (ii)
    $w^{u}_{1}=\textsf{root}^{U}(u)$ and $w^{u}_{d}=\textsf{id}(u)$;
  • (iii)
    if $\textsf{id}(u)\neq\textsf{root}^{U}(u)$ and $v\in N(u)$ is such that $\textsf{id}(v)=\textsf{parent}^{U}(u)$, then $P^{U}_{v}$ is a sequence of length $d-1$ satisfying $w_{v}^{i}=w_{u}^{i}$ for every $i\in[d-1]$;
  • (iv)
    if every $v\in N(u)$ is such that $C_{p}(v)\in U$, then either $\textsf{id}(u)\in P^{U}_{v}$ or $\textsf{id}(v)\in P^{U}_{u}$.
Conditions (i) to (iii) are satisfied if and only if $P_{u}^{U}$ represents the path from the root of $T_{u}^{U}$ to $u$, for all $u$. Assuming that (i) to (iii) are satisfied, we have that (iv) is satisfied if and only if $T_{u}^{U}$ is a decomposition tree of $G[U]$, for all nodes $u$ of $G[U]$. Indeed, for every $u$ in $G[U]$, and for every $v\in N(u)$ that is also in $G[U]$, we have that either $u$ is an ancestor of $v$ (i.e. $\textsf{id}(u)$ belongs to $P^{U}_{v}$), or $v$ is an ancestor of $u$ (i.e. $\textsf{id}(v)$ belongs to $P^{U}_{u}$).
Then, for every $U\in\textsf{Col}(p)$, one can certify that $T^{U}_{u}$ is an decomposition tree of $G[U]$ with a certificate containing at most $2^{p}$ node identifiers, which can be encoded on $\mathcal{O}(2^{p}\cdot\log n)=\mathcal{O}(\log n)$ bits. Furthermore, one can certify that the union over all nodes $u$ and all $U\in\textsf{Col}(p)$, of $T^{U}_{u}$ is a $p$-skeleton of $G$ using certificates of size $\mathcal{O}(|\textsf{Col}(p)|\cdot 2^{p}\cdot\log n)=\mathcal{O}(\log n)$ bits. ∎
6.3 Certifying Reductions on Graphs of Bounded Expansion
Let $f:\mathbb{N}\to\mathbb{N}$, and let $\mathcal{G}$ be a class of connected graphs with expansion $f$. We first define a notion of reducibility that admits an efficient distributed certification scheme.
Definition 6.6.
Let $\varphi\in\mathrm{FO}[\Lambda]$. For every labeled graph $(G,\ell)$ with $G\in\mathcal{G}$, let us assume that every node $u$ of $G$ is provided with the following values:
  1. 1.
    The set $\Lambda$, the formula $\varphi$, and the label $\ell(u)\subseteq\Lambda$;
  2. 2.
    A positive integer $p$;
  3. 3.
    A set $\widehat{\Lambda}$, a formula $\widehat{\varphi}$, and a label ^$\widehat{\ell}(u)\subseteq\widehat{\Lambda}$;
  4. 4.
    The encoding of a $p$-skeleton $S$ of $G$.
We say that $\varphi$ admits a distributedly certifiable reduction on $\mathcal{G}$ if there is a certification scheme using $\mathcal{O}(\log n)$-bits certificates for certifying that $(p,\widehat{\Lambda},\widehat{\varphi},\widehat{\ell})$ is the reduction of $(\Lambda,\varphi,S,\ell)$ on $G$.
In a way similar to what was done in Section 4, we show that every first-order formula over labeled graphs admits a certifiable reduction. We fist establish this result for existential formulas.
Lemma 6.7.
Every existential formula in $\mathrm{FO}[\Lambda]$ admits a distributed certifiable reduction on $\mathcal{G}$.
Proof.
Let $\varphi\in\mathrm{FO}[\Lambda]$ be an existential formula. Consider that each node is provided with the values listed in Definition 6.6. Let $\varphi(\overline{x})=\exists\overline{y}\,\psi(\overline{x},\overline{y})$, where $\psi$ is quantifier free. Every node checks that $p=|\overline{x}|+|\overline{y}|$. The scheme of Lemma 6.5 enables to certify that the encoding $S$ correctly defines a $p$-skeleton of $G$. Let $C_{p}$ be the $(p,f(p))$-treedepth coloring that defines $S$. The same construction as in Lemma 4.3 can then be used. More precisely, for every $U\in\textsf{Col}(p)$, let $\widetilde{\Lambda}^{U}$, $\widetilde{\psi}^{U}$, ~$\widetilde{\ell}^{U}$, $\Lambda^{U}$, $\psi^{U}$, and $\ell^{U}$ be defined as in the proof of Lemma 4.3. Instead of computing these values, each node receives them from the prover as certificates. The values of $\widehat{\Lambda}$, $\widehat{\varphi}$, $\widetilde{\Lambda}^{U}$, $\widetilde{\psi}^{U}$, $\Lambda^{U}$, and $\psi^{U}$ do not depend on the input graph, and can be computed by each node. The labeling ~$\tilde{\ell}^{U}$ can be computed from $\ell$ by a 1-round algorithm in the BROADCAST CONGEST model. Hence, using Lemma 6.2, the correctness of each labeling ~$\tilde{\ell}^{U}$ can be certified with $\mathcal{O}(\log n)$-bit certificates. The correctness of each labeling ^$\widehat{\ell}^{U}$ can be certified using Lemma 6.3, using $\mathcal{O}(\log n)$-bit certificates. Finally, each node can check the correctness of $\widehat{\ell}$, and accept or reject accordingly. ∎
We can now proceed with the general case.
Lemma 6.8.
Every formula in $\mathrm{FO}[\Lambda]$ admits a distributed certifiable reduction on $\mathcal{G}$.
Proof.
Let $\varphi\in\mathrm{FO}[\Lambda]$ be an existential formula. Let us assume that each node is provided with the values described in Definition 6.6, i.e. the tuples $(\Lambda,\varphi,\ell)$ and $(p,\hat{\Lambda},\hat{\varphi},\hat{\ell})$, and the encoding of a $p$-skeleton $S$. We follow the same induction as in Lemma 4.3, over the quantifier depth of $\varphi$. The base case is trivial as the construction requires no communication. Furthermore, if $\varphi=\neg\psi$ where $\psi$ admits a distributed certifiable reduction, then $\varphi$ admits a distributed certifiable reduction as well, as we described in Lemma 4.3.
It remains to consieder the case $\varphi(\overline{x})=\exists y\,\psi(\overline{x},y)$. Let $(p^{\psi},\widehat{\Lambda}^{\psi},\widehat{\psi},\widehat{\ell}^{\psi}(u))$, $(\Lambda^{\xi},\xi,\ell^{\xi})$ as it was done in Case 1 in the proof of Lemma 4.5. Again, instead of computing these values, each node receives them as certificates, and checks them. The values of $p,p^{\psi},\widehat{\Lambda}^{\psi}$, $\widehat{\psi}$, $\Lambda^{\xi}$, and $\xi$ do not depend on the input graph, and can be computed by each node. Then, each node can construct the input values (1) to (4) given in Definition 6.6 with respect to to $\xi$ and $\psi$. By the induction hypothesis, the correctness of ^$\widehat{\ell}^{\psi}(u)$ at each node $u$ can be certified with $\mathcal{O}(\log n)$-bit certificates.
Recall that, in Lemma 4.5, the value of $\ell^{\xi}(u)$ on each node $u$ was computed using Lemma 4.4, without any communication, under the assumption that each node knows a encoding a $p^{\psi}$-skeleton $S$ of $G$. The prover provides the encoding of a $p^{\psi}$-skeleton to each node as a $\mathcal{O}(\log n)$-bit certificate, which can be checked using Lemma 6.5. Finally, the correctness of $\widehat{\ell}$ can be certified with $\mathcal{O}(\log n)$-bit certificates, using the scheme of Lemma 6.7 with given values $(\Lambda^{\xi},(\exists y,\xi(\overline{x},y)),\ell^{\xi})$ and $(p,\widehat{\Lambda},\widehat{\varphi},\widehat{\ell})$ and $S$. ∎
6.4 Proof of Theorem 1.5
We are now ready to prove Theorem 1.5.
Proof.
Let $\varphi\in\mathrm{FO}[\Lambda]$. We proceed the same way as in Theorem 1.2. Specifically, let us redefine $\varphi(x)$, as $\varphi$ with a dummy free-variable $x$. The prover provides each node with the following certificate:
  • (i)
    a positive integer $p$, a set $\widehat{\Lambda}$, a $p$-lca-reduced formula $\widehat{\varphi}$, and a label set ^$\widehat{\ell}(u)$;
  • (ii)
    the certificate of $u$ used to certify the $p$-skeleton $S$ of $G$;
  • (iii)
    the certificate $\textsf{Cert}(u)$ for certifying that $(p,\widehat{\Lambda},\widehat{\varphi},\widehat{\ell})$ is the reduction of $(\Lambda,\varphi,S,\ell)$.
The part (i) of the certificate can be encoded in $\mathcal{O}(1)$ bits, the part (ii) can be encoded in $\mathcal{O}(\log n)$ bits, and Lemma 6.8 asserts that a certificate for (iii) exists and, furthermore, it can be encoded in $\mathcal{O}(\log n)$ bits.
The verifier proceeds as follows. Each node checks that $(p,\widehat{\Lambda},\widehat{\varphi},\widehat{\ell})$ is the reduction of $(\Lambda,\varphi,S,\ell)$ using the certificates given in (iii) using the verifier $\mathcal{A}$ in Lemma 6.8, with certificates (i) and (ii) provided to the nodes. If a node rejects in $\mathcal{A}$, then the node rejects, otherwise it carries on the verification as described hereafter.
Assuming that no nodes have rejected so far, we have that $\textsf{true}(\varphi,G,\ell)=\textsf{true}(\widehat{\varphi},G,\widehat{\ell})$. Since $\varphi$ has only one dummy variable, the truth value of $\varphi(x)$, as well as the one of $\widehat{\varphi}(x)$, is the same for every node in $G$. Moreover, the formula $\widehat{\varphi}(x)$ is a Boolean combination of label predicates of $x$. Therefore, every node $u$ can check $\widehat{\varphi}(u)$ using ^$\widehat{\ell}(u)$, and it accepts if $\widehat{\varphi}(u)$ holds, and rejects otherwise.
It remains to analyze soundness and completeness.
For establishing completeness of the certification scheme, let us suppose that $G\models\varphi$. Then, by Lemma 4.5, we have that $\varphi$ is reducible in $\mathcal{G}$. Then there exist certificates for (i) and (ii), and, by Lemma 6.8, there exist certificates for (iii) such that every node accepts when running verifier $\mathcal{A}$. Since $G\models\varphi$ we have that $\textsf{true}(\varphi,G)=\textsf{true}(\widehat{\varphi},G,\widehat{\ell})=V(G)$. It follows that $\widehat{\varphi}(u)$ is true at every node $u$, and thus every node accepts.
For establishing soundness, let us suppose that $G\not\models\varphi$, and let us suppose that the nodes received certificates (i)-(iii) such that every node accepts when renning the verifier $\mathcal{A}$. Then $\textsf{true}(\varphi,G,\ell)=\textsf{true}(\widehat{\varphi},G,\widehat{\ell})$. However, in this case $\textsf{true}(\varphi,G,\ell)=\emptyset$. Therefore $\widehat{\varphi}(u)$ does not hold at any node $u$, and thus all nodes reject. ∎
7 Lower Bounds and Impossibility Results
In this section we give evidence that our results somehow represent the limit of tractability of distributed model checking and distributed certification of properties on graphs of bounded expansion.
All our reductions are based on classical results in communication complexity. Let $f:X\times Y\rightarrow\{0,1\}$ be a function. The deterministic communication complexity of $f$ $D(f)$ corresponds to the minimum number of bits that must be exchanged in any deterministic protocol between a player Alice holding an input $x\in X$ and another player Bob $y\in Y$ in order to output the value $f(x,y)$. The non-deterministic communication complexity of $f$, $N(f)$ is defined similarly, but for non-deterministic communication protocols. We refer to the book of Kushilevitz and Nissan for more detailed definitions [51].
In important problem in communication complexity is set disjointedness. It is defined by function $\textsf{Disj}_{n}$. For each $n$ the inputs of $\textsf{Disj}_{n}$ are subsets of $[n]$. Given $A,B\subseteq[n]$ we have that $\textsf{Disj}(A,B)=1$ if and only if $A\cap B=\emptyset$.
Proposition 7.1 ([51]).
$D(\textsf{Disj}_{n})=\Omega(n)$ and $N(\textsf{Disj}_{n})=\Omega(n)$.
7.1 No Extension to Monadic Second-Order Logic
We first show that it is not possible to extend our results to the whole set of MSO properties on graphs of bounded expansion. We illustrate this fact by considering the property of being non-three-colorable. We define Non-3-Coloring the set of graphs that do not admit a proper three-coloring (their chromatic number is at least $4$). The membership in Non-3-Coloring can be represented by the following MSO formula:
$\forall C_{1},C_{2},C_{3},\exists x,y,~{}\textsf{adj}(x,y)\wedge\big{(}(x\in C_{1}\wedge y\in C_{1})\vee(x\in C_{2}\wedge y\in C_{2})\vee(x\in C_{3}\wedge y\in C_{3})\big{)}$
In the following theorem, we show that there are no efficient CONGEST algorithms for checking nor compact certification scheme for certifying the membership in Non-3-Coloring, even under the promise that the input graph belongs to a class of graphs of bounded expansion and of logarithmic diameter.
Theorem 7.2.
Let $\mathcal{G}$ the class of graphs of maximum degree $5$ and diameter $\mathcal{O}(\log n)$. Every CONGEST algorithm deciding the membership in Non-3-Coloring of a $n$-node input graph in $\mathcal{G}$ requires $\Omega(n/\log^{2}n)$ rounds of communication.
Every certification scheme that certifies membership to Non-3-Coloring of a $n$-node input graph in $\mathcal{G}$ requires certificates of size $\Omega(n/\log n)$.
Proof.
[36] show that every locally checkable proof, and therefore every proof-labeling scheme for Non-3-Coloring requires certificates of size $\Omega(n^{2}/\log n)$. The same proof directly implies the same lowerbound for the number of rounds of a CONGEST algorithm deciding the membership in Non-3-Coloring. We give an sketch of the proof of Göös and Suomella and then we show how to adapt it to the case when the input graph has maximum degree $4$.
The proof of [36] is consists in a reduction to $\textsf{Disj}_{n^{2}}$. In the reduction each instance $(A,B)$ of $\textsf{Disj}_{n^{2}}$ is associated with a $\mathcal{O}(n)$-node graph $G_{A,B}$. The set of vertices of $G_{A,B}$ is divided in two sets of vertices, denoted $V_{A}$ and $V_{B}$, and satisfies the following conditions:
  • (C1)
    The graph induced by $V_{A}$ only depends on $A$.
  • (C2)
    The graph induced by $V_{B}$ only depends on $B$.
  • (C3)
    The set of edges $E_{A,B}$ with one endpoint in $V_{A}$ and the other in $V_{B}$ contains at most $\mathcal{O}(\log n)$ edges, and does not depend on $A$ and $B$, only on $n$. The nodes with endpoints in $E_{A,B}$ are denoted $V_{A,B}$.
  • (C4)
[…]
The construction is also useful for a lowerbound in the CONGEST model. Let $\mathcal{A}$ be an algorithm deciding the membership in Non-3-Coloring in $\mathcal{R}(n)$ rounds. Then Alice and Bob construct $G_{A,B}$ as above, and simulate the $\mathcal{R}(n)$ rounds of $\mathcal{A}$ on it, communicating on each round to other player the messages transmitted over the edges $E_{A,B}$. The players accept if all the nodes of their sides accept. On each communication round the players send $\mathcal{O}(\log n)$ bits per edge of $E_{A,B}$. Hence, the nodes interchange $\mathcal{O}(R(n)\cdot|E_{A,B}|\cdot\log n)=\mathcal{O}(R(n)\cdot\log^{2}n)$ bits in total. From Proposition 7.1 we obtain that $\mathcal{O}(R(n)\cdot\log^{2}n)=D(\textsf{Disj}_{n^{2}})=\Omega(n^{2})$. Therefore, $R(n)=\Omega(n^{2}/\log^{2}n)$.
[…]
$\ v$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{1}$$v_{1}$$v_{2}$$w_{2}$$v_{3}$$w_{3}$$v_{5}$$v_{4}$$w_{4}$$w_{5}$
Figure 3: Node gadget. In this figure is depicted the gadget for a node of degree $5$. Notice that in every proper three-coloring of the gadget the nodes $v_{1},\dots,v_{5}$ receive the same colors. Conversely, if we color the nodes $v_{1},\dots,v_{5}$ with the same color, we can extend the coloring into a three-coloring of all the nodes of the gadget.
[…]
We now reduce the diameter constructing a graph $\widehat{G}_{A,B}$ obtained from $\widetilde{G}_{A,B}$ by plugging to it two rooted binary trees $T_{A}$ and $T_{B}$, where each edge of $T_{A}$ and $T_{B}$ is subdivided. Tree $T_{A}$ has depth $\mathcal{O}(\log|\widetilde{V_{A}}|)$, and each leaf corresponds to a node of $\widetilde{V_{A}}$. The tree $T_{B}$ is defined analogously for $\widetilde{V}_{B}$. Finally, add an edge between the root of $T_{A}$ and the root of $T_{B}$. Let us define $\widehat{V}_{A}=\widetilde{V}_{A}\cup V(T_{A})$ and $\widetilde{V}_{B}=\widetilde{V}_{B}\cup V(T_{B})$. We obtain that $\widehat{G}_{A,B}$ also satisfies conditions (C1)-(C4) and belongs to $\mathcal{G}$. Conditions C1 and C2 follow from the construction of $\widehat{V}_{A}$ and $\widehat{V}_{B}$. Condition C3 holds because in $\widehat{G}_{A,B}$ there is only one more edge in the cut than $\widetilde{G}_{A,B}$ consisting in the edge between the roots. Finally, for condition (C4) we observe that any $3$-coloring of the nodes in $\widetilde{G}_{A,B}$ can be extended to a three coloring of the trees $T_{A}$ and $T_{B}$. First, pick two different colors for the roots of $T_{A}$ and $T_{B}$. Then, color arbitrarily the other nodes of degree $3$ in $T_{A}$ and $T_{B}$. Then, color the nodes of degree two (the subdivided edges) with a color different than their endpoints. We deduce that $\widehat{G}_{A,B}$ satisfies conditions (C1)-(C4) and belongs to $\mathcal{G}$.
Let $\mathcal{P}$ be a certification scheme (resp. let $\mathcal{A}$ be a CONGEST algorithm) for Non-3-Coloring with certificates of size $c(n)$ (resp. running in $R(n)$ rounds) under the promise that the input graph is a $n$-node graph contained in $\mathcal{G}$. Let $A,B\subseteq[n]$ be an instance of $\textsf{Disj}_{n^{2}}$. Consider the communication protocol where Alice and Bob construct $\widehat{G}_{A,B}$ and simulate $\mathcal{P}$ (resp. $\mathcal{A}$) in the same way that we explained for $G_{A,B}$. Since $\widehat{G}_{A,B}$ is a graph with $\Theta(n^{2})$ nodes, we obtain that $\mathcal{O}(c(n^{2})\cdot\log n)=N(\textsf{Disj}_{n^{2}})=\Omega(n^{2})$, hence $c(n^{2})=\Omega(n^{2}/\log n)$, and then $c(n)=\Omega(n/\log n)$. Similarly, we obtain that $R(n)=\Omega(\sqrt{n}/\log^{2}(n))$. ∎
Remark.
Theorem 7.2 implies that the difficulty of certifying MSO properties over graphs of max degree 5 is (up to logarithmic factors) as hard as the hardest problem over that family. Indeed, a graph of max degree $5$ can be encoded using $\mathcal{O}(n\log n)$ bits, e.g. by adjacency lists of all nodes. We can use such encoding to design a certification scheme that certifies any property on graphs of maximum degree $5$: every node interprets its certificate as an encoding of a graph of maximum degree $5$, and checks that every neighbor received the same encoding, and that its local view corresponds with the one in the encoding. If the conditions are verified, the nodes can compute any property of the input graph without any further communication.
We can also show that we cannot extend the certification of first-order properties to monadic second-order properties for planar graphs. However, in this case we obtain a weaker lowerbound.
Theorem 7.3.
Every certification scheme for membership to Non-3-Coloring under the promise that the input graph is planar requires certificates of size $\Omega(\sqrt{n}/\log n)$ bits.
Similar to the proof of Theorem 7.2 we use the construction of Göös and Suomella and the reduction of Garey, Johnson and Stockmeyer for planar graphs. Given an $n$-node graph $G$, the reduction of [30] for planar graphs first pick an arbitrary embedding of $G$ in the plane. In this embedding, there might be a number $\textsf{cr}(G)$ of edge crossings (they must exist if $G$ is non-planar). Then, a planar graph $\widetilde{G}$ is constructed from $G$ by replacing on point where a pair of edges cross, by the crossing gadget shown in Figure 4. This gadget allows to transmit the coloring of one of the endpoints of the edge in a planar way. For each crossing, we add 13 new vertices. In the worst case, $\textsf{cr}(G)=\Theta(n^{4})$. Hence, the obtained graph has $\widetilde{G}$ has $\Theta(n^{4})$ nodes.
Proof.
$\ x_{2}$$\ x_{1}$$\ y_{1}$$\ y_{2}$
Figure 4: Crossing gadget. In every proper three-coloring of the gadget node label $x_{1}$ has the same color of node labeled $x_{2}$, as well as node labeled $y_{1}$ has the same color of $y_{2}$. Conversely, every three-coloring of $x_{1},x_{2},y_{1}$ and $y_{2}$ that assigns the same color to $x_{1}$ and $x_{2}$, and the same color to $y_{1}$ and $y_{2}$ can be extended into a three coloring of the gadget.
Let $\mathcal{G}$ be the class of planar graphs. As we did for Theorem 7.2, we plug the reduction construction of Garey, Johnson and Stockmeyer to the construction of Göös and Suomella. More precisely, given an instance $(A,B)$ of $\textsf{Disj}_{n}$, we build graph $G_{A,B}$ (c.f. proof of Theorem 7.2). Then, using the crossing gadget we obtain a planar graph $\widetilde{G}_{A,B}$. We partition the nodes of $\widetilde{G}_{A,B}$ into $\widetilde{V}_{A}$ and $\widetilde{V}_{B}$. All the nodes inside gadgets created for edges with both endpoint in $V_{A}$ (respectively $V_{B}$) belong to $\widetilde{V}_{A}$ (respectively $\widetilde{V}_{B}$). We also consider that the nodes inside the gadgets used for the edges in the set $E_{A,B}$ of $G_{A,B}$ belong to $V_{A}$. We obtain that $\widetilde{G}_{A,B}$ satisfies conditions (C1)-(C4).
The remaining of the proof is analogous to Theorem 7.2. Let $\mathcal{P}$ be a certification scheme for Non-3-Coloring with certificates of size $c(n)$ under the promise that the input graph is planar. On input $A,B\subseteq[n]$, Alice and Bob construct $\widetilde{G}_{A,B}$ and simulate $\mathcal{P}$ in the same way that we explained for $G_{A,B}$. Since $\widetilde{G}_{A,B}$ is a graph with $\Theta(n^{4})$ nodes, we obtain that $\mathcal{O}(c(n^{4})\cdot\log n)=N(\textsf{Disj}_{n^{2}})=\Omega(n^{2})$, hence $c(n^{4})=\Omega(n^{2}/\log n)$, and then $c(n)=\Omega(\sqrt{n}/\log n)$. ∎
7.2 No Extension to First Order Logic on Graphs of Bounded Degeneracy
[…]
Theorem 7.4.
[…]
Proof.
The result for CONGEST algorithms and graph of degeneracy $2$ was already shown in Korhonen et al. [47], which in turn use a construction of Druker et al. [17]. Roughly, the lower-bound is based on a reduction to $\textsf{Disj}_{m}$, for $m=\Theta(n^{1+1/2})$. Each instance $A,B\in[m]$ is associated with a $\mathcal{O}(n)$-node graph $G_{A,B}$ of degeneracy $2$. The set of vertices of $G_{A,B}$ are divided in two sets of vertices, denoted $V_{A}$ (and called the nodes owned by Alice) and $V_{B}$ (the nodes owned by Bob), satisfying the following conditions:
  • (B1)
    The graph induced by $V_{A}$ only depends on $A$.
  • (B2)
    The graph induced by $V_{B}$ only depends on $B$.
  • (B3)
    The set of edges $E_{A,B}$ with one endpoint in $V_{A}$ and the other in $V_{B}$ contains at most $\mathcal{O}(n)$ edges, and does not depend on $A$ and $B$.
  • (B4)
[…]
8 Conclusion
In centralized model checking, the fixed-parameter tractability of FO was extended from graphs of bounded expansion to the more general class of nowhere dense graphs in [39]. Notably, nowhere dense graphs represent the ultimate frontier in the program of designing fixed-parameter tractable algorithms for model checking FO formulas. Indeed, as shown in [18], for any graph class $\mathcal{G}$ that is not nowhere dense (i.e., $\mathcal{G}$ is somewhere dense) and closed under taking subgraphs, FO model checking on $\mathcal{G}$ is not fixed-parameter tractable (unless $\text{FPT}=\text{W}[1]$). The tractability frontier for FO model checking in the CONGEST model remains unknown. Moreover, there is no reason to expect that the tractability landscape in the distributed setting should mirror the one of the centralized setting. This raises several concrete and intriguing questions, discussed hereafter.
Is it possible to design model-checking algorithms for FO formulas running in $\tilde{O}(D)$ rounds in CONGEST on nowhere dense graphs of diameter $D$? From a theoretical perspective, a negative answer to this question would be even more insightful than a positive one. More broadly, what is the precise boundary for $\tilde{O}(D)$-round FO model checking in the CONGEST model? Does it coincide with the class of bounded expansion graphs, extend to nowhere dense graphs, or even include a larger family of graphs? There are several difficulties in extending our approach to nowhere dense graphs. First, it is unclear how to implement the combinatorial arguments from [39], particularly the constructions of sparse neighborhood covers, in a distributed model. Second, the proof of [39] builds on a weak variant of Gaifman’s local form, and not a quantifier elimination procedure. While this is sufficient for fixed-parameter tractability of FO model-checking, we do not know whether it could be adapted in the distributed setting.
Another question is whether a meta-theorem with $\tilde{O}(D)$ round complexity can be obtained for some extensions of FO logic on planar graphs. Theorem 7.2 shows that we cannot hope to extend Theorem 1.2 to MSO, even on planar graphs. On the other hand, in centralized computing, the fixed-parameter tractability of logics lying strictly between FO and MSO is an active area of research. For example, the extension of FO known as separator logic, denoted FOL+conn, was introduced independently by Schirrmacher, Siebertz, and Vigny [60], and by Bojańczyk and Pilipczuk [3]. This logic extends FO, for every $k\geq 1$, with the general predicate $\mathsf{conn}_{k}(x,y,z_{1},\ldots,z_{k})$, which evaluates to true on a graph $G$ if the nodes corresponding to $x$ and $y$ are connected by a path that avoids the nodes corresponding to $\{z_{1},\ldots,z_{k}\}$. Another example is disjoint paths logic, denoted FOL+DP, which extends FO with the atomic predicate $\mathsf{dp}_{k}(x_{1},y_{1},\ldots,x_{k},y_{k})$, expressing the existence of $k$ pairwise internally vertex-disjoint paths between $x_{i}$ and $y_{i}$ for $i\in\{1,\ldots,k\}$. The main obstacle to developing distributed meta-theorems for these logics lies in the absence of distributed counterparts of the core algorithmic tools used in the centralized setting—such as tree decompositions which underlies tractability results for FOL+conn [58], or the irrelevant vertex technique used for FOL+DP [35].